### v1

# Misspecification uncertainties in near-deterministic regression

Abstract
The expected loss is an upper bound to the model generalization error which admits robust PAC-Bayes bounds for learning. However, loss minimization is known to ignore misspecification, where models cannot exactly reproduce observations. This leads to significant underestimates of parameter uncertainties in the large data, or underparameterized, limit. We analyze the generalization error of near-deterministic, misspecified and underparametrized surrogate models, a regime of broad relevance in science and engineering. We show posterior distributions must cover every training point to avoid a divergent generalization error and derive an ensemble ansatz that respects this constraint, which for linear models incurs minimal overhead. The efficient approach is demonstrated on model problems before application to high dimensional datasets in atomistic machine learning. Parameter uncertainties from misspecification survive in the underparametrized limit, giving accurate prediction and bounding of test errors.
I Introduction
Surrogate models are widely used across science and engineering, notably to efficiently approximate the action of computationally intensive simulation engines  (Alizadeh et al., 2020; Deringer et al., 2019; Lapointe et al., 2020; Nyshadham et al., 2019; Montes de Oca Zapiain et al., 2022; Bonatti et al., 2022). Domain expertise is leveraged to design optimal features and model architecture before parameter inference is carried out based on training data generated by the engines. In this paper, we consider the common scenario where the learning problems share three key characteristics:
  • simulation engines are near deterministic: simulation outputs have vanishing aleatoric uncertainty. For example, the atomic energy $V({\bf X})$ is a deterministic function of atomic positions ${\bf X}$.
  • surrogate models are misspecified: no single choice of the $P$ model parameters can match all $N$ training observations, such that model averaging is required to capture uncertainties.
[…]
Misspecification affects both finite capacity models and deep learning approaches with finite training resources  (Lahlou et al., 2021; Psaros et al., 2023). A misspecification-aware learning scheme should minimize the cross entropy between predicted and observed data distributions, known as the generalization error or predictive risk (see (Alquier, 2021; Morningstar et al., 2022)), but this loss is typically numerically intractable for high dimensional models.
Bayesian learning schemes instead minimize the expected loss (log likelihood), an upper bound to the generalization error, which is known to ignore misspecification. Epistemic uncertainties are therefore typically severely underestimated and can be shown to vanish in the underparametrized limit, with broad implications for surrogate model selection, uncertainty quantification and error propagation, as noted by multiple groups (Masegosa, 2020; Kato et al., 2022; Lotfi et al., 2022; Lahlou et al., 2021; Psaros et al., 2023; Imbalzano et al., 2021). We analyze the generalization error of deterministic, misspecified and under-parametrized surrogate models, i.e. those typically used to approximate simulation engines in science and engineering. The deterministic limit is atypical in Bayesian analysis, as log likelihoods are typically required to be bounded and smooth, requiring a finite aleatoric uncertainty. Here, we take the deterministic limit to derive conditions that any parameter distribution must obey to avoid a divergent generalization error, then derive a simple ensemble ansatz which admits a robust learning scheme. Our main contributions are:
  • We define a pointwise optimal parameter sets (POPS) for each training point, for which model predictions are exact. Parameter distributions must have mass in every POPS to avoid a divergent generalization error. An ensemble ansatz is proposed that respects POPS occupancy and gives a finite generalization error.
  • For linear models, our ansatz can be efficiently evaluated via rank-one updates to a leverage-weighted loss minimizer. This is our main result, an efficient scheme to incorporate misspecification into linear regression with minimal computational overhead.
  • We compare our ansatz to direct numerical minimization of the generalization error, showing how our scheme provides robust bounds on test errors. This is confirmed in numerical experiments on challenging datasets from atomic machine learning.
(Germain et al., 2016) established an important connection between Bayesian inference and probability approximately correct (PAC-Bayes) analysis, which bounds true expectations by empirical likelihood expectations through concentration inequalities. This provides a robust learning scheme without test-train splitting, and gives a rationalization for Bayesian inference as a form of regularization. In particular, minimizing the PAC-Bayes bound for the expected loss gives the familiar relation between posterior, prior and likelihood from Bayesian inference.
In recent years, multiple groups have used the PAC-Bayes framework to derive tighter bounds for the generalization error than that provided by the expected loss  (Masegosa, 2020; Germain et al., 2016; Lahlou et al., 2021; Psaros et al., 2023; Lotfi et al., 2022; Morningstar et al., 2022). However, these efforts focus on probabilistic regression settings, often with neural network models far from the underparametrized limit, meaning aleatoric, epistemic and misspecification errors must be considered jointly.
(Masegosa, 2020) derived second order PAC-Bayes bounds for the generalization error, minimized through a variational or ensemble ansatz. (Morningstar et al., 2022) developed estimators to assess disagreement between the generalization error and expected loss, deriving specialized PAC bounds for theoretical guarantees. (Lahlou et al., 2021) considered misspecification of neural networks, jointly training a minimum loss surrogate and an independent predictor of misspecification error. (Lotfi et al., 2022) considered misspecification in the context of deep model selection, deriving conditional likelihoods better aligned with the generalization error.
As misspecification has received only recent attention in the literature, various definitions for the distinction between epistemic and misspecification uncertainties exist (Kato et al., 2022). Here, misspecification is defined as non-aleatoric uncertainties which survive in the underparametrized limit $N/P\to\infty$, where PAC-Bayes bounds for the expected loss become tight. Our epistemic uncertainties thus coincide with standard estimates from loss minimization.
II Bayesian surrogate models
II.1 Deterministic Simulation Engines
A simulation engine ${\mathcal{E}}:{\mathcal{X}}\to{\mathcal{Y}}$ takes input ${\mathbf{X}}\in{\mathcal{X}}$ and returns ${\mathbf{Y}}\in{\mathcal{Y}}$. We denote the dimension of ${\mathcal{Y}}$ and ${\mathcal{X}}$ by $Y$ and $X$ respectively. We consider the regime where the aleatoric uncertainty of ${\mathcal{E}}$ is weak and can be described as a Gaussian distribution with mean $\langle{\mathbf{Y}}|{\mathbf{X}}\rangle\equiv{\mathcal{E}}({\mathbf{X}})$ and covariance $\langle{\mathbf{Y}}\otimes{\mathbf{Y}}|{\mathbf{X}}\rangle-{\mathcal{E}}({\mathbf{X}})\otimes{\mathcal{E}}({\mathbf{X}})\equiv{\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})$, where $\|{\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})\|=\mathcal{O}(\epsilon^{2Y})$, and $\epsilon$ is assumed small. The true conditional output distribution thus reads
$\rho_{\mathcal{E}}({\mathbf{Y}}|{\mathbf{X}})=\mathcal{N}({\mathbf{Y}}|{\mathcal{E}}({\mathbf{X}}),{\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})).$ (1)
We assume homoskedatiscity ${\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})={\bm{\Sigma}}_{\mathcal{Y}}$ of this weak aleatoric error to investigate the deterministic, underparametrised limit $\epsilon\to 0$, $N/P\to\infty$ for $\epsilon=\mathcal{O}(\sqrt{P/N})$. With weights $\sum_{i=1}^{N}{\rm w}_{i}=1$, the training configurations ${\mathcal{D}}_{N}=\{{\mathbf{X}}_{i}\}_{i=1}^{N}$ form an input distribution
$\rho_{N}({\mathbf{X}})\equiv\sum_{i}{\rm w}_{i}\delta({\mathbf{X}}-{\mathbf{X}}_{i}),\,\,\rho_{{\mathcal{D}}}({\mathbf{X}})\equiv\lim_{N\to\infty}\rho_{N}({\mathbf{X}}).$ (2)
The non-parametric estimate of $\rho_{N}$ will suffer from a curse of dimensionality (Tsybakov, 2008), i.e. the worst case error scales as $N^{1/(1+{\rm Dim}({\mathbf{X}}))}$$\simeq$1. However, learning only requires that we can bound expectations over independent identically distributed (i.i.d.) samples ${\mathbf{X}},{\mathbf{Y}}\in{\mathcal{D}}$
$\langle f({\mathbf{Y}},{\mathbf{X}})\rangle_{{\mathcal{D}}}\equiv\int_{\mathcal{D}}f({\mathbf{Y}},{\mathbf{X}})\rho_{\mathcal{E}}({\mathbf{Y}}|{\mathbf{X}})\rho_{{\mathcal{D}}}({\mathbf{X}}){\rm d}{\mathbf{X}}{\rm d}{\mathbf{Y}}.$ (3)
For timeseries or other signal data the i.i.d. property is only satisfied for decorrelated segments ${\mathbf{X}}=\{{\bf x}_{t}\}_{t=1}^{T}$, meaning ${\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})$ captures signal correlations as in generalized least squares regression.
II.2 Deterministic Surrogate Models and the Generalization Error
A surrogate model ${\mathcal{M}}:{\mathcal{X}}\to{\mathcal{Y}}$ for ${\mathcal{E}}$ is defined by $P$ parameters ${\bm{\Theta}}\in{\mathcal{P}}$. Under a parameter distribution $\pi({\bm{\Theta}})$, the predicted distribution has the same form as (1), reading
$\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})$ $=\mathcal{N}({\mathbf{Y}}|{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}}),{\bm{\Sigma}}_{\mathcal{Y}}),$ (4)
$\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}})$ $=\int_{\mathcal{P}}\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}.$ (5)
Unlike typical Bayesian schemes, we do not treat the covariance term ${\bm{\Sigma}}_{\mathcal{Y}}$ as fitting parameter but as an intrinsic property of the quasi-deterministic simulation engine ${\mathcal{E}}$. This has implications for model selection criteria, as we discuss in appendix A.
The generalization error ${\mathcal{G}}[\pi]$ is simply the cross entropy of predicted to observed distributions. The $\rho_{N}({\mathbf{X}})$ terms cancel inside the log and, omiting the constant ${\mathcal{G}}_{\mathcal{E}}\equiv\left\langle\ln\left|\rho_{\mathcal{E}}({\mathbf{Y}}|{\mathbf{X}})\right|\right\rangle_{\mathcal{D}}$, we have
${\mathcal{G}}[\pi]=-\left\langle\ln\left|\int_{\mathcal{P}}\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}\right|\right\rangle_{\mathcal{D}},$ (6)
PAC-Bayes analysis (Alquier, 2021) provides concentration inequalities (Hoeffding, 1994) that bound expectations (3) over ${\mathcal{D}}$ by empirical expectations over ${\mathcal{D}}_{N}$. These require $f({\mathbf{Y}},{\mathbf{X}})$ to be bounded (i.e. $\epsilon>0$) and expressed as an expectation over $\pi({\bm{\Theta}})$, i.e. $f({\mathbf{Y}},{\mathbf{X}})=\int_{\mathcal{P}}f({\mathbf{Y}},{\mathbf{X}},{\bm{\Theta}})\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}$. As ${\mathcal{G}}[\pi]$ is not in this form, PAC concentration inequalities can only be applied to some upper bound for ${\mathcal{G}}[\pi]$ which is an expectation over $\pi({\bm{\Theta}})$.
II.3 The expected loss and PAC estimation
The bound used in Bayesian schemes is for the expected loss ${\mathcal{L}}[\pi]\geq{\mathcal{G}}[\pi]$, often known as the negative log likelihood. The Jensen inequality for convex functions $f(\langle x\rangle)\leq\langle f(x)\rangle$ gives, with $f(x)=-\ln|x|$,
$\mathcal{G}[\pi]\leq{\mathcal{L}}[\pi]=-\int_{\mathcal{P}}\left\langle\ln\left|\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\right|\right\rangle_{\mathcal{D}}\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}.$ (7)
The empirical expected loss ^$\hat{{\mathcal{L}}}_{N}[\pi]$ over $\rho_{N}({\mathbf{X}})$ can estimate the true expected loss ${\mathcal{L}}[\pi]$ through the PAC-Bayes concentration inequality, as shown by (Germain et al., 2016)
${\mathcal{L}}[\pi]\leq\hat{{\mathcal{L}}}_{N}[\pi]$ $\equiv\int_{\mathcal{P}}\hat{L}_{N}({\bm{\Theta}})\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}$
$+\frac{1}{N}\int_{\mathcal{P}}\ln\left|\frac{\pi({\bm{\Theta}})}{\pi_{0}({\bm{\Theta}})}\right|\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}+\frac{C_{0}}{N}$ (8)
where $\hat{L}_{N}({\bm{\Theta}})=-\langle\ln|\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})|\rangle_{N}$ is the empirical loss over $\rho_{N}({\mathbf{X}})$, the distribution $\pi_{0}({\bm{\Theta}})$ is to some prior and $C_{0}$ is independent of $\pi$ but depends on $\epsilon$ and $\pi_{0}$. A well chosen prior thus tightens the PAC bound. Minimization of ^$\hat{{\mathcal{L}}}_{N}[\pi]$ yields the posterior familiar from Bayesian inference, as shown by  (Germain et al., 2016), but with an additional term $\lambda$ that ensures normalization-
$\pi_{{\mathcal{L}},N}({\bm{\Theta}})=\pi_{0}({\bm{\Theta}})\exp(\lambda-N\hat{L}_{N}({\bm{\Theta}})).$ (9)
Appendix A discusses this bound in the context of Bayesian model selection for deterministic models. Using Laplace’s method as $N/P\to\infty$ gives $\hat{L}_{N}({\bm{\Theta}})\to L({\bm{\Theta}})$, with
[…]
$L({\bm{\Theta}})=-\langle\ln|\mathcal{N}({\mathcal{E}}({\mathbf{X}})|{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}}),{\bm{\Sigma}}_{\mathcal{Y}})|\rangle_{\mathcal{X}},$ (11)
where $\langle\dots\rangle_{\mathcal{X}}\equiv\int\dots\rho_{\mathcal{D}}({\mathbf{X}}){\rm d}{\mathbf{X}}$. For linear problems, ${\bm{\Theta}}^{*}_{\mathcal{L}}$ is a generalized least squares solution, with the minimizing set a connected region in $\mathbb{R}^{P}$. In the underparametrized limit $N/P\to\infty$, expected loss is thus minimized by the sharp distribution $\pi_{\mathcal{L}}({\bm{\Theta}})\to\delta({\bm{\Theta}}-{\bm{\Theta}}^{*}_{\mathcal{L}})$, giving vanishing parameter uncertainty. The generalization error of the minimium loss solution clearly diverges as $1/\epsilon^{2}$:
$\lim_{N/P\to\infty}{\mathcal{G}}[\pi_{\mathcal{L}}]=\frac{1}{2}{\rm Tr}({\bm{\Sigma}^{*}_{\mathcal{L}}}{\bm{\Sigma}}_{\mathcal{Y}}^{-1})=\mathcal{O}(1/\epsilon^{2}),$ (12)
where ${\bm{\Sigma}}^{*}_{\mathcal{L}}=\langle[{\mathcal{E}}({\mathbf{X}})-{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}}^{*}_{\mathcal{L}})]\otimes[{\mathcal{E}}({\mathbf{X}})-{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}}^{*}_{\mathcal{L}})]\rangle$ is the error covariance around ${\bm{\Theta}}^{*}_{\mathcal{L}}$ and we use ${\rm Tr}({\bm{\Sigma}}_{\mathcal{Y}})=\mathcal{O}(\epsilon^{2})$. As noted by (Masegosa, 2020), it is clear that the maximum likelihood parameter distribution $\pi_{\mathcal{L}}({\bm{\Theta}})$ is not a minimizer of the generalization error under misspecification.
III Analysis of the generalization error in the near-deterministic limit
We consider the generalization error (6) in the near-deterministic limit, ${\mathcal{G}}_{0}[\pi]\equiv\lim_{\epsilon\to 0}{\mathcal{G}}[\pi]$. The integral over ${\mathbf{Y}}$ in (6) concentrates at ${\mathbf{Y}}={\mathcal{E}}({\mathbf{X}})$; using $\langle\dots\rangle_{\mathcal{X}}\equiv\int\dots\rho_{\mathcal{D}}({\mathbf{X}}){\rm d}{\mathbf{X}}$ yields, using Laplace’s method
${\mathcal{G}}_{0}[\pi]=-\left\langle\ln\left|\int_{\mathcal{P}}\rho_{\mathcal{M}}({\mathcal{E}}({\mathbf{X}})|{\mathbf{X}},{\bm{\Theta}})\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}\right|\right\rangle_{\mathcal{X}}.$
Whilst ${\mathbf{Y}}={\mathcal{E}}({\mathbf{X}})$ is a unique maximum, as $\epsilon\to 0$ the marginal distribution (5) concentrates ${\bm{\Theta}}$ over pointwise optimal sets (POPS) of parameters
${\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})\equiv\left\{{\bm{\Theta}}\,|{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}})={\mathcal{E}}({\mathbf{X}})\right\}.$ (13)
Any model with a constant term will ensure ${\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})\neq\emptyset$, but under misspecification it clear that $\cap_{{\mathbf{X}}\in{\mathcal{X}}}{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})=\emptyset$. The conditional distribution $\rho_{\mathcal{M}}({\mathcal{E}}({\mathbf{X}})|{\mathbf{X}},{\bm{\Theta}})$ will sharpen as $\epsilon\to 0$, introducing a Jacobian into the integral over ${\bm{\Theta}}$. This can be contained in a pointwise mass function
${\rm m}_{\mathcal{E}}({\bm{\Theta}},{\mathbf{X}})$ $\equiv\int_{\mathcal{P}}\delta({\bm{\Theta}}-{\bm{\Theta}}^{\prime})\rho_{\mathcal{M}}({\mathcal{E}}({\mathbf{X}})|{\mathbf{X}},{\bm{\Theta}}^{\prime}){\rm d}{\bm{\Theta}}^{\prime}$
$=\int_{{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})}\frac{\delta({\bm{\Theta}}-{\bm{\Theta}}^{\prime})}{\sqrt{\|{\bm{\nabla}}_{\bm{\Theta}}{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}}^{\prime})\|}}{\rm d}{\bm{\Theta}}^{\prime}.$ (14)
This gives our first main result, a generalization error for deterministic surrogate models
${\mathcal{G}}_{0}[\pi]\equiv-\left\langle\ln\int_{\mathcal{P}}{\rm m}_{\mathcal{E}}({\bm{\Theta}},{\mathbf{X}})\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}\right\rangle_{\mathcal{X}}.$ (15)
To avoid $1/\epsilon^{2}$ divergence in ${\mathcal{G}}_{0}[\pi]$ a valid $\pi({\bm{\Theta}})$ must satisfy the ‘covering’ constraint as $\epsilon\to 0$ of
$\min_{{\mathbf{X}}\in{\mathcal{X}}}\max_{{\bm{\Theta}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})}\pi({\bm{\Theta}})>\lim_{\epsilon\to 0}\exp(-c/\epsilon^{2}),$ (16)
for any $c>0$. Equation (16) is our second main result, a demonstration that any candidate minimizer of the generalization error must have probability mass in every POPS ${\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})$, whilst the minimizing distribution concentrates on regions where multiple POPS intersect. As shown above, the minimum loss solution $\pi_{\mathcal{L}}({\bm{\Theta}})$ does not satisfy this requirement and thus ${\mathcal{G}}[\pi_{\mathcal{L}}]=\mathcal{O}(1/\epsilon^{2})$. Importantly, when (16) is satisfied, the predictive distribution $\rho_{\mathcal{M}}({\mathbf{Y}},{\mathbf{X}})$ covers every training data point, i.e. model predictions are guaranteed to envelope observations as $N/P\to\infty$.
III.1 POPS-constrained loss minimizing ensemble ansatz
The covering constraint (16) shows, in the limit $\epsilon\rightarrow 0$, the minimizer of ${\mathcal{G}}_{0}$ concentrates onto the union of POPS for each training point. This suggests that a representation in terms of a discrete set of models is likely to capture its key features. Selecting some ${\bm{\Theta}}_{\mathbf{X}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})$ for each ${\mathbf{X}}\in{\mathcal{X}}$ and possibly some weighting ${\rm w}({\mathbf{X}})>0$, equation (16) is satisfied by the ensemble ansatz
$\pi_{E}({\bm{\Theta}})\equiv\langle{\rm w}({\mathbf{X}})\delta({\bm{\Theta}}-{\bm{\Theta}}_{\mathbf{X}})\rangle_{\mathcal{X}}.$ (17)
It is simple to show the generalization error is finite and bounded by
${\mathcal{G}}_{0}[\pi_{E}]\leq\left\langle\ln\left|{\rm w}({\mathbf{X}}){\rm m}_{\mathcal{E}}({\bm{\Theta}}_{\mathbf{X}},{\mathbf{X}})\right|\right\rangle_{\mathcal{X}},$ (18)
and the expected loss is ${\mathcal{L}}[\pi_{E}]=\langle{\rm w}({\mathbf{X}})L({\bm{\Theta}}_{\mathbf{X}})\rangle_{\mathcal{X}}$. However, whilst ${\mathcal{G}}_{0}[\pi_{E}]$ is finite, the vast majority of the pointwise fits ${\bm{\Theta}}_{\mathbf{X}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})$ will give very poor predictions away from ${\mathbf{X}}$. We therefore select each pointwise fit to minimize the (POPS-constrained) expected loss ${\mathcal{L}}[\pi_{E}]$, i.e.
${\bm{\Theta}}^{*}_{\mathbf{X}}\equiv\arg\min_{{\bm{\Theta}}_{\mathbf{X}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})}L({\bm{\Theta}}_{\mathbf{X}}).$ (19)
The goal of this POPS-constrained loss-minimization is to ensure the pointwise fits will be as close as possible; for convex loss landscapes, this is measured by the Fisher information metric ${\bm{\mathcal{I}}}^{*}_{\mathcal{L}}$. By minimizing the ensemble spread, we aim to distribute an (albeit small) mass to other POPS, thereby reducing the generalization error. In addition, for specified regression problems $\pi_{E}({\bm{\Theta}})$ then reduces to the true minimizer $\pi^{*}_{\mathcal{L}}({\bm{\Theta}})=\delta({\bm{\Theta}}-{\bm{\Theta}}^{*}_{\mathcal{L}})$.
When the loss landscape is not convex additional constraints will have to be imposed to minimize (and measure) the ensemble spread. More generally, POPS-constrained loss minimization with our ensemble ansatz is potentially sub-optimal in that some other ${\bm{\Theta}}_{\mathbf{X}}$, which are not loss minimizers, could be members of multiple POPS, i.e. ${\bm{\Theta}}_{\mathbf{X}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}}^{\prime})$, which is a priori beneficial to minimizing the generalization error. The loss minimization scheme (19) will minimize the average prediction error of each selected ${\bm{\Theta}}_{\mathbf{X}}$ across all other ${\mathbf{X}}^{\prime}\in{\mathcal{X}}$. Exploring this tradeoff (multiple POPS occupancy vs global prediction error) would entail a combinatorial scheme which our simple ensemble ansatz avoids. A key limitation of (19) is that many pointwise fits may be near-degenerate, raising the generalization error and, in general, underestimating the average deviation of any test point from the minimum loss prediction. However, the predicted max/min range of any test point is expected to be robust in the limit $N/P\to\infty$, as this is invariant to how we weight the ensemble.
Furthermore, even if it were possible to find the true minimizer $\pi^{*}_{\mathcal{G}}({\bm{\Theta}})$ of ${\mathcal{G}}_{0}$, it is not clear a priori that the resulting distribution of predicted test errors would be close to observed test errors. The covering constraint (16) only ensures that any minimizer of ${\mathcal{G}}_{0}$ should produce predictions which bound any test point, i.e. provide a max/min envelope on model errors. In practice, we will show that uniformly sampling from either the parameters or posterior predictions of the POPS-constrained loss minimizing ensemble yields very good results for a diverse set of linear regression problems, which we demonstrate in the remainder of this paper.
III.2 POPS-constrained ensemble for linear regression
Linear models use feature functions $\mathcal{F}:{\mathcal{X}}\to{\mathcal{P}}\times{\mathcal{Y}}$ to map inputs ${\mathbf{X}}\in{\mathcal{X}}$ to features ${\mathbf{F}}({\mathbf{X}})\in{\mathcal{P}}\times{\mathcal{Y}}$, giving ${\mathcal{M}}({\mathbf{X}},{\bm{\Theta}})={\bm{\Theta}}\cdot{\mathbf{F}}({\mathbf{X}})$ with parameters ${\bm{\Theta}}\in{\mathcal{P}}$. Minimizing the PAC-Bayes bound (8) with a Gaussian prior $\pi_{0}({\bm{\Theta}})=\mathcal{N}({\bm{\Theta}}|{\bf 0},{\bm{\Sigma}}_{0})$ gives a minimizer ${\bm{\Theta}}^{*}_{\mathcal{L}}$ of
${\bm{\Theta}}^{*}_{\mathcal{L}}$ $={\bf A}\langle{\mathbf{F}}({\mathbf{X}}){\mathcal{E}}({\mathbf{X}})\rangle_{\mathcal{X}},$ (20)
${\bf A}$ $=[{\bm{\Sigma}}_{\mathcal{Y}}{\bm{\Sigma}}^{-1}_{0}/N+\langle{\mathbf{F}}({\mathbf{X}})[{\mathbf{F}}({\mathbf{X}})]^{\top}\rangle]^{-1}_{\mathcal{X}}.$
We discuss how the epistemic covariance for small $N/P$ can be incorporated in a variational PAC-Bayes scheme in Appendix B. However, in the following, we concentrate on misspecification uncertainties in the limit $N/P\to\infty$, where the influence of $\bm{\Sigma}_{0}$ is negligble.
For linear models the POPS-constraint is simply ${\bm{\Theta}}_{\mathbf{X}}\cdot{\mathbf{F}}({\mathbf{X}})\equiv{\mathcal{E}}({\mathbf{X}})$. The loss minmizing pointwise fits ${\bm{\Theta}}^{*}_{{\mathbf{X}}}$ can be evaluated analytically via the efficient rank-one updates
${\bm{\Theta}}^{*}_{{\mathbf{X}}}={\bm{\Theta}}^{*}_{\mathcal{L}}+\frac{{\mathcal{E}}({\mathbf{X}})-{\bm{\Theta}}^{*}_{\mathcal{L}}\cdot{\mathbf{F}}({\mathbf{X}})}{{\rm h}({\mathbf{X}})}{\bf A}{\mathbf{F}}({\mathbf{X}}),$ (21)
where ${\rm h}({\mathbf{X}})=[{\mathbf{F}}({\mathbf{X}})]^{\top}{\bf A}{\mathbf{F}}({\mathbf{X}})$ is the leverage, an outlier measure closely related to the distance of (Mahalanobis, 1936). The minimum loss solution is the leverage-weighted centroid of pointwise optima, ${\bm{\Theta}}^{*}_{\mathcal{L}}=\langle{\rm h}({\mathbf{X}}){\bm{\Theta}}^{*}_{{\mathbf{X}}}\rangle_{\mathcal{X}}$. Equation (21) is our third main result, a scheme to incorporate misspecification uncertainty into least-squares linear regression for minimal overhead, requiring only $\mathcal{O}(N)$ inner products over feature vectors. Implementation in least squares linear regression is trivial, as shown in the following python code, where A.T is the transpose of A and we assume a uniform weighting ${\rm w}({\bf X})=1/N$:
1import numpy as np
2def pointwise_optima(X,y):
3    # X.shape = (N,P), y.shape = (N,)
4    A = np.linalg.inv(np.dot(X.T,X))
5    # Minimum loss solution
6    Theta_L = np.dot(A,np.dot(X.T,y))
7    # Minimum loss errors
8    e = y - np.dot(X, Theta_L)
9    # Leverage of each point
10    h = ( X * np.dot(X, A) ).sum(1)
11    # Pointwise fits, shape = (P,N)
12    T = np.multiply(np.dot(X,A).T, e/h )
13    return T
An ensemble of test predictions can then be made simply by contracting with a constant tensor, a process which can easily be vectorized/parallelized. We now investigate how our POPS-constrained ensemble performs in a variety of high-dimensional regression tasks.
IV Numerical Experiments
The supplementary material provides python/JaX code to reproduce all figures, with the exception of the large atomic machine learning datasets.
IV.1 Comparison with direct minimization of $\mathcal{G}$
For low dimensional problems, it is possible to approximately minimize an empirical generalization error $\mathcal{G}$ over discrete model ensembles, using a finite aleatoric uncertainty ${\bm{\Sigma}}_{\mathcal{Y}}$ for regularization. With increasing ${\bm{\Sigma}}_{\mathcal{Y}}$ it is clear that $\mathcal{G}\to\mathcal{L}$ and thus we expect the minimum loss solution $\pi_{\mathcal{L}}({\bm{\Theta}})$ to emerge. However, at small ${\bm{\Sigma}}_{\mathcal{Y}}$ direct minimization typically fails, requiring careful parameter tuning and remaining at low $P<6$ to avoid vanishing gradients or other numerical issues. The POPS-ensemble, in contrast, can easily sample $P=\mathcal{O}(1000)$ or higher dimensional data.
Figure 1 shows a simple example fitting a one dimensional quadratic polynomial to a sinusoidal function, with the constant, linear and quadratic features giving $P=3$ model parameters. However, the features are highly correlated, simplifying the regression task. At large ${\bm{\Sigma}}_{\mathcal{Y}}$ the ensemble solution agrees closely with the minimum loss solution, whilst at small ${\bm{\Sigma}}_{\mathcal{Y}}$ the range of fits broadens, with the max/min range closely matching that of the POPS-ensemble. We also plot the result of Bayesian ridge regression as implemented in scikit-learn, showing that even at the ±$\pm 3\sigma$ range, predicted parameter uncertainties fail to capture the true test errors, which, as discussed above, will worsen with increasing $N/P$.
Figure 1: Regression of a deterministic quadratic polynomial ($P=3$) model onto a sinusoidal ”simulation engine”, trained on $N=100$ points. Clockwise from top left: mean and $3\sigma$ confidence interval from Bayesian ridge regression , direct minimization of $\mathcal{G}_{E}$, regularized with $\Sigma_{\mathcal{Y}}=1$, the min/max range of the POPS-ensemble (21), and $\mathcal{G}_{E}$ with $\Sigma_{\mathcal{Y}}=0.1$.
We applied the same procedure to a model system with $P=6$ decorrelated features, unlike the polynomial example, to simulate more realistic datasets. The minimization routine could only treat a small range of $\Sigma_{\mathcal{Y}}$ values, which we normalize by the loss minimizer RMS error $\sigma_{MLE}$. Figure 2 shows the resulting parameter distributions. At large $\Sigma_{\mathcal{Y}}$ the minimizer quickly concentrates all ensemble members on the minimum loss solution, whilst at $\Sigma_{\mathcal{Y}}=\sigma_{MLE}$ the final distribution is very close to the POPS-ensemble. At lower values the final distribution was a little broader, but we attribute this largely to numerical instabilities as the final result is sensitive to initial conditions (unlike the first two cases). In addition, a minimizer initialized with the POPS-ensemble did not change appreciably. These albeit simple results give strong evidence that the POPS-ensemble is indeed a stable and highly efficient means to find approximate minimizers of ${\mathcal{G}}$ that would otherwise be intractable.
Figure 2: Regression of a $P=6$ linear model onto a cubic ”simulation engine”, trained on $N=500$ points. The panels show the culmulative distribution of each parameter across the ensemble (same size as $N$) for the loss minimizer (MLE), the POPS-ensemble (21) and from direct minimization of ${\mathcal{G}}$ with $\Sigma_{\mathcal{Y}}/\sigma_{MLE}=2,1,1/2$.
IV.2 Test errors in medium-dimensional linear regression
We now look at higher/medium dimensional problems, $P=10-100$, where direct minimization of ${\mathcal{G}}$ is not practical. The simulation engines (see supplementary material) were constrained to be misspecified from the linear surrogate models by combining features in quadratic, cubic or sinusoidal functions or randomly sampling the linear feature coefficients from some predetermined subset of models. A noise term that simulates the presence of variability orthogonal to the chosen feature set was also used. In all cases, independent test and training data were generated, typically with a 10:90 test:train split. However, all conclusions remain robust to varying the precise nature of data generation and the classes of engine considered. Additional examples are provided in the supplementary material and appendix C.
IV.2.1 POPS-ensemble prediction of test errors
For each test point, it is simple to generate an ‘envelope’, or max/min range, from the POPS-ensemble. We find this envelope provides a highly informative bound on test errors for large $N/P$. The top left of figure 3 is a parity plot for the minimum loss model, showing around 93% coverage, or 7% envelope violation. The violation rate drops with $N/P$, as shown on the bottom of figure 3 for a range of $P$ and $N/P$ and two simulation engines.
Whilst the envelope bounds are highly informative, it is also desirable to predict test error distributions. As discussed above, the limitation of our POPS-ensemble is that the weighting of ensemble members is unlikely to be optimal, whilst the max/min range can be expected to be robust. As a result, a conservative estimate of the test error can be generated by uniformly resampling the max/min bounds for each test point to generate predicted error histograms. As shown in the top (right) of figure 3, this prediction from the POPS-ensemble results in excellent agreement with the test error ensemble, as confirmed across multiple engines in appendix A. Whilst a single (re)sample per test point is sufficient, it is possible to effectively bootstrap error histograms through repeated resampling.
IV.2.2 Resampled POPS (rPOPS) ensemble
The parameter vectors of the POPS-ensemble across all simulation engines studied are observed to be dense across their simplex, or hypercube of the max/min range per parameter. Inspired by the success of the uniform resampling approach used to produce test error predictions of the POPS ensemble in figure 3, we have also investigated the effect of uniformly resampling this ‘POPS-simplex’, a process we name as ‘resampled’ POPS, or rPOPS. We emphasize that the rPOPS approach makes a strong assumption that all parameter vectors inside the POPS- simplex are close to valid POPS-ensemble members. We thus include these results as a motivation for further study; in particular, whilst the $\mathcal{O}(N)$ cost of the POPS-ensemble remains highly managable, the cost of the rPOPS-ensemble can be tailored through the number of resamples, offering large efficiencies as $N/P\to\infty$ when producing envelopes and test error distributions.
As shown in figure 3, the rPOPS results are excellent, giving tight agreement with test error distributions and only slightly more conservative errors bounds, a feature which is again robust across simulation engines (appendix C). A detailed investigation of the POPS-simplex and theoretical analysis of the rPOPS-ensemble is left as a topic for future study. We conclude with an application to challenging high-dimensional $P=\mathcal{O}(1000)$ datasets from atomic machine learning.
Figure 3: Top: test error predictions for a misspecified linear surrogate model (P=20) on a cubic simulation engine. Left: parity plot for the minimum loss model, along with envelope bounds from the POPS-ensemble and rPOPS-ensemble (see text), showing 93-98% coverage at $N/P=750$. Right: Test error histogram, along with predicted test errors from the POPS- and rPOPS-ensemble, as explained in the text. Below: POPS-ensemble envelope violation rates with $P$ and $N/P$ for two simulation engines.
IV.3 Interatomic potentials for material simulations
Figure 4: Characteristics of the ensemble method for the application to interatomic potentials. Left: Parity plot for energy and force predictions for $N/P=50$. The errorbars reflect the min/max envelope of the POPS-ensemble. Right: cumulative distribution of the ratio of the residual of the minimum-loss model to the envelope width (max/min range per test point) for different conditions.
Atomic-scale simulations of materials were traditionally broadly classified into two types:
  • Empirical simulations with physically-motivated parametric models of interatomic interactions. These allow for large simulations over long timescales at a qualitative level of accuracy.
  • First-principle quantum simulations where interatomic interactions are obtained from approximate solutions of the full Schrodinger equation. Such calculations can provide quantitative predictions at a high computational cost that strongly limits the accessible time- and length-scales.
The advent of machine learning has blurred the lines between these two categories by promising near-quantum accuracy at a much more favorable computational cost and scaling (typically cubic in the number of atoms for first-principles, and linear for empirical approaches) (Deringer et al., 2019). In the most common approaches, the total energy of the system of atoms $V$ is factored into a sum of per-atom contributions $V_{i}$, which are themselves parameterized in term of a vector of features that describe the local atomic environment around each atom ${\bf F}_{i}(\{{\bf x}_{j}|r_{ij}<r_{c}\})$, where $r_{c}$ is a interaction range. Training an ML model then proceeds by minimizing a squared error loss between the energies and forces (which are given by elements of the negative gradient of the total energy, $-\nabla V$) as obtained by first-principle methods and those predicted by the ML model. While certain ML approaches are in principle systematically improvable, the computational cost associated with evaluating energies and forces repeatedly for large systems and long times often limits the complexity of the models that are selected for applications. Further, quantum calculations can be tightly converged, so that their intrinsic error $\epsilon$ (w.r.t. a fully converged version of the same calculation) can be made small compared to typical errors incurred by the models that are selected in practice. This regime corresponds to the underparametrized deterministic limit considered above. Estimating the errors incurred by ML models of interatomic interactions (Wen and Tadmor, 2020; Li et al., 2018; Gabriel et al., 2021; Wan et al., 2021) is required when one wishes to assess the robustness of conclusions drawn from the simulations, especially when quantitative predictions are sought.
In the following, training data was generated using an information-theoretic approach (Karabin and Perez, 2020; Montes de Oca Zapiain et al., 2022) that maximizes the information entropy of the training data distribution, as measured in feature space. 10,000 atomistic configurations containing between 2 and 32 tungsten atoms were generated and characterized using Density Functional Theory (Kohn and Sham, 1965), yielding 167,922 training data points in total (10,000 energies and 157,922 force components, 3 Cartesian for each atom in every configuration).
Per-atom energies are expressed as a linear combination of $P=1596$ so-called bispectrum components within the ‘Quadratic’ Spectral Neighborhood Analysis Potential (SNAP) formalism (Wood and Thompson, 2018), giving the linear surrogate model $V_{i}={\bm{\Theta}}\cdot{\bf F}_{i}(\{{\bf x}_{j}|r_{ij}<r_{c}\})$.
Numerical experiments were carried out by randomly sub-selecting $N$ training points (either energies or gradient components), and by generating a discrete ensemble of $N$ pointwise-optimal models according to Eq. 21. The properties of the ensemble are then evaluated on a complementary set of 1,000 randomly-selected testing points (again either energies or gradient components). The random test/train partition is repeated 10 times for each value of $N/P$.
The results show that the envelope violation on test data varies between $40\%$ and $10\%$ for $N/P=1$ and $50$, respectively, in qualitative agreement with the results shown in figure 3, demonstrating that worst-case model errors are increasingly well captured by the ensemble envelope. Indeed, the scaling is actually superior to the medium test examples. As seen in these examples, the envelope induced by the model ensemble yields a useful estimate of the actual errors, e.g., as incurred by the ordinary minimum-loss model, as reported in figure 4. For example, the mean ratio of the residual of the minimum-loss model to the envelope width (max/min range) is around $0.45$ at $N/P=1$, i.e, the envelope width overestimates the actual minimum-loss residual only by a factor of around 2 on average. This figure drops to about $0.25$ at $N/P=50$, following the broadening of the envelopes that occurs as rare ”outliers” that are particularly difficult to capture within the model class are gradually added to the training set. This was empirically verified by sampling ensembles containing only $P$ models, independently of the size of the training set: in this case, the mean ratio of the residual of the minimum-loss model to the envelope width remains around $0.45$ (c.f., figure 4). This numerical experiment indicates that the width of the model envelope predictions provides reliable worst-case confidence intervals as well as representative (albeit somewhat conservative) estimates of the actual errors at testing points.
This information thus generated is invaluable to estimate the uncertainties on physical quantities inferred from simulations that encounter local atomic environments that significantly differ from those contained in the training set. We however note that accurate uncertainty quantification relies on the training sets being representative of possible testing sets, which highlights the fundamental importance training data diversity (Karabin and Perez, 2020) and the perils of narrowly tailored training sets.
V Conclusion
In spite of its ubiquity in practical applications, misspecification is often ignored in Bayesian regression approaches based on expected loss minimization, which can lead to erroneous conclusions regarding parameter uncertainty quantification, surrogate model selection, and error propagation. In this work, we show that the important near-deterministic underparameterized regime is amenable to approximation by a formally simple discrete ensemble ansatz that exploits the concentration of the cross-entropy minimizing ensemble onto pointwise optimal parameter sets. For linear regression problems, this ensemble can be obtained extremely efficiently using rank-one perturbations of the expected loss solution. POPS-constrained loss-minimizing ensembles are shown to produce excellent parameter distributions, worst-case bounds on model predictions, and approximations of the pointwise errors.
Software and Data
Will be released after peer review
Acknowledgements
This work was initiated during a visit at the Institute for Pure and Applied Mathematics (IPAM) at the University of California in Los Angeles (UCLA). Their hospitality is graciously acknowledged. TDS gratefully acknowledges support from ANR grant ANR-19-CE46-0006-1, IDRIS allocations A0120913455 and an Emergence@INP grant from the CNRS. D.P. was supported by the Laboratory Directed Research and Development program of Los Alamos National Laboratory under project number 20220063DR. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001).
Appendix A Model selection criteria for deterministic models
The Bayesian information criterion (BIC), used to discriminate between possible surrogate models, is derived by approximating (twice) the negative log evidence as $N/P\to\infty$, assuming a slowly varying prior $\pi_{0}({\bm{\Theta}})=\mathcal{O}(1)$. Using the above definitions we find
[…]
We note that the value of the PAC-Bayes loss bound (8) satifies $BIC=2N\hat{\mathcal{L}}_{N}[\pi_{{\mathcal{L}},N}]-2C_{0}$, with $C_{0}$ model independent. In standard derivations ${\bm{\Sigma}}_{\mathcal{Y}}$ is set to the variational maximum likelihood solution ${\bm{\Sigma}}_{\mathcal{Y}}={\bm{\Sigma}}^{*}_{\mathcal{L}}$, yielding the familar expression $BIC=P\ln N+N\ln\|{\bm{\Sigma}}^{*}_{\mathcal{L}}\|+\ln\|{\bm{\mathcal{I}}}^{*}_{\mathcal{L}}\|$, where model independent terms are neglected. However, for deterministic regression problems, ${\bm{\Sigma}}_{\mathcal{Y}}$ is fixed. Neglecting model-independent terms, this gives an alternative BIC for deterministic models of
$BIC=N{\rm Tr}({\bm{\Sigma}}^{*}_{\mathcal{E}}{\bm{\Sigma}}^{-1}_{\mathcal{Y}})+P\ln N+\ln\|{\bm{\mathcal{I}}}^{*}_{\mathcal{L}}\|.$ (23)
Appendix B POPS-constrained variational ansatz for PAC-Bayesian linear regression
The POPS-constrained ensemble ansatz can also be used to incorporate both epistemic and misspecification parameter uncertainties in a variational scheme robust at small $N/P$, which we detail here for completeness. We assume a Gaussian posterior $\pi_{V}({\bm{\Theta}})=\mathcal{N}({\bm{\Theta}}|{\bm{\mu}}_{V},{\bm{\Sigma}}_{V})$, such that the ensemble ansatz is characterized by the ensemble mean ${\bm{\mu}}_{E}=\langle{\bm{\Theta}}^{*}_{\mathbf{X}}\rangle$ and covariance ${\bm{\Sigma}}_{E}$. A first approximation simply combines the covariance from minimization of the PAC-Bayes loss bound (8), ${\bm{\Sigma}}^{*}_{\mathcal{L}}=\left[{\bm{\Sigma}}_{\mathcal{Y}}{\bm{\Sigma}}^{-1}_{0}+N\langle{\mathbf{F}}({\mathbf{X}})[{\mathbf{F}}({\mathbf{X}})]^{\top}\rangle_{\mathcal{X}}\right]^{-1}$, with the ensemble covariance ${\bm{\Sigma}}_{E}$. This yield a posterior covariance ${\bm{\Sigma}}^{*}_{\mathcal{L}}+{\bm{\Sigma}}_{E}$, which remains robust at finite $N/P$ and limits to ${\bm{\Sigma}}_{E}$ as $N/P\to\infty$.
However, the full variational scheme has comparable complexity to Bayesian ridge regression for linear models [Tipping, 2001]. Our variational ansatz replaces ${\bm{\Theta}}^{*}_{\mathcal{L}}$ in the pointwise fits (21) with a variational parameter ${\bm{\Theta}}_{V}$, giving a variational posterior mean and covariance of
${\bm{\mu}}_{V}({\bm{\Theta}}_{V})\equiv\left\langle{\bf v}_{\mathbf{X}}\right\rangle_{\mathcal{X}}+\left\langle{\bf W}_{\mathbf{X}}\right\rangle_{\mathcal{X}}{\bm{\Theta}}_{V},\quad{\bm{\Sigma}}_{V}({\bm{\Theta}}_{V})\equiv\left\langle\left[\delta{\bf v}_{\mathbf{X}}+\delta{\bf W}_{\mathbf{X}}{\bm{\Theta}}_{V}\right]\left[\delta{\bf v}_{\mathbf{X}}+\delta{\bf W}_{\mathbf{X}}{\bm{\Theta}}_{V}\right]^{\top}\right\rangle_{\mathcal{X}},$ (24)
where ${\bf v}_{\mathbf{X}},{\bf W}_{\mathbf{X}}$ are derived from (21) such that ${\bm{\mu}}_{V}({\bm{\Theta}}^{*}_{\mathcal{L}})={\bm{\mu}}_{E}$ and ${\bm{\Sigma}}_{V}({\bm{\Theta}}^{*}_{\mathcal{L}})={\bm{\Sigma}}_{E}$, reading
${\bf v}_{\mathbf{X}}\equiv{\bf A}\frac{{\mathbf{F}}({\mathbf{X}}){\mathcal{E}}({\mathbf{X}})}{{\rm h}({\mathbf{X}})},\quad{\bf W}_{\mathbf{X}}\equiv\mathbb{I}-{\bf A}\frac{{\mathbf{F}}({\mathbf{X}}){\mathbf{F}}^{\top}({\mathbf{X}})}{{\rm h}({\mathbf{X}})},\quad\delta{\bf v}_{\mathbf{X}}\equiv{\bf v}_{\mathbf{X}}-\left\langle{\bf v}_{\mathbf{X}}\right\rangle_{\mathcal{X}},\quad\delta{\bf W}_{\mathbf{X}}\equiv{\bf W}_{\mathbf{X}}-\left\langle{\bf W}_{\mathbf{X}}\right\rangle_{\mathcal{X}}.$
This gives a variational distribution which can be used in the PAC-Bayes loss bound ^$\hat{{\mathcal{L}}}_{N}[\pi]$ (8)
$\pi_{V}({\bm{\Theta}}|{\bm{\Theta}}_{V})\equiv\mathcal{N}({\bm{\Theta}}|{\bm{\mu}}_{V}({\bm{\Theta}}_{V}),{\bm{\Sigma}}_{V}({\bm{\Theta}}_{V})),\quad{\bm{\Theta}}_{V}^{*}\equiv\arg\min_{{\bm{\Theta}}_{V}\in{\mathcal{P}}}\hat{{\mathcal{L}}}_{N}[\pi_{V}(.|{\bm{\Theta}}_{V})].$ (25)
Using the empirical loss $\hat{L}_{N}({\bm{\Theta}})$, quadratic in ${\bm{\Theta}}$ for linear models, the variational minimum satisfies
$\left\langle{\bf W}_{\mathbf{X}}\right\rangle_{\mathcal{X}}[N{\bm{\nabla}}_{\bm{\Theta}}\hat{L}_{N}({\bm{\Theta}}^{*}_{V})+{\bm{\Sigma}}_{0}^{-1}{\bm{\mu}}_{V}({\bm{\Theta}}^{*}_{V})]+\left[N{\bf A}^{-1}+{\bm{\Sigma}}_{0}^{-1}+{\bm{\Sigma}}^{-1}_{V}({\bm{\Theta}}^{*}_{V})\right]:{\bm{\Sigma}}^{\prime}_{V}({\bm{\Theta}}^{*}_{V})={\bf 0},$ (26)
where :::: denotes a contraction over the first two indicies of a third-order tensor ${\bm{\Sigma}}^{\prime}_{V}({\bm{\Theta}})$ linear in ${\bm{\Theta}}$:
${\bm{\Sigma}}^{\prime}_{V}({\bm{\Theta}})=\left\langle\delta{\bf W}_{\mathbf{X}}\delta{\bf v}^{\top}_{\mathbf{X}}+\delta{\bf v}_{\mathbf{X}}\delta{\bf W}_{\mathbf{X}}+\delta{\bf W}_{\mathbf{X}}{\bm{\Theta}}\delta{\bf W}_{\mathbf{X}}+\delta{\bf W}_{\mathbf{X}}\delta{\bf W}_{\mathbf{X}}{\bm{\Theta}}\right\rangle_{\mathcal{X}}.$ (27)
As for Bayesian ridge regression, the only non-linearity in (26) is the term ${\bm{\Sigma}}^{-1}_{V}({\bm{\Theta}}):{\bm{\Sigma}}^{\prime}_{V}({\bm{\Theta}}^{*}_{V})$, meaning any implementation can use existing iterative schemes e.g. [Tipping, 2001],[Wipf and Nagarajan, 2007]. A full study of the low data regime is left for future work.
Appendix C Additional plots of test error prediction
The plots below are identical in presentation to the top row of figure 3, but for the quadratic and random linear simulation engines described in the main text.
Figure 5: Test error prediction for medium dimensional problems. From left: Observed test errors for a misspecified linear surrogate model (P=20) on a quadratic simulation engine, along with predicted test errors from the POPS-ensemble. Parity plots for the minimum loss model, along with envelope bounds from the POPS-ensemble.

### v2

@@
 I Introduction
-Surrogate models are widely used across science and engineering, notably to efficiently approximate the action of computationally intensive simulation engines  (Alizadeh et al., 2020; Deringer et al., 2019; Lapointe et al., 2020; Nyshadham et al., 2019; Montes de Oca Zapiain et al., 2022; Bonatti et al., 2022).
+Surrogate models are widely used across science and engineering, notably to efficiently approximate the action of computationally intensive simulation engines  (Alizadeh et al., 2020; Deringer et al., 2019; Lapointe et al., 2020; Nyshadham et al., 2019; Montes de Oca Zapiain et al., 2022; Bonatti et al., 2022; Kudela and Matousek, 2022).
 Domain expertise is leveraged to design optimal features and model architecture before parameter inference is carried out based on training data generated by the engines.
 In this paper, we consider the common scenario where the learning problems share three key characteristics:
-  • simulation engines are near deterministic: simulation outputs have vanishing aleatoric uncertainty.
-For example, the atomic energy $V({\bf X})$ is a deterministic function of atomic positions ${\bf X}$.
+  • Simulation engines are near deterministic: simulation outputs have vanishing aleatoric uncertainty, e.g. the atomic energy $V({\bf X})$ is a deterministic function of atomic positions ${\bf X}$.
   • Surrogate models are misspecified: no single choice of the $P$ model parameters can match all $N$ training observations, such that model averaging is required to capture uncertainties.
-[…]
+  • Surrogate models are underparametrized: large quantities of training data are available, such that the standard posterior from Bayesian inference gives vanishing parameter uncertainties.
+This typically occurs in the limit $P/N\ll 1$ for finite capacity models.
 Misspecification affects both finite capacity models and deep learning approaches with finite training resources  (Lahlou et al., 2021; Psaros et al., 2023).
-A misspecification-aware learning scheme should minimize the cross entropy between predicted and observed data distributions, known as the generalization error or predictive risk (see (Alquier, 2021; Morningstar et al., 2022)), but this loss is typically numerically intractable for high dimensional models.
-Bayesian learning schemes instead minimize the expected loss (log likelihood), an upper bound to the generalization error, which is known to ignore misspecification.
-Epistemic uncertainties are therefore typically severely underestimated and can be shown to vanish in the underparametrized limit, with broad implications for surrogate model selection, uncertainty quantification and error propagation, as noted by multiple groups (Masegosa, 2020; Kato et al., 2022; Lotfi et al., 2022; Lahlou et al., 2021; Psaros et al., 2023; Imbalzano et al., 2021).
+A misspecification-aware learning scheme should minimize the cross entropy between predicted and observed data distributions, known as the generalization error, or population risk (see (Alquier, 2021; Morningstar et al., 2022)).
+However, this is typically not tractable for learning, both due to numerical instabilities and the lack of theoretical bounds for convergence.
+Bayesian learning schemes instead minimize the expected loss (log likelihood), an upper bound to the generalization error, which is known to ignore misspecification but admits a robust learning scheme.
+Epistemic uncertainties are thus severely underestimated, vanishing in the underparametrized limit, with broad implications for surrogate model selection, uncertainty quantification and error propagation, as noted by multiple groups(Masegosa, 2020; Kato et al., 2022; Lotfi et al., 2022; Lahlou et al., 2021; Psaros et al., 2023; Imbalzano et al., 2021).
 We analyze the generalization error of deterministic, misspecified and under-parametrized surrogate models, i.e.
 those typically used to approximate simulation engines in science and engineering.
-The deterministic limit is atypical in Bayesian analysis, as log likelihoods are typically required to be bounded and smooth, requiring a finite aleatoric uncertainty.
-Here, we take the deterministic limit to derive conditions that any parameter distribution must obey to avoid a divergent generalization error, then derive a simple ensemble ansatz which admits a robust learning scheme.
+Our focus is the underparametrized regime, meaning we do not focus on epistemic errors important for finite $N$.
+As posterior distributions become unbounded in the deterministic limit, this is only taken as a theoretical device to derive conditions that any posterior must obey to avoid a divergent generalization error.
+We then design an ensemble ansatz that respects this condition and admits a robust learning scheme at finite $N$, which reduces to a constrained loss minimization problem for large $N/P$.
 Our main contributions are:
@@
 An ensemble ansatz is proposed that respects POPS occupancy and gives a finite generalization error.
-  • For linear models, our ansatz can be efficiently evaluated via rank-one updates to a leverage-weighted loss minimizer.
+  • For linear models, the variational minimum of our ansatz can be efficiently evaluated via rank-one updates to a leverage-weighted loss minimizer.
 This is our main result, an efficient scheme to incorporate misspecification into linear regression with minimal computational overhead.
-  • We compare our ansatz to direct numerical minimization of the generalization error, showing how our scheme provides robust bounds on test errors.
+  • We compare our variational minimum to numerial optimization of our ensemble ansatz and show how our scheme provides robust bounds on test errors.
 This is confirmed in numerical experiments on challenging datasets from atomic machine learning.
-(Germain et al., 2016) established an important connection between Bayesian inference and probability approximately correct (PAC-Bayes) analysis, which bounds true expectations by empirical likelihood expectations through concentration inequalities.
+(Germain et al., 2016) established an important connection between Bayesian inference and probability approximately correct (PAC-Bayes) analysis for regression problems with unbounded losses, which bounds true expectations by empirical likelihood expectations through concentration inequalities.
 This provides a robust learning scheme without test-train splitting, and gives a rationalization for Bayesian inference as a form of regularization.
@@
 (Lotfi et al., 2022) considered misspecification in the context of deep model selection, deriving conditional likelihoods better aligned with the generalization error.
-As misspecification has received only recent attention in the literature, various definitions for the distinction between epistemic and misspecification uncertainties exist (Kato et al., 2022).
-Here, misspecification is defined as non-aleatoric uncertainties which survive in the underparametrized limit $N/P\to\infty$, where PAC-Bayes bounds for the expected loss become tight.
+As various definitions for the distinction between epistemic and misspecification uncertainties exist (Kato et al., 2022), we define misspecification as non-aleatoric uncertainties which survive in the underparametrized limit, where the minimizer of PAC-Bayes bounds for the expected loss, which coincides with the posterior from Bayesian inference, predict vanishing parameter uncertainties.
+This definition is made more precise in section II.3.
 Our epistemic uncertainties thus coincide with standard estimates from loss minimization.
+We summarize known results for the generalization error and expected loss, PAC-Bayes bounds for the regression loss, and the sub-optimality of loss minimization for misspecified problems in section II.
+Section III contains our central theoretical contribution, an analysis of the generalization error for near-deterministic surrogate models and a simple ensemble ansatz which can be efficiently evaluated for linear models.
+Section IV applies our ansatz to high-dimensional linear regression problems, using synthetic data and challenging applications in atomic machine learning.
 II Bayesian surrogate models
@@
 A simulation engine ${\mathcal{E}}:{\mathcal{X}}\to{\mathcal{Y}}$ takes input ${\mathbf{X}}\in{\mathcal{X}}$ and returns ${\mathbf{Y}}\in{\mathcal{Y}}$.
-We denote the dimension of ${\mathcal{Y}}$ and ${\mathcal{X}}$ by $Y$ and $X$ respectively.
-We consider the regime where the aleatoric uncertainty of ${\mathcal{E}}$ is weak and can be described as a Gaussian distribution with mean $\langle{\mathbf{Y}}|{\mathbf{X}}\rangle\equiv{\mathcal{E}}({\mathbf{X}})$ and covariance $\langle{\mathbf{Y}}\otimes{\mathbf{Y}}|{\mathbf{X}}\rangle-{\mathcal{E}}({\mathbf{X}})\otimes{\mathcal{E}}({\mathbf{X}})\equiv{\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})$, where $\|{\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})\|=\mathcal{O}(\epsilon^{2Y})$, and $\epsilon$ is assumed small.
+We consider the regime where the aleatoric uncertainty of ${\mathcal{E}}$ is weak and can be described as a Gaussian distribution with mean $\langle{\mathbf{Y}}|{\mathbf{X}}\rangle\equiv{\mathcal{E}}({\mathbf{X}})$ and covariance $\langle{\mathbf{Y}}{\mathbf{Y}}^{\top}|{\mathbf{X}}\rangle-{\mathcal{E}}({\mathbf{X}}){\mathcal{E}}^{\top}({\mathbf{X}})\equiv{\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})$, where $\|{\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})\|=\mathcal{O}(\epsilon^{2Y})$, and $\epsilon$ is assumed small.
 The true conditional output distribution thus reads
@@
 We assume homoskedatiscity ${\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})={\bm{\Sigma}}_{\mathcal{Y}}$ of this weak aleatoric error to investigate the deterministic, underparametrised limit $\epsilon\to 0$, $N/P\to\infty$ for $\epsilon=\mathcal{O}(\sqrt{P/N})$.
-With weights $\sum_{i=1}^{N}{\rm w}_{i}=1$, the training configurations ${\mathcal{D}}_{N}=\{{\mathbf{X}}_{i}\}_{i=1}^{N}$ form an input distribution
+With weights $\sum_{i=1}^{N}{\rm w}_{i}=1$, the training configurations ${\mathcal{D}}_{N}=\{{\mathbf{Y}}_{i},{\mathbf{X}}_{i}\}_{i=1}^{N}$ form an input distribution in ${\mathbf{X}}$ of
 $\rho_{N}({\mathbf{X}})\equiv\sum_{i}{\rm w}_{i}\delta({\mathbf{X}}-{\mathbf{X}}_{i}),\,\,\rho_{{\mathcal{D}}}({\mathbf{X}})\equiv\lim_{N\to\infty}\rho_{N}({\mathbf{X}}).$ (2)
-The non-parametric estimate of $\rho_{N}$ will suffer from a curse of dimensionality (Tsybakov, 2008), i.e.
-the worst case error scales as $N^{1/(1+{\rm Dim}({\mathbf{X}}))}$$\simeq$1.
-However, learning only requires that we can bound expectations over independent identically distributed (i.i.d.) samples ${\mathbf{X}},{\mathbf{Y}}\in{\mathcal{D}}$
+Learning requires that we can bound expectations over independent identically distributed (i.i.d.) samples ${\mathbf{X}},{\mathbf{Y}}\in{\mathcal{D}}$
 $\langle f({\mathbf{Y}},{\mathbf{X}})\rangle_{{\mathcal{D}}}\equiv\int_{\mathcal{D}}f({\mathbf{Y}},{\mathbf{X}})\rho_{\mathcal{E}}({\mathbf{Y}}|{\mathbf{X}})\rho_{{\mathcal{D}}}({\mathbf{X}}){\rm d}{\mathbf{X}}{\rm d}{\mathbf{Y}}.$ (3)
@@
 Unlike typical Bayesian schemes, we do not treat the covariance term ${\bm{\Sigma}}_{\mathcal{Y}}$ as fitting parameter but as an intrinsic property of the quasi-deterministic simulation engine ${\mathcal{E}}$.
-This has implications for model selection criteria, as we discuss in appendix A.
-The generalization error ${\mathcal{G}}[\pi]$ is simply the cross entropy of predicted to observed distributions.
-The $\rho_{N}({\mathbf{X}})$ terms cancel inside the log and, omiting the constant ${\mathcal{G}}_{\mathcal{E}}\equiv\left\langle\ln\left|\rho_{\mathcal{E}}({\mathbf{Y}}|{\mathbf{X}})\right|\right\rangle_{\mathcal{D}}$, we have
-${\mathcal{G}}[\pi]=-\left\langle\ln\left|\int_{\mathcal{P}}\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}\right|\right\rangle_{\mathcal{D}},$ (6)
-PAC-Bayes analysis (Alquier, 2021) provides concentration inequalities (Hoeffding, 1994) that bound expectations (3) over ${\mathcal{D}}$ by empirical expectations over ${\mathcal{D}}_{N}$.
-These require $f({\mathbf{Y}},{\mathbf{X}})$ to be bounded (i.e.
-$\epsilon>0$) and expressed as an expectation over $\pi({\bm{\Theta}})$, i.e.
-$f({\mathbf{Y}},{\mathbf{X}})=\int_{\mathcal{P}}f({\mathbf{Y}},{\mathbf{X}},{\bm{\Theta}})\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}$.
-As ${\mathcal{G}}[\pi]$ is not in this form, PAC concentration inequalities can only be applied to some upper bound for ${\mathcal{G}}[\pi]$ which is an expectation over $\pi({\bm{\Theta}})$.
+This has implications for model selection criteria, as we discuss in appendix B.
+The generalization error ${\mathcal{G}}[\pi]$ is simply the cross entropy of the predicted distribution $\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}})\rho_{N}({\mathbf{X}})$ to the observed distribution $\rho({\mathbf{Y}}|{\mathbf{X}})\rho_{N}({\mathbf{X}})$; omiting the constant term ${\mathcal{G}}_{\mathcal{E}}\equiv\left\langle\ln\left|\rho_{\mathcal{E}}({\mathbf{Y}}|{\mathbf{X}})\right|\right\rangle_{\mathcal{D}}$ we have
+${\mathcal{G}}[\pi]=-\left\langle\ln\left|\int_{\mathcal{P}}\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}\right|\right\rangle_{\mathcal{D}}.$ (6)
 II.3 The expected loss and PAC estimation
-The bound used in Bayesian schemes is for the expected loss ${\mathcal{L}}[\pi]\geq{\mathcal{G}}[\pi]$, often known as the negative log likelihood.
+PAC-Bayes analysis (Alquier, 2021) provides concentration inequalities (Hoeffding, 1994) that bound expectations (3) over ${\mathcal{D}}$ by empirical expectations over ${\mathcal{D}}_{N}$ with some probability $1-\xi$.
+Whilst originally designed for classification problems where the loss is bounded to $[0,1]$, the treatment can be extended to regression settings with unbounded losses, as discussed in (Germain et al., 2016) and appendix A.
+In this setting, available PAC-Bayes bounds are for $\langle f({\mathbf{Y}},{\mathbf{X}})\rangle_{\mathcal{D}}$, where $f({\mathbf{Y}},{\mathbf{X}})$ to be bounded (i.e.
+$\epsilon>0$) and can expressed as an expectation over $\pi({\bm{\Theta}})$, i.e.
+$\langle f({\mathbf{Y}},{\mathbf{X}})\rangle_{\mathcal{D}}=\int_{\mathcal{P}}\langle f({\mathbf{Y}},{\mathbf{X}},{\bm{\Theta}})\rangle_{\mathcal{D}}\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}$.
+As ${\mathcal{G}}[\pi]$ is not in this form, the bound used in PAC-Bayesian schemes is for the expected loss ${\mathcal{L}}[\pi]\geq{\mathcal{G}}[\pi]$, often known as the negative log likelihood.
 The Jensen inequality for convex functions $f(\langle x\rangle)\leq\langle f(x)\rangle$ gives, with $f(x)=-\ln|x|$,
-$\mathcal{G}[\pi]\leq{\mathcal{L}}[\pi]=-\int_{\mathcal{P}}\left\langle\ln\left|\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\right|\right\rangle_{\mathcal{D}}\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}.$ (7)
-The empirical expected loss ^$\hat{{\mathcal{L}}}_{N}[\pi]$ over $\rho_{N}({\mathbf{X}})$ can estimate the true expected loss ${\mathcal{L}}[\pi]$ through the PAC-Bayes concentration inequality, as shown by (Germain et al., 2016)
-${\mathcal{L}}[\pi]\leq\hat{{\mathcal{L}}}_{N}[\pi]$ $\equiv\int_{\mathcal{P}}\hat{L}_{N}({\bm{\Theta}})\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}$
-$+\frac{1}{N}\int_{\mathcal{P}}\ln\left|\frac{\pi({\bm{\Theta}})}{\pi_{0}({\bm{\Theta}})}\right|\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}+\frac{C_{0}}{N}$ (8)
-where $\hat{L}_{N}({\bm{\Theta}})=-\langle\ln|\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})|\rangle_{N}$ is the empirical loss over $\rho_{N}({\mathbf{X}})$, the distribution $\pi_{0}({\bm{\Theta}})$ is to some prior and $C_{0}$ is independent of $\pi$ but depends on $\epsilon$ and $\pi_{0}$.
-A well chosen prior thus tightens the PAC bound.
-Minimization of ^$\hat{{\mathcal{L}}}_{N}[\pi]$ yields the posterior familiar from Bayesian inference, as shown by  (Germain et al., 2016), but with an additional term $\lambda$ that ensures normalization-
+$\mathcal{G}[\pi]\leq{\mathcal{L}}[\pi]=-\int_{\mathcal{P}}\left\langle\ln\left|\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\right|\right\rangle_{\mathcal{D}}\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}.$
+As discussed by (Germain et al., 2016) the expected regression loss ${\mathcal{L}}[\pi]$ admits the PAC-Bayes upper bound, holding with probability $1-\xi$, of
+${\mathcal{L}}[\pi]$$\leq\hat{{\mathcal{L}}}_{N}[\pi]+\frac{1}{N}\left(KL(\pi|\pi_{0})-\ln|\xi|\right)+C_{0}[\pi_{0}],$
+^$\hat{{\mathcal{L}}}_{N}[\pi]$ $\equiv\int_{\mathcal{P}}\hat{L}_{N}({\bm{\Theta}}){\rm d}\pi({\bm{\Theta}}),$ (7)
+where $\hat{L}_{N}({\bm{\Theta}})=-\langle\ln|\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\rangle_{N}$ is the empirical loss over $\rho_{N}({\mathbf{X}})$, $KL(\pi|\pi_{0})=\int_{\mathcal{P}}\ln|\pi/\pi_{0}|{\rm d}\pi({\bm{\Theta}})$ is the cross entropy between $\pi$ and some prior $\pi_{0}({\bm{\Theta}})$ and $C_{0}[\pi_{0}]$ is constant with respect to $\pi$ and has a finite upper bound discussed in appendix A.
+The minimizer of the PAC-Bayes upper bound (7) is that of ^$\hat{{\mathcal{L}}}_{N}[\pi]+KL(\pi|\pi_{0})/N$, giving the posterior from Bayesian inference (Germain et al., 2016)-
 $\pi_{{\mathcal{L}},N}({\bm{\Theta}})=\pi_{0}({\bm{\Theta}})\exp(\lambda-N\hat{L}_{N}({\bm{\Theta}})).$ (8)
-Appendix A discusses this bound in the context of Bayesian model selection for deterministic models.
-Using Laplace’s method as $N/P\to\infty$ gives $\hat{L}_{N}({\bm{\Theta}})\to L({\bm{\Theta}})$, with
+Appendix B discusses this bound in the context of Bayesian model selection. Using Laplace’s method (appendix C) as $N/P\to\infty$ gives $\hat{L}_{N}({\bm{\Theta}})\to L({\bm{\Theta}})$, with
 […]
@@
 $L({\bm{\Theta}})=-\langle\ln|\mathcal{N}({\mathcal{E}}({\mathbf{X}})|{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}}),{\bm{\Sigma}}_{\mathcal{Y}})|\rangle_{\mathcal{X}},$ (10)
-where $\langle\dots\rangle_{\mathcal{X}}\equiv\int\dots\rho_{\mathcal{D}}({\mathbf{X}}){\rm d}{\mathbf{X}}$.
-For linear problems, ${\bm{\Theta}}^{*}_{\mathcal{L}}$ is a generalized least squares solution, with the minimizing set a connected region in $\mathbb{R}^{P}$.
-In the underparametrized limit $N/P\to\infty$, expected loss is thus minimized by the sharp distribution $\pi_{\mathcal{L}}({\bm{\Theta}})\to\delta({\bm{\Theta}}-{\bm{\Theta}}^{*}_{\mathcal{L}})$, giving vanishing parameter uncertainty.
-The generalization error of the minimium loss solution clearly diverges as $1/\epsilon^{2}$:
-$\lim_{N/P\to\infty}{\mathcal{G}}[\pi_{\mathcal{L}}]=\frac{1}{2}{\rm Tr}({\bm{\Sigma}^{*}_{\mathcal{L}}}{\bm{\Sigma}}_{\mathcal{Y}}^{-1})=\mathcal{O}(1/\epsilon^{2}),$ (12)
-where ${\bm{\Sigma}}^{*}_{\mathcal{L}}=\langle[{\mathcal{E}}({\mathbf{X}})-{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}}^{*}_{\mathcal{L}})]\otimes[{\mathcal{E}}({\mathbf{X}})-{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}}^{*}_{\mathcal{L}})]\rangle$ is the error covariance around ${\bm{\Theta}}^{*}_{\mathcal{L}}$ and we use ${\rm Tr}({\bm{\Sigma}}_{\mathcal{Y}})=\mathcal{O}(\epsilon^{2})$.
+[…]
+$\lim_{N/P\to\infty}{\mathcal{G}}[\pi_{\mathcal{L}}]=\frac{1}{2}{\rm Tr}({\bm{\Sigma}^{*}_{\mathcal{L}}}{\bm{\Sigma}}_{\mathcal{Y}}^{-1})=\mathcal{O}(1/\epsilon^{2}).$ (11)
+${\bm{\Sigma}}^{*}_{\mathcal{L}}=\langle[{\mathcal{E}}({\mathbf{X}})-{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}}^{*}_{\mathcal{L}})][{\mathcal{E}}({\mathbf{X}})-{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}}^{*}_{\mathcal{L}})]^{\top}\rangle$ is the error covariance around ${\bm{\Theta}}^{*}_{\mathcal{L}}$ and we use ${\rm Tr}({\bm{\Sigma}}_{\mathcal{Y}})=\mathcal{O}(\epsilon^{2})$.
 As noted by (Masegosa, 2020), it is clear that the maximum likelihood parameter distribution $\pi_{\mathcal{L}}({\bm{\Theta}})$ is not a minimizer of the generalization error under misspecification.
+We emphasize that the primary motivation of this section is to show that whilst the generalization error is the central quantity of interest for learning under misspecification, we typically only have access to an upper bound in the form of a PAC-Bayes concentration inequality (7) for the expected loss.
+Under model misspecification it is clear that minimization of this upper bound in sub-optimal, in that the generalization error diverges as $1/\epsilon^{2}$.
+In the next section we analyze the generalization error as $\epsilon\to 0$ to derive a condition for $\pi({\bm{\Theta}})$ that ensures a finite generalization error.
+We then design an ensemble ansatz which respects this condition and can be parametrized through a variational minimization of the PAC-Bayes upper bound (7).
 III Analysis of the generalization error in the near-deterministic limit
@@
 The integral over ${\mathbf{Y}}$ in (6) concentrates at ${\mathbf{Y}}={\mathcal{E}}({\mathbf{X}})$; using $\langle\dots\rangle_{\mathcal{X}}\equiv\int\dots\rho_{\mathcal{D}}({\mathbf{X}}){\rm d}{\mathbf{X}}$ yields, using Laplace’s method
-${\mathcal{G}}_{0}[\pi]=-\left\langle\ln\left|\int_{\mathcal{P}}\rho_{\mathcal{M}}({\mathcal{E}}({\mathbf{X}})|{\mathbf{X}},{\bm{\Theta}})\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}\right|\right\rangle_{\mathcal{X}}.$
-Whilst ${\mathbf{Y}}={\mathcal{E}}({\mathbf{X}})$ is a unique maximum, as $\epsilon\to 0$ the marginal distribution (5) concentrates ${\bm{\Theta}}$ over pointwise optimal sets (POPS) of parameters
+${\mathcal{G}}_{0}[\pi]$ $=-\left\langle\ln\left|\rho_{\mathcal{M}}({\mathcal{E}}({\mathbf{X}})|{\mathbf{X}})\right|\right\rangle_{\mathcal{X}}$ (12)
+$=-\left\langle\ln\left|\int_{\mathcal{P}}\rho_{\mathcal{M}}({\mathcal{E}}({\mathbf{X}})|{\mathbf{X}},{\bm{\Theta}})\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}\right|\right\rangle_{\mathcal{X}}.$
+Whilst ${\mathbf{Y}}={\mathcal{E}}({\mathbf{X}})$ is a unique maximum, to ensure $\min_{{\mathbf{X}}\in{\mathcal{X}}}\rho_{\mathcal{M}}({\mathcal{E}}({\mathbf{X}})|{\mathbf{X}})>0$ and thus avoid divergence in ${\mathcal{G}}_{0}[\pi]$ as $\epsilon\to 0$ we require $\pi({\bm{\Theta}})$ has mass in every pointwise optimal parameter set (POPS)
 ${\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})\equiv\left\{{\bm{\Theta}}\,|{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}})={\mathcal{E}}({\mathbf{X}})\right\}.$ (13)
 Any model with a constant term will ensure ${\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})\neq\emptyset$, but under misspecification it clear that $\cap_{{\mathbf{X}}\in{\mathcal{X}}}{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})=\emptyset$.
-The conditional distribution $\rho_{\mathcal{M}}({\mathcal{E}}({\mathbf{X}})|{\mathbf{X}},{\bm{\Theta}})$ will sharpen as $\epsilon\to 0$, introducing a Jacobian into the integral over ${\bm{\Theta}}$.
-This can be contained in a pointwise mass function
-${\rm m}_{\mathcal{E}}({\bm{\Theta}},{\mathbf{X}})$ $\equiv\int_{\mathcal{P}}\delta({\bm{\Theta}}-{\bm{\Theta}}^{\prime})\rho_{\mathcal{M}}({\mathcal{E}}({\mathbf{X}})|{\mathbf{X}},{\bm{\Theta}}^{\prime}){\rm d}{\bm{\Theta}}^{\prime}$
-$=\int_{{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})}\frac{\delta({\bm{\Theta}}-{\bm{\Theta}}^{\prime})}{\sqrt{\|{\bm{\nabla}}_{\bm{\Theta}}{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}}^{\prime})\|}}{\rm d}{\bm{\Theta}}^{\prime}.$ (14)
-This gives our first main result, a generalization error for deterministic surrogate models
+To express the final generalization error, we introduce a pointwise mass function
+${\rm m}_{\mathcal{E}}({\bm{\Theta}},{\mathbf{X}})$ $\equiv\int_{\mathcal{P}}\delta({\bm{\Theta}}-{\bm{\Theta}}^{\prime})\rho_{\mathcal{M}}({\mathcal{E}}({\mathbf{X}})|{\mathbf{X}},{\bm{\Theta}}^{\prime}){\rm d}{\bm{\Theta}}^{\prime},$
+which gives our first main result, a generalization error for deterministic surrogate models of
 ${\mathcal{G}}_{0}[\pi]\equiv-\left\langle\ln\int_{\mathcal{P}}{\rm m}_{\mathcal{E}}({\bm{\Theta}},{\mathbf{X}})\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}\right\rangle_{\mathcal{X}}.$ (14)
-To avoid $1/\epsilon^{2}$ divergence in ${\mathcal{G}}_{0}[\pi]$ a valid $\pi({\bm{\Theta}})$ must satisfy the ‘covering’ constraint as $\epsilon\to 0$ of
+To avoid $1/\epsilon^{2}$ divergence in ${\mathcal{G}}_{0}[\pi]$ a valid $\pi({\bm{\Theta}})$ must satisfy the ‘POPS covering’ constraint as $\epsilon\to 0$ of
 $\min_{{\mathbf{X}}\in{\mathcal{X}}}\max_{{\bm{\Theta}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})}\pi({\bm{\Theta}})>\lim_{\epsilon\to 0}\exp(-c/\epsilon^{2}),$ (15)
@@
 III.1 POPS-constrained loss minimizing ensemble ansatz
-The covering constraint (16) shows, in the limit $\epsilon\rightarrow 0$, the minimizer of ${\mathcal{G}}_{0}$ concentrates onto the union of POPS for each training point.
+The covering constraint (15) shows the minimizer of ${\mathcal{G}}_{0}$ must have mass in the POPS of each training point.
 This suggests that a representation in terms of a discrete set of models is likely to capture its key features.
-Selecting some ${\bm{\Theta}}_{\mathbf{X}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})$ for each ${\mathbf{X}}\in{\mathcal{X}}$ and possibly some weighting ${\rm w}({\mathbf{X}})>0$, equation (16) is satisfied by the ensemble ansatz
-$\pi_{E}({\bm{\Theta}})\equiv\langle{\rm w}({\mathbf{X}})\delta({\bm{\Theta}}-{\bm{\Theta}}_{\mathbf{X}})\rangle_{\mathcal{X}}.$ (17)
-It is simple to show the generalization error is finite and bounded by
-${\mathcal{G}}_{0}[\pi_{E}]\leq\left\langle\ln\left|{\rm w}({\mathbf{X}}){\rm m}_{\mathcal{E}}({\bm{\Theta}}_{\mathbf{X}},{\mathbf{X}})\right|\right\rangle_{\mathcal{X}},$ (18)
-and the expected loss is ${\mathcal{L}}[\pi_{E}]=\langle{\rm w}({\mathbf{X}})L({\bm{\Theta}}_{\mathbf{X}})\rangle_{\mathcal{X}}$.
-However, whilst ${\mathcal{G}}_{0}[\pi_{E}]$ is finite, the vast majority of the pointwise fits ${\bm{\Theta}}_{\mathbf{X}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})$ will give very poor predictions away from ${\mathbf{X}}$.
-We therefore select each pointwise fit to minimize the (POPS-constrained) expected loss ${\mathcal{L}}[\pi_{E}]$, i.e.
-${\bm{\Theta}}^{*}_{\mathbf{X}}\equiv\arg\min_{{\bm{\Theta}}_{\mathbf{X}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})}L({\bm{\Theta}}_{\mathbf{X}}).$ (19)
+Selecting some set of pointwise fits $\mathcal{T}_{E}=\{{\bm{\Theta}}_{\mathbf{X}}\}_{{\mathbf{X}}\in{\mathcal{X}}}$, ${\bm{\Theta}}_{\mathbf{X}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})$ for each ${\mathbf{X}}\in{\mathcal{X}}$ and some weighting ${\rm w}_{E}({\mathbf{X}})>0$, $\langle{\rm w}_{E}({\mathbf{X}})\rangle_{\mathcal{X}}=1$, equation (15) is satisfied by the Gaussian ensemble ansatz
+$\pi_{E}({\bm{\Theta}})\equiv\langle{\rm w}_{E}({\mathbf{X}})\mathcal{N}({\bm{\Theta}}|{\bm{\Theta}}_{{\mathbf{X}}},\sigma^{2}\mathbb{I})\rangle_{\mathcal{X}},$ (16)
+where $\sigma$ is fixed and assumed small.
+Defining mass function for the ensemble which measures the number of pointwise fits ${\bm{\Theta}}_{{\mathbf{X}}^{\prime}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})$
+$m_{E}({\mathbf{X}})\equiv\left\langle{\rm w}_{E}({\mathbf{X}}^{\prime}){\rm m}_{\mathcal{E}}({\bm{\Theta}}_{{\mathbf{X}}^{\prime}},{\mathbf{X}})\right\rangle_{{\mathcal{X}}},$ (17)
+the generalization error of (16) reads, in the limit $\sigma\to 0$,
+${\mathcal{G}}_{0}[\pi_{E}]=-\left\langle\ln|m_{E}({\mathbf{X}})|\right\rangle_{{\mathcal{X}}}.$ (18)
+The above result shows to minimize ${\mathcal{G}}_{0}[\pi_{E}]$ we should select fits ${\bm{\Theta}}_{\mathbf{X}}\in\mathcal{T}_{E}$ to have maximum occupancy across all other POPS.
+In the worst case $\mathcal{T}_{E}\cap{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})={\bm{\Theta}}_{{\mathbf{X}}}$, $\forall{\mathbf{X}}\in{\mathcal{X}}$, ${\mathcal{G}}_{0}[\pi_{E}]$ still has a finite upper bound as $\sigma\to 0$ of
+${\mathcal{G}}_{0}[\pi_{E}]\leq-\left\langle\ln|{\rm w}_{E}({\mathbf{X}})|\right\rangle_{{\mathcal{X}}}+\mathcal{S}_{{\mathcal{X}}},$ (19)
+where $\mathcal{S}_{{\mathcal{X}}}=-\left\langle\ln|\rho({\mathbf{X}})|\right\rangle_{{\mathcal{X}}}$ is the entropy of the input distribution.
+In appendix D we show how each ${\bm{\Theta}}_{\mathbf{X}}\in\mathcal{T}_{E}$ can be determined in a variational setting at finite $\sigma$, using (16) in the PAC-Bayes loss bound (7) for the regression loss ${\mathcal{L}}[\pi_{E}]=\langle{\rm w}_{E}({\mathbf{X}})L({\bm{\Theta}}_{\mathbf{X}})\rangle_{\mathcal{X}}$, which as $P/N\to 0$, $\sigma\to 0$ reduces to POPS-constrained loss minimization
+${\bm{\Theta}}^{*}_{\mathbf{X}}$ $\equiv\arg\min_{{\bm{\Theta}}_{\mathbf{X}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})}L({\bm{\Theta}}_{\mathbf{X}}),$ (20)
+$\pi^{*}_{E}({\bm{\Theta}})$ $\equiv\langle{\rm w}_{E}({\mathbf{X}})\delta({\bm{\Theta}}-{\bm{\Theta}}^{*}_{{\mathbf{X}}})\rangle_{\mathcal{X}}.$ (21)
 The goal of this POPS-constrained loss-minimization is to ensure the pointwise fits will be as close as possible; for convex loss landscapes, this is measured by the Fisher information metric ${\bm{\mathcal{I}}}^{*}_{\mathcal{L}}$.
-By minimizing the ensemble spread, we aim to distribute an (albeit small) mass to other POPS, thereby reducing the generalization error.
-In addition, for specified regression problems $\pi_{E}({\bm{\Theta}})$ then reduces to the true minimizer $\pi^{*}_{\mathcal{L}}({\bm{\Theta}})=\delta({\bm{\Theta}}-{\bm{\Theta}}^{*}_{\mathcal{L}})$.
+By minimizing the ensemble spread, we aim to distribute an (albeit small) mass to other POPS, thereby reducing ${\mathcal{G}}_{0}$.
+In addition, for specified regression problems $\pi^{*}_{E}({\bm{\Theta}})$ then reduces to the true minimizer $\pi^{*}_{\mathcal{L}}({\bm{\Theta}})=\delta({\bm{\Theta}}-{\bm{\Theta}}^{*}_{\mathcal{L}})$.
 When the loss landscape is not convex additional constraints will have to be imposed to minimize (and measure) the ensemble spread.
-More generally, POPS-constrained loss minimization with our ensemble ansatz is potentially sub-optimal in that some other ${\bm{\Theta}}_{\mathbf{X}}$, which are not loss minimizers, could be members of multiple POPS, i.e.
-${\bm{\Theta}}_{\mathbf{X}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}}^{\prime})$, which is a priori beneficial to minimizing the generalization error.
-The loss minimization scheme (19) will minimize the average prediction error of each selected ${\bm{\Theta}}_{\mathbf{X}}$ across all other ${\mathbf{X}}^{\prime}\in{\mathcal{X}}$.
-Exploring this tradeoff (multiple POPS occupancy vs global prediction error) would entail a combinatorial scheme which our simple ensemble ansatz avoids.
-A key limitation of (19) is that many pointwise fits may be near-degenerate, raising the generalization error and, in general, underestimating the average deviation of any test point from the minimum loss prediction.
+More generally, POPS-constrained loss minimization with our ensemble ansatz is potentially sub-optimal in that $\pi^{*}_{E}({\bm{\Theta}})$ chooses ${\bm{\Theta}}^{*}_{\mathbf{X}}$ to minimize the average prediction error across all other ${\mathbf{X}}^{\prime}\in{\mathcal{X}}$ but is not guaranteed to be a minimizer of (18).
+Exploring this tradeoff (multiple POPS occupancy vs global prediction error) would entail a combinatorial scheme which (21) avoids.
+One strategy, to be investigated in future work, would be to cluster the loss-minimizing pointwise fits, try to find POPS-intersections for each cluster, then replace all members of that cluster with a single POPS-intersecting fit.
+A further limitation of (21) is that many pointwise fits may be near-degenerate, raising the generalization error and, in general, underestimating the average deviation of any test point from the minimum loss prediction.
 However, the predicted max/min range of any test point is expected to be robust in the limit $N/P\to\infty$, as this is invariant to how we weight the ensemble.
@@
 ${\bf A}$ $=[{\bm{\Sigma}}_{\mathcal{Y}}{\bm{\Sigma}}^{-1}_{0}/N+\langle{\mathbf{F}}({\mathbf{X}})[{\mathbf{F}}({\mathbf{X}})]^{\top}\rangle]^{-1}_{\mathcal{X}}.$
-We discuss how the epistemic covariance for small $N/P$ can be incorporated in a variational PAC-Bayes scheme in Appendix B.
-However, in the following, we concentrate on misspecification uncertainties in the limit $N/P\to\infty$, where the influence of $\bm{\Sigma}_{0}$ is negligble.
+We discuss how the epistemic covariance for small $N/P$ can be incorporated in a variational PAC-Bayes scheme in Appendix D.
+However, in the following, we concentrate on misspecification uncertainties in the limit $N/P\to\infty$, where the influence of $\bm{\Sigma}_{0}$ is negligible.
 For linear models the POPS-constraint is simply ${\bm{\Theta}}_{\mathbf{X}}\cdot{\mathbf{F}}({\mathbf{X}})\equiv{\mathcal{E}}({\mathbf{X}})$.
-The loss minmizing pointwise fits ${\bm{\Theta}}^{*}_{{\mathbf{X}}}$ can be evaluated analytically via the efficient rank-one updates
+The loss minimizing pointwise fits ${\bm{\Theta}}^{*}_{{\mathbf{X}}}$ can be evaluated analytically via the efficient rank-one updates
 ${\bm{\Theta}}^{*}_{{\mathbf{X}}}={\bm{\Theta}}^{*}_{\mathcal{L}}+\frac{{\mathcal{E}}({\mathbf{X}})-{\bm{\Theta}}^{*}_{\mathcal{L}}\cdot{\mathbf{F}}({\mathbf{X}})}{{\rm h}({\mathbf{X}})}{\bf A}{\mathbf{F}}({\mathbf{X}}),$ (23)
@@
 Equation (23) is our third main result, a scheme to incorporate misspecification uncertainty into least-squares linear regression for minimal overhead, requiring only $\mathcal{O}(N)$ inner products over feature vectors.
-Implementation in least squares linear regression is trivial, as shown in the following python code, where A.T is the transpose of A and we assume a uniform weighting ${\rm w}({\bf X})=1/N$:
-1import numpy as np
-2def pointwise_optima(X,y):
-3    # X.shape = (N,P), y.shape = (N,)
-4    A = np.linalg.inv(np.dot(X.T,X))
-5    # Minimum loss solution
-6    Theta_L = np.dot(A,np.dot(X.T,y))
-7    # Minimum loss errors
-8    e = y - np.dot(X, Theta_L)
-9    # Leverage of each point
-10    h = ( X * np.dot(X, A) ).sum(1)
-11    # Pointwise fits, shape = (P,N)
-12    T = np.multiply(np.dot(X,A).T, e/h )
-13    return T
-An ensemble of test predictions can then be made simply by contracting with a constant tensor, a process which can easily be vectorized/parallelized.
+Implementation in least squares linear regression is trivial, as shown in appendix F.
 We now investigate how our POPS-constrained ensemble performs in a variety of high-dimensional regression tasks.
@@
 With increasing ${\bm{\Sigma}}_{\mathcal{Y}}$ it is clear that $\mathcal{G}\to\mathcal{L}$ and thus we expect the minimum loss solution $\pi_{\mathcal{L}}({\bm{\Theta}})$ to emerge.
-However, at small ${\bm{\Sigma}}_{\mathcal{Y}}$ direct minimization typically fails, requiring careful parameter tuning and remaining at low $P<6$ to avoid vanishing gradients or other numerical issues.
+We also compare these appraoches to Bayesian ridge regression, corresponding to minimizing a first-order PAC-Bayes loss bound, and an ensemble approach from citepmasegosa2020learning, section C.3, which corresponds to minimizing a second-order PAC-Bayes loss bound, whose loss function, as for $\mathcal{G}_{E}$, also requires exponentiating model errors.
+At small ${\bm{\Sigma}}_{\mathcal{Y}}$ direct minimization of $\mathcal{G}_{E}$ and the Masegosa ensemble bound typically fails, requiring careful parameter tuning and remaining at low $P<6$ to avoid vanishing gradients or other numerical issues.
 The POPS-ensemble, in contrast, can easily sample $P=\mathcal{O}(1000)$ or higher dimensional data.
@@
 However, the features are highly correlated, simplifying the regression task.
-At large ${\bm{\Sigma}}_{\mathcal{Y}}$ the ensemble solution agrees closely with the minimum loss solution, whilst at small ${\bm{\Sigma}}_{\mathcal{Y}}$ the range of fits broadens, with the max/min range closely matching that of the POPS-ensemble.
-We also plot the result of Bayesian ridge regression as implemented in scikit-learn, showing that even at the ±$\pm 3\sigma$ range, predicted parameter uncertainties fail to capture the true test errors, which, as discussed above, will worsen with increasing $N/P$.
+We show results for ${\bm{\Sigma}}_{\mathcal{Y}}=1/3$, at the lower limit for which minimization converges for all methods.
+As expected from section II.3 Bayesian ridge regression significantly underestimates test errors even at the ±$\pm 3\sigma$ level, which as discussed above will worsen with increasing $N/P$.
+We find similar behaviour for the Masegosa ensemble, which we note is primarily designed for cases with finite aleatoric uncertainty.
+Direct minimization of $\mathcal{G}_{E}$ gives better bounding of test errors, though still with notable underestimates in high-density regions, though bounnds became tighter with lower ${\bm{\Sigma}}_{\mathcal{Y}}$ (whilst stable) However, the POPS-constrained ensemble gives almost perfect bounding of test errors, being stable for arbitrarly small ${\bm{\Sigma}}_{\mathcal{Y}}$ values.
 Figure 1: Regression of a deterministic quadratic polynomial ($P=3$) model onto a sinusoidal ”simulation engine”, trained on $N=100$ points.
@@
 However, all conclusions remain robust to varying the precise nature of data generation and the classes of engine considered.
-Additional examples are provided in the supplementary material and appendix C.
+Additional examples are provided in the supplementary material and appendix G.
 IV.2.1 POPS-ensemble prediction of test errors
@@
 As a result, a conservative estimate of the test error can be generated by uniformly resampling the max/min bounds for each test point to generate predicted error histograms.
-As shown in the top (right) of figure 3, this prediction from the POPS-ensemble results in excellent agreement with the test error ensemble, as confirmed across multiple engines in appendix A.
+As shown in the top (right) of figure 3, this prediction from the POPS-ensemble results in excellent agreement with the test error ensemble, as confirmed across multiple engines in appendix B.
 Whilst a single (re)sample per test point is sufficient, it is possible to effectively bootstrap error histograms through repeated resampling.
@@
 We thus include these results as a motivation for further study; in particular, whilst the $\mathcal{O}(N)$ cost of the POPS-ensemble remains highly managable, the cost of the rPOPS-ensemble can be tailored through the number of resamples, offering large efficiencies as $N/P\to\infty$ when producing envelopes and test error distributions.
-As shown in figure 3, the rPOPS results are excellent, giving tight agreement with test error distributions and only slightly more conservative errors bounds, a feature which is again robust across simulation engines (appendix C).
+As shown in figure 3, the rPOPS results are excellent, giving tight agreement with test error distributions and only slightly more conservative errors bounds, a feature which is again robust across simulation engines (appendix G).
 A detailed investigation of the POPS-simplex and theoretical analysis of the rPOPS-ensemble is left as a topic for future study.
@@
 Atomic-scale simulations of materials were traditionally broadly classified into two types:
-  • Empirical simulations with physically-motivated parametric models of interatomic interactions. These allow for large simulations over long timescales at a qualitative level of accuracy.
-  • First-principle quantum simulations where interatomic interactions are obtained from approximate solutions of the full Schrodinger equation.
-Such calculations can provide quantitative predictions at a high computational cost that strongly limits the accessible time- and length-scales.
-The advent of machine learning has blurred the lines between these two categories by promising near-quantum accuracy at a much more favorable computational cost and scaling (typically cubic in the number of atoms for first-principles, and linear for empirical approaches) (Deringer et al., 2019).
-In the most common approaches, the total energy of the system of atoms $V$ is factored into a sum of per-atom contributions $V_{i}$, which are themselves parameterized in term of a vector of features that describe the local atomic environment around each atom ${\bf F}_{i}(\{{\bf x}_{j}|r_{ij}<r_{c}\})$, where $r_{c}$ is a interaction range.
-Training an ML model then proceeds by minimizing a squared error loss between the energies and forces (which are given by elements of the negative gradient of the total energy, $-\nabla V$) as obtained by first-principle methods and those predicted by the ML model.
-While certain ML approaches are in principle systematically improvable, the computational cost associated with evaluating energies and forces repeatedly for large systems and long times often limits the complexity of the models that are selected for applications.
+  • Empirical simulations with physically-motivated parametric models of interatomic interactions. These allow for fast simulations at a qualitative level of accuracy.
+  • First-principle quantum simulations obtained from approximate solutions of the full Schrodinger equation. These incur a high computational cost and exhibit poor scalability with system size.
+The advent of ML has blurred the lines between these two limits by promising near-quantum accuracy at a much more favorable computational cost and scaling (Deringer et al., 2019).
+The total energy of the system $V$ is typically factored into a sum of per-atom contributions $V_{i}$, which are themselves parameterized in term of a feature vector that describes the local atomic environment ${\bf F}_{i}(\{{\bf x}_{j}|r_{ij}<r_{c}\})$, where $r_{c}$ is a interaction range.
+Training an ML model then proceeds by minimizing a squared error loss between reference energies and gradients.
+While certain ML approaches are systematically improvable, the computational cost associated with repeated evaluations can limit the complexity of the models that are selected for applications.
 Further, quantum calculations can be tightly converged, so that their intrinsic error $\epsilon$ (w.r.t.
-a fully converged version of the same calculation) can be made small compared to typical errors incurred by the models that are selected in practice.
-This regime corresponds to the underparametrized deterministic limit considered above.
-Estimating the errors incurred by ML models of interatomic interactions (Wen and Tadmor, 2020; Li et al., 2018; Gabriel et al., 2021; Wan et al., 2021) is required when one wishes to assess the robustness of conclusions drawn from the simulations, especially when quantitative predictions are sought.
-In the following, training data was generated using an information-theoretic approach (Karabin and Perez, 2020; Montes de Oca Zapiain et al., 2022) that maximizes the information entropy of the training data distribution, as measured in feature space.
-10,000 atomistic configurations containing between 2 and 32 tungsten atoms were generated and characterized using Density Functional Theory (Kohn and Sham, 1965), yielding 167,922 training data points in total (10,000 energies and 157,922 force components, 3 Cartesian for each atom in every configuration).
-Per-atom energies are expressed as a linear combination of $P=1596$ so-called bispectrum components within the ‘Quadratic’ Spectral Neighborhood Analysis Potential (SNAP) formalism (Wood and Thompson, 2018), giving the linear surrogate model $V_{i}={\bm{\Theta}}\cdot{\bf F}_{i}(\{{\bf x}_{j}|r_{ij}<r_{c}\})$.
+a fully converged calculation) can be made small compared to typical errors incurred by computationally-efficient models.
+This regime corresponds to the underparametrized deterministic setting considered here.
+Estimating the errors incurred by such ML models is highly sought after to assess the robustness of conclusions drawn from the simulations (Wen and Tadmor, 2020; Li et al., 2018; Gabriel et al., 2021; Wan et al., 2021) .
+In the following, training data was generated using an information-theoretic approach (Karabin and Perez, 2020; Montes de Oca Zapiain et al., 2022).
+10,000 atomistic configurations containing between 2 and 32 tungsten atoms were generated and characterized using Density Functional Theory (Kohn and Sham, 1965), yielding 167,922 training data points in total (10,000 energies and 157,922 gradient components).
+Per-atom energies are expressed as a linear combination of $P=1596$ so-called bispectrum components within the ‘Quadratic’ Spectral Neighborhood Analysis Potential (SNAP) formalism (Wood and Thompson, 2018), which can be expressed as a linear surrogate model $V_{i}={\bm{\Theta}}\cdot{\bf F}_{i}(\{{\bf x}_{j}|r_{ij}<r_{c}\})$.
 Numerical experiments were carried out by randomly sub-selecting $N$ training points (either energies or gradient components), and by generating a discrete ensemble of $N$ pointwise-optimal models according to Eq.
-21.
-The properties of the ensemble are then evaluated on a complementary set of 1,000 randomly-selected testing points (again either energies or gradient components).
+23.
+The properties of the ensemble are then evaluated on a complementary set of 1,000 testing points.
 The random test/train partition is repeated 10 times for each value of $N/P$.
@@
 This numerical experiment indicates that the width of the model envelope predictions provides reliable worst-case confidence intervals as well as representative (albeit somewhat conservative) estimates of the actual errors at testing points.
-This information thus generated is invaluable to estimate the uncertainties on physical quantities inferred from simulations that encounter local atomic environments that significantly differ from those contained in the training set.
-We however note that accurate uncertainty quantification relies on the training sets being representative of possible testing sets, which highlights the fundamental importance training data diversity (Karabin and Perez, 2020) and the perils of narrowly tailored training sets.
 V Conclusion
-In spite of its ubiquity in practical applications, misspecification is often ignored in Bayesian regression approaches based on expected loss minimization, which can lead to erroneous conclusions regarding parameter uncertainty quantification, surrogate model selection, and error propagation.
+stIn spite of its ubiquity in practical applications, misspecification is often ignored in Bayesian regression approaches based on expected loss minimization, which can lead to erroneous conclusions regarding parameter uncertainty quantification, surrogate model selection, and error propagation.
 In this work, we show that the important near-deterministic underparameterized regime is amenable to approximation by a formally simple discrete ensemble ansatz that exploits the concentration of the cross-entropy minimizing ensemble onto pointwise optimal parameter sets.
@@
 89233218CNA000001).
-Appendix A Model selection criteria for deterministic models
+Appendix A PAC-Bayes concentration inequalities for regression
+PAC-Bayes inequalities were originally designed for classification problems where the loss function is bounded to $[0,1]$.
+The extension to regression requires treatment of unbounded losses, which is discussed in detail in (Germain et al., 2016).
+Determining the value of the upper bound for $C_{0}[\pi_{0}]$ in the PAC-Bayes concentration inequality (7) requires certain assumptions on the distribution of the error between the empirical and true expected loss (Germain et al., 2016).
+Following (Germain et al., 2016), we assume that this error distribution is unbounded but sub-normal, i.e.
+all error fluctuations are bounded by a normal distribution of variance $s^{2}$, giving $C_{0}[\pi_{0}]<s^{2}/2$.
+We refer the reader to (Germain et al., 2016) and subsequent works (Shalaeva et al., 2020; Masegosa, 2020) for further discussion.
+Appendix B Model selection criteria for deterministic models
 The Bayesian information criterion (BIC), used to discriminate between possible surrogate models, is derived by approximating (twice) the negative log evidence as $N/P\to\infty$, assuming a slowly varying prior $\pi_{0}({\bm{\Theta}})=\mathcal{O}(1)$.
@@
 […]
-We note that the value of the PAC-Bayes loss bound (8) satifies $BIC=2N\hat{\mathcal{L}}_{N}[\pi_{{\mathcal{L}},N}]-2C_{0}$, with $C_{0}$ model independent.
+We note that the value of the PAC-Bayes loss bound (7) satifies $BIC=2N\hat{\mathcal{L}}_{N}[\pi_{{\mathcal{L}},N}]-2C_{0}$, with $C_{0}$ model independent, as discussed in (Germain et al., 2016).
 In standard derivations ${\bm{\Sigma}}_{\mathcal{Y}}$ is set to the variational maximum likelihood solution ${\bm{\Sigma}}_{\mathcal{Y}}={\bm{\Sigma}}^{*}_{\mathcal{L}}$, yielding the familar expression $BIC=P\ln N+N\ln\|{\bm{\Sigma}}^{*}_{\mathcal{L}}\|+\ln\|{\bm{\mathcal{I}}}^{*}_{\mathcal{L}}\|$, where model independent terms are neglected.
@@
 $BIC=N{\rm Tr}({\bm{\Sigma}}^{*}_{\mathcal{E}}{\bm{\Sigma}}^{-1}_{\mathcal{Y}})+P\ln N+\ln\|{\bm{\mathcal{I}}}^{*}_{\mathcal{L}}\|.$ (25)
-Appendix B POPS-constrained variational ansatz for PAC-Bayesian linear regression
+Appendix C Summary of Laplace’s method
+Laplace’s method, also known as the steepest descents method, is a well-known identity allowing the evaluation of an integral of an exponentiated function multiplied by a large number.
+The method applies to a function $f(x)$ which is twice differentiable and has a unique maximum in some closed interval $[a,b]$ which may be the entire real line.
+This encompasses all applications in this paper; the extension to multidimensional functions is straightforward.
+Laplace’s method then reads
+$\lim_{N\to\infty}\int_{a}^{b}\exp[Nf(x)]{\rm d}{x}=\sqrt{\frac{2\pi}{N}}\sqrt{\frac{{\rm d}^{2}f((x)}{{\rm d}x^{2}}}^{-1}\exp[N\max_{x\in[a,b]}f(x)].$ (26)
+Appendix D POPS-constrained variational minimization of ensemble ansatz
+For finite $N/P$, we have access to training data ${\mathcal{D}}_{N}=\{{\mathbf{Y}}_{i},{\mathbf{X}}_{i}\}_{i=1}^{N}$ to give averages over $\langle\dots\rangle_{N}$, with some set of pointwise fits $\mathcal{T}_{E}=\{{\bm{\Theta}}_{i}\}_{i=1}^{N}$, ${\bm{\Theta}}_{i}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}}_{i})$.
+Our ensemble ansatz then reads
+$\pi_{E}({\bm{\Theta}})=\langle{\rm w}_{E,i}\mathcal{N}({\bm{\Theta}}|{\bm{\Theta}}_{i},\sigma^{2}\mathbb{I})\rangle_{N}=\sum_{{\mathbf{X}}_{i}\in{\mathcal{D}}_{N}}{\rm w}_{E,i}{\rm w}_{i}\mathcal{N}({\bm{\Theta}}|{\bm{\Theta}}_{i},\sigma^{2}\mathbb{I}),$ (27)
+where ${\rm w}_{E,i}$ are the discrete ensemble weights and ${\rm w}_{i}$ are the training data weights.
+The empirical loss ^$\hat{\mathcal{L}}_{N}[\pi_{E}]$, which contributes All $\pi$-dependent terms in the PAC-Bayes upper bound (7) then read
+^$\hat{\mathcal{L}}_{N}[\pi_{E}]=\int\hat{L}_{N}({\bm{\Theta}})\pi_{E}({\bm{\Theta}}),\quad KL(\pi_{E}|\pi_{0})=\int\ln\left|\frac{\pi_{E}({\bm{\Theta}})}{\pi_{0}({\bm{\Theta}})}\right|\pi_{E}({\bm{\Theta}}){\rm d}{\bm{\Theta}}.$ (28)
+To proceed, we define a POPS-constrained variation $\delta_{i}{\bm{\Theta}}_{i}$ such that ${\bm{\Theta}}_{i}+\delta_{i}{\bm{\Theta}}_{i}\in\mathcal{P}_{\mathcal{M}}({\mathbf{X}}_{i})$, which gives a variation in $\pi_{E}$ of
+$\delta_{i}\pi_{E}({\bm{\Theta}})\equiv\frac{{\rm w}_{E,i}{\rm w}_{i}}{\sigma^{2}}\delta{\bm{\Theta}}_{i}^{\top}\left[{\bm{\Theta}}-{\bm{\Theta}}_{i}\right]\mathcal{N}({\bm{\Theta}}|{\bm{\Theta}}_{i},\sigma^{2}\mathbb{I}).$ (29)
+The upper bound (7) thus has a POPS-constrained variation
+$\delta_{i}\left(\hat{\mathcal{L}}_{N}[\pi_{E}]+\frac{1}{N}KL(\pi_{E}|\pi_{0})\right)$ $=\frac{{\rm w}_{E,i}{\rm w}_{i}}{\sigma^{2}}\int\delta{\bm{\Theta}}_{i}^{\top}\left[{\bm{\Theta}}-{\bm{\Theta}}_{i}\right]\left(\hat{L}_{N}({\bm{\Theta}})+\frac{1}{N}\ln\left|\frac{\pi_{E}({\bm{\Theta}})}{\pi_{0}({\bm{\Theta}})}\right|\right)\mathcal{N}({\bm{\Theta}}|{\bm{\Theta}}_{i},\sigma^{2}\mathbb{I}){\rm d}{\bm{\Theta}},$
+$={\rm w}_{E_{i}}{\rm w}_{i}\delta_{i}{\bm{\Theta}}_{i}^{\top}\int\nabla_{\bm{\Theta}}\left(\hat{L}_{N}({\bm{\Theta}})+\frac{1}{N}\ln\left|\frac{\pi_{E}({\bm{\Theta}})}{\pi_{0}({\bm{\Theta}})}\right|\right)\mathcal{N}({\bm{\Theta}}|{\bm{\Theta}}_{i},\sigma^{2}\mathbb{I}){\rm d}{\bm{\Theta}},$ (30)
+where we use integration by parts.
+Aside from the POPS constrained variation, this is a standard variational result which in principle allows for variational minimization of $\pi_{E}({\bm{\Theta}})$ at finite $N$ for both linear and non-linear models.
+An alternative POPS-constrained variational approach for linear regression at finite $N$ is provided in appendix E.
+In the limit $\sigma\to 0$ we make the approximation that, for ${\bm{\Theta}}-{\bm{\Theta}}_{i}=\delta{\bm{\Theta}}$, $\pi_{E}({\bm{\Theta}}_{i}+\delta{\bm{\Theta}})\simeq(1/N)\mathcal{N}(\delta{\bm{\Theta}}|{\bf 0},\sigma^{2}\mathbb{I})$.
+With a Gaussian prior of zero mean and covariance ${\bm{\Sigma}}_{0}$ we have
+$\int\nabla_{\bm{\Theta}}\ln\left|\frac{\pi_{E}({\bm{\Theta}})}{\pi_{0}({\bm{\Theta}})}\right|\mathcal{N}({\bm{\Theta}}|{\bm{\Theta}}_{i},\sigma^{2}\mathbb{I}){\rm d}{\bm{\Theta}}=\int\left(\frac{{\bm{\Theta}}-{\bm{\Theta}}_{i}}{\sigma^{2}}+{\bm{\Sigma}}_{0}^{-1}{\bm{\Theta}}\right)\mathcal{N}({\bm{\Theta}}|{\bm{\Theta}}_{i},\sigma^{2}\mathbb{I}){\rm d}{\bm{\Theta}}\to{\bm{\Sigma}}_{0}^{-1}{\bm{\Theta}}_{i},\quad\sigma\to 0.$ (31)
+In the present work we are primarily concerned with the underparametrized, or large data limit $P/N\to 0$, where the influence of the prior distribution is negligible, as demonstrated in section II.3.
+In this regime, for small $\sigma$ the variation of the upper bound (7) has a limiting form of
+$\delta_{i}\hat{\mathcal{L}}_{N}[\pi_{E}]={\rm w}_{E_{i}}{\rm w}_{i}\delta_{i}{\bm{\Theta}}_{i}^{\top}\left[\nabla_{\bm{\Theta}}\hat{L}_{N}({\bm{\Theta}}_{i})+\frac{1}{N}{\bm{\Sigma}}_{0}^{-1}{\bm{\Theta}}_{i}\right]+\mathcal{O}(\sigma^{2}).$ (32)
+This result shows the POPS-constrained variation of the PAC-Bayes loss upper bound converges to the gradient of the squared loss evaluated at a POPS-constrained parameter choice, to within a positive multiplicative constant .
+As a result, in the underparametrized limit the variational optimal choice of of POPS-constrained parameter ${\bm{\Theta}}_{\mathbf{X}}$ will be given by
+${\bm{\Theta}}^{*}_{i}\equiv\arg\min_{{\bm{\Theta}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}}_{i})}L({\bm{\Theta}}),$ (33)
+in agreement with the result given in equation (21).
+Appendix E POPS-constrained variational ansatz for PAC-Bayesian linear regression
 The POPS-constrained ensemble ansatz can also be used to incorporate both epistemic and misspecification parameter uncertainties in a variational scheme robust at small $N/P$, which we detail here for completeness.
@@
 This yield a posterior covariance ${\bm{\Sigma}}^{*}_{\mathcal{L}}+{\bm{\Sigma}}_{E}$, which remains robust at finite $N/P$ and limits to ${\bm{\Sigma}}_{E}$ as $N/P\to\infty$.
-However, the full variational scheme has comparable complexity to Bayesian ridge regression for linear models [Tipping, 2001].
+However, the full variational scheme has comparable complexity to Bayesian ridge regression for linear models (Tipping, 2001).
 Our variational ansatz replaces ${\bm{\Theta}}^{*}_{\mathcal{L}}$ in the pointwise fits (23) with a variational parameter ${\bm{\Theta}}_{V}$, giving a variational posterior mean and covariance of
@@
 As for Bayesian ridge regression, the only non-linearity in (36) is the term ${\bm{\Sigma}}^{-1}_{V}({\bm{\Theta}}):{\bm{\Sigma}}^{\prime}_{V}({\bm{\Theta}}^{*}_{V})$, meaning any implementation can use existing iterative schemes e.g.
-[Tipping, 2001],[Wipf and Nagarajan, 2007].
+(Tipping, 2001),(Wipf and Nagarajan, 2007).
 A full study of the low data regime is left for future work.
-Appendix C Additional plots of test error prediction
+Appendix F Implementation of the ensemble ansatz in Python
+The following python code, where A.T is the transpose of A and we assume a uniform weighting ${\rm w}_{E}({\bf X})=1$:
+1import numpy as np
+2def pointwise_optima(X,y):
+3    # X.shape = (N,P), y.shape = (N,)
+4    # Inverse of A
+5    A_i = np.dot(X.T,X) +
+6    # Minimum loss solution
+7    T_L = np.linalg.solve(A_i,np.dot(X.T,y))
+8    # Minimum loss errors
+9    e = y - np.dot(X, T_L)
+10    # Leverage of each point
+11    X_dot_A = np.linalg.solve(A_i,X.T)
+12    h = ( X * X_dot_A ).sum(1)
+13    # Pointwise fits, shape = (P,N)
+14    T = np.multiply(X_dot_A, e/h )
+15    return T
+An ensemble of test predictions can then be made simply by contracting with a constant tensor, a process which can easily be vectorized/parallelized.
+Appendix G Additional plots of test error prediction
 The plots below are identical in presentation to the top row of figure 3, but for the quadratic and random linear simulation engines described in the main text.

### v3

@@
-Misspecification uncertainties in near-deterministic regression
+Parameter uncertainties for imperfect surrogate models in the low-noise regime
 Abstract
-The expected loss is an upper bound to the model generalization error which admits robust PAC-Bayes bounds for learning.
-However, loss minimization is known to ignore misspecification, where models cannot exactly reproduce observations.
-This leads to significant underestimates of parameter uncertainties in the large data, or underparameterized, limit.
-We analyze the generalization error of near-deterministic, misspecified and underparametrized surrogate models, a regime of broad relevance in science and engineering.
-We show posterior distributions must cover every training point to avoid a divergent generalization error and derive an ensemble ansatz that respects this constraint, which for linear models incurs minimal overhead.
-The efficient approach is demonstrated on model problems before application to high dimensional datasets in atomistic machine learning.
-Parameter uncertainties from misspecification survive in the underparametrized limit, giving accurate prediction and bounding of test errors.
-I Introduction
-Surrogate models are widely used across science and engineering, notably to efficiently approximate the action of computationally intensive simulation engines  (Alizadeh et al., 2020; Deringer et al., 2019; Lapointe et al., 2020; Nyshadham et al., 2019; Montes de Oca Zapiain et al., 2022; Bonatti et al., 2022; Kudela and Matousek, 2022).
-Domain expertise is leveraged to design optimal features and model architecture before parameter inference is carried out based on training data generated by the engines.
-In this paper, we consider the common scenario where the learning problems share three key characteristics:
-  • Simulation engines are near deterministic: simulation outputs have vanishing aleatoric uncertainty, e.g. the atomic energy $V({\bf X})$ is a deterministic function of atomic positions ${\bf X}$.
-  • Surrogate models are misspecified: no single choice of the $P$ model parameters can match all $N$ training observations, such that model averaging is required to capture uncertainties.
-  • Surrogate models are underparametrized: large quantities of training data are available, such that the standard posterior from Bayesian inference gives vanishing parameter uncertainties.
-This typically occurs in the limit $P/N\ll 1$ for finite capacity models.
-Misspecification affects both finite capacity models and deep learning approaches with finite training resources  (Lahlou et al., 2021; Psaros et al., 2023).
-A misspecification-aware learning scheme should minimize the cross entropy between predicted and observed data distributions, known as the generalization error, or population risk (see (Alquier, 2021; Morningstar et al., 2022)).
-However, this is typically not tractable for learning, both due to numerical instabilities and the lack of theoretical bounds for convergence.
-Bayesian learning schemes instead minimize the expected loss (log likelihood), an upper bound to the generalization error, which is known to ignore misspecification but admits a robust learning scheme.
-Epistemic uncertainties are thus severely underestimated, vanishing in the underparametrized limit, with broad implications for surrogate model selection, uncertainty quantification and error propagation, as noted by multiple groups(Masegosa, 2020; Kato et al., 2022; Lotfi et al., 2022; Lahlou et al., 2021; Psaros et al., 2023; Imbalzano et al., 2021).
-We analyze the generalization error of deterministic, misspecified and under-parametrized surrogate models, i.e.
+Bayesian regression determines model parameters by minimizing the expected loss, an upper bound to the true generalization error.
+However, the loss ignores misspecification, where models are imperfect.
+Parameter uncertainties from Bayesian regression are thus significantly underestimated and vanish in the large data limit.
+This is particularly problematic when building models of low-noise, or near-deterministic, calculations, as the main source of uncertainty is neglected.
+We analyze the generalization error of misspecified, near-deterministic surrogate models, a regime of broad relevance in science and engineering.
+We show posterior distributions must cover every training point to avoid a divergent generalization error and design an ansatz that respects this constraint, which for linear models incurs minimal overhead.
+This is demonstrated on model problems before application to thousand dimensional datasets in atomistic machine learning.
+Our efficient misspecification-aware scheme gives accurate prediction and bounding of test errors where existing schemes fail, allowing this important source of uncertainty to be incorporated in computational workflows.
+Surrogate models are widely used across science and engineering to efficiently approximate the action of computationally intensive simulation engines  Alizadeh et al.
+(2020); Deringer et al.
+(2019); Lapointe et al.
+(2020); Nyshadham et al.
+(2019); Montes de Oca Zapiain et al.
+(2022); Bonatti et al.
+(2022); Kudela and Matousek (2022).
+Domain expertise is leveraged to design optimal features and model architecture before parameters are inferred from training data generated by the engines.
+In a broad range of applications, the inference problem shares three key characteristics:
+a) Simulation engines are near deterministic: outputs have vanishing aleatoric uncertainty, e.g. the atomic energy $V({\bf X})$ is deterministic to atomic positions ${\bf X}$.
+b) Surrogate models are misspecified: no single choice of the $P$ parameters can match all $N$ observations, meaning parameters are intrinsically uncertain.
+c) Surrogate models are underparametrized: large quantities of training data are available, such that standard epistemic uncertainty estimates from Bayesian inference vanish, typically corresponding to the regime $P/N\ll 1$.
+Misspecification affects both finite capacity models and deep learning approaches with finite training resources  Lahlou et al.
+(2021); Psaros et al.
+(2023).
+A misspecification-aware learning scheme should minimize the cross entropy between predicted and observed data distributions, known as the generalization error, or under certain technical definitions, the population risk (see Alquier (2021); Morningstar et al.
+(2022)).
+However, the generalization error is typically not tractable for learning, both due to numerical instabilities and the lack of theoretical bounds for convergence.
+Bayesian learning schemes instead minimize an upper bound, the expected loss (log likelihood), which ignores misspecification but admits a robust learning scheme.
+Epistemic uncertainties are thus severely underestimated, vanishing in the underparametrized limit, with broad implications for surrogate model selection, uncertainty quantification and error propagation, as noted by multiple groups Masegosa (2020); Kato et al.
+(2022); Lotfi et al.
+(2022); Lahlou et al.
+(2021); Psaros et al.
+(2023); Imbalzano et al.
+(2021).
+In this paper, we analyze the generalization error of deterministic, misspecified and under-parametrized surrogate models, i.e.
 those typically used to approximate simulation engines in science and engineering.
-Our focus is the underparametrized regime, meaning we do not focus on epistemic errors important for finite $N$.
-As posterior distributions become unbounded in the deterministic limit, this is only taken as a theoretical device to derive conditions that any posterior must obey to avoid a divergent generalization error.
-We then design an ensemble ansatz that respects this condition and admits a robust learning scheme at finite $N$, which reduces to a constrained loss minimization problem for large $N/P$.
+As posterior distributions strictly become unbounded in the deterministic limit, this is only taken as a theoretical device to derive conditions that any posterior must obey to avoid a divergent generalization error.
+We then design an ansatz that respect this condition and admits a robust learning scheme at finite $N$, with standard treatment of epistemic ($N$-dependent) errors, reducing to a constrained loss minimization problem for large $N/P$.
+Our final misspecification-aware parameter posterior has finite uncertainty even as $N/P\to\infty$ with a provably lower generalization error than the Gibbs posterior of Bayesian inference.
 Our main contributions are:
-  • We define a pointwise optimal parameter sets (POPS) for each training point, for which model predictions are exact.
+a) We define pointwise optimal parameter sets (POPS) for each training point, within which model predictions are exact.
 Parameter distributions must have mass in every POPS to avoid a divergent generalization error.
-An ensemble ansatz is proposed that respects POPS occupancy and gives a finite generalization error.
-  • For linear models, the variational minimum of our ansatz can be efficiently evaluated via rank-one updates to a leverage-weighted loss minimizer.
-This is our main result, an efficient scheme to incorporate misspecification into linear regression with minimal computational overhead.
-  • We compare our variational minimum to numerial optimization of our ensemble ansatz and show how our scheme provides robust bounds on test errors.
-This is confirmed in numerical experiments on challenging datasets from atomic machine learning.
-(Germain et al., 2016) established an important connection between Bayesian inference and probability approximately correct (PAC-Bayes) analysis for regression problems with unbounded losses, which bounds true expectations by empirical likelihood expectations through concentration inequalities.
+b) We use the ensemble of loss minimizers from each POPS to design two ansatz posteriors, a reweighted $\mathcal{O}(N)$ ensemble and a bounding $\mathcal{O}(P)$ hypercube, both of which respect POPS occupancy and give a finite generalization error, lower than the Gibbs posterior from inference.
+c) For linear models, our ansatz can be efficiently evaluated via rank-one updates to a leverage-weighted loss minimizer.
+This gives an efficient scheme to incorporate misspecification into linear regression with minimal computational overhead.
+d) We test the approach on synthetic datasets, showing how our scheme provides robust bounds on test errors.
+This performance is maintained under application to thousand-dimensional datasets from atomic machine learning, untreatable with existing methods.
+.1 Related work
+Germain and coworkers Germain et al.
+(2016) established an important connection between Bayesian inference and probability approximately correct (PAC-Bayes) analysis for regression problems with unbounded losses, which bounds true expectations by empirical likelihood expectations through concentration inequalities.
 This provides a robust learning scheme without test-train splitting, and gives a rationalization for Bayesian inference as a form of regularization.
 In particular, minimizing the PAC-Bayes bound for the expected loss gives the familiar relation between posterior, prior and likelihood from Bayesian inference.
-In recent years, multiple groups have used the PAC-Bayes framework to derive tighter bounds for the generalization error than that provided by the expected loss  (Masegosa, 2020; Germain et al., 2016; Lahlou et al., 2021; Psaros et al., 2023; Lotfi et al., 2022; Morningstar et al., 2022).
+In recent years, multiple groups have used the PAC-Bayes framework to derive tighter bounds for the generalization error than that provided by the expected loss  Masegosa (2020); Germain et al.
+(2016); Lahlou et al.
+(2021); Psaros et al.
+(2023); Lotfi et al.
+(2022); Morningstar et al.
+(2022).
 However, these efforts focus on probabilistic regression settings, often with neural network models far from the underparametrized limit, meaning aleatoric, epistemic and misspecification errors must be considered jointly.
-(Masegosa, 2020) derived second order PAC-Bayes bounds for the generalization error, minimized through a variational or ensemble ansatz.
-(Morningstar et al., 2022) developed estimators to assess disagreement between the generalization error and expected loss, deriving specialized PAC bounds for theoretical guarantees.
-(Lahlou et al., 2021) considered misspecification of neural networks, jointly training a minimum loss surrogate and an independent predictor of misspecification error.
-(Lotfi et al., 2022) considered misspecification in the context of deep model selection, deriving conditional likelihoods better aligned with the generalization error.
-As various definitions for the distinction between epistemic and misspecification uncertainties exist (Kato et al., 2022), we define misspecification as non-aleatoric uncertainties which survive in the underparametrized limit, where the minimizer of PAC-Bayes bounds for the expected loss, which coincides with the posterior from Bayesian inference, predict vanishing parameter uncertainties.
-This definition is made more precise in section II.3.
+Masegosa Masegosa (2020) derived second order PAC-Bayes bounds for the generalization error minimized through a variational or ensemble ansatz.
+Morningstar et al.
+Morningstar et al.
+(2022) developed estimators to assess disagreement between the generalization error and expected loss, deriving specialized PAC bounds for theoretical guarantees.
+Lahlou et al.
+Lahlou et al.
+(2021) considered misspecification of neural networks, jointly training a minimum loss surrogate and an independent predictor of misspecification error.
+Lofti et al.
+Lotfi et al.
+(2022) considered misspecification in the context of deep model selection, deriving conditional likelihoods better aligned with the generalization error.
+As various definitions for the distinction between epistemic and misspecification uncertainties exist Kato et al.
+(2022), we define misspecification as non-aleatoric uncertainties which survive in the underparametrized limit, where the minimizer of PAC-Bayes bounds for the expected loss, which coincides with the posterior from Bayesian inference, predict vanishing parameter uncertainties.
+This definition is made more precise in section I.4.
 Our epistemic uncertainties thus coincide with standard estimates from loss minimization.
-We summarize known results for the generalization error and expected loss, PAC-Bayes bounds for the regression loss, and the sub-optimality of loss minimization for misspecified problems in section II.
-Section III contains our central theoretical contribution, an analysis of the generalization error for near-deterministic surrogate models and a simple ensemble ansatz which can be efficiently evaluated for linear models.
-Section IV applies our ansatz to high-dimensional linear regression problems, using synthetic data and challenging applications in atomic machine learning.
-II Bayesian surrogate models
-II.1 Deterministic Simulation Engines
+I Bayesian surrogate models
+I.1 Deterministic simulation engines
 A simulation engine ${\mathcal{E}}:{\mathcal{X}}\to{\mathcal{Y}}$ takes input ${\mathbf{X}}\in{\mathcal{X}}$ and returns ${\mathbf{Y}}\in{\mathcal{Y}}$.
@@
 property is only satisfied for decorrelated segments ${\mathbf{X}}=\{{\bf x}_{t}\}_{t=1}^{T}$, meaning ${\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})$ captures signal correlations as in generalized least squares regression.
-II.2 Deterministic Surrogate Models and the Generalization Error
+I.2 Deterministic surrogate models
 A surrogate model ${\mathcal{M}}:{\mathcal{X}}\to{\mathcal{Y}}$ for ${\mathcal{E}}$ is defined by $P$ parameters ${\bm{\Theta}}\in{\mathcal{P}}$.
@@
 $\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}})$ $=\int_{\mathcal{P}}\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}.$ (5)
-Unlike typical Bayesian schemes, we do not treat the covariance term ${\bm{\Sigma}}_{\mathcal{Y}}$ as fitting parameter but as an intrinsic property of the quasi-deterministic simulation engine ${\mathcal{E}}$.
-This has implications for model selection criteria, as we discuss in appendix B.
+Unlike typical Bayesian schemes, we do not treat the covariance term ${\bm{\Sigma}}_{\mathcal{Y}}$ as fitting parameter but as an intrinsic property of ${\mathcal{E}}$.
+This has implications for model selection criteria, discussed in the supplementary material (SM) Swinburne and Perez (2024).
+I.3 The generalization error and expected loss
 The generalization error ${\mathcal{G}}[\pi]$ is simply the cross entropy of the predicted distribution $\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}})\rho_{N}({\mathbf{X}})$ to the observed distribution $\rho({\mathbf{Y}}|{\mathbf{X}})\rho_{N}({\mathbf{X}})$; omiting the constant term ${\mathcal{G}}_{\mathcal{E}}\equiv\left\langle\ln\left|\rho_{\mathcal{E}}({\mathbf{Y}}|{\mathbf{X}})\right|\right\rangle_{\mathcal{D}}$ we have
 ${\mathcal{G}}[\pi]=-\left\langle\ln\left|\int_{\mathcal{P}}\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}\right|\right\rangle_{\mathcal{D}}.$ (6)
-II.3 The expected loss and PAC estimation
-PAC-Bayes analysis (Alquier, 2021) provides concentration inequalities (Hoeffding, 1994) that bound expectations (3) over ${\mathcal{D}}$ by empirical expectations over ${\mathcal{D}}_{N}$ with some probability $1-\xi$.
-Whilst originally designed for classification problems where the loss is bounded to $[0,1]$, the treatment can be extended to regression settings with unbounded losses, as discussed in (Germain et al., 2016) and appendix A.
-In this setting, available PAC-Bayes bounds are for $\langle f({\mathbf{Y}},{\mathbf{X}})\rangle_{\mathcal{D}}$, where $f({\mathbf{Y}},{\mathbf{X}})$ to be bounded (i.e.
-$\epsilon>0$) and can expressed as an expectation over $\pi({\bm{\Theta}})$, i.e.
-$\langle f({\mathbf{Y}},{\mathbf{X}})\rangle_{\mathcal{D}}=\int_{\mathcal{P}}\langle f({\mathbf{Y}},{\mathbf{X}},{\bm{\Theta}})\rangle_{\mathcal{D}}\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}$.
-As ${\mathcal{G}}[\pi]$ is not in this form, the bound used in PAC-Bayesian schemes is for the expected loss ${\mathcal{L}}[\pi]\geq{\mathcal{G}}[\pi]$, often known as the negative log likelihood.
-The Jensen inequality for convex functions $f(\langle x\rangle)\leq\langle f(x)\rangle$ gives, with $f(x)=-\ln|x|$,
-$\mathcal{G}[\pi]\leq{\mathcal{L}}[\pi]=-\int_{\mathcal{P}}\left\langle\ln\left|\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\right|\right\rangle_{\mathcal{D}}\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}.$
-As discussed by (Germain et al., 2016) the expected regression loss ${\mathcal{L}}[\pi]$ admits the PAC-Bayes upper bound, holding with probability $1-\xi$, of
-${\mathcal{L}}[\pi]$$\leq\hat{{\mathcal{L}}}_{N}[\pi]+\frac{1}{N}\left(KL(\pi|\pi_{0})-\ln|\xi|\right)+C_{0}[\pi_{0}],$
-^$\hat{{\mathcal{L}}}_{N}[\pi]$ $\equiv\int_{\mathcal{P}}\hat{L}_{N}({\bm{\Theta}}){\rm d}\pi({\bm{\Theta}}),$ (7)
-where $\hat{L}_{N}({\bm{\Theta}})=-\langle\ln|\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\rangle_{N}$ is the empirical loss over $\rho_{N}({\mathbf{X}})$, $KL(\pi|\pi_{0})=\int_{\mathcal{P}}\ln|\pi/\pi_{0}|{\rm d}\pi({\bm{\Theta}})$ is the cross entropy between $\pi$ and some prior $\pi_{0}({\bm{\Theta}})$ and $C_{0}[\pi_{0}]$ is constant with respect to $\pi$ and has a finite upper bound discussed in appendix A.
-The minimizer of the PAC-Bayes upper bound (7) is that of ^$\hat{{\mathcal{L}}}_{N}[\pi]+KL(\pi|\pi_{0})/N$, giving the posterior from Bayesian inference (Germain et al., 2016)-
+In principle any learning scheme should aim to minimize the non-convex functional ${\mathcal{G}}[\pi]$, but in practice learning schemes target a convex upper bound, the expected loss ${\mathcal{L}}[\pi]\geq{\mathcal{G}}[\pi]$, often known as the negative log likelihood.
+Using the Jensen inequality $-\langle\ln x\rangle\leq-\ln\langle x\rangle$ gives
+$\mathcal{G}[\pi]\leq{\mathcal{L}}[\pi]=-\int_{\mathcal{P}}\left\langle\ln\left|\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\right|\right\rangle_{\mathcal{D}}\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}.$ (7)
+${\mathcal{L}}[\pi]$ is clearly minimized by a delta function $\delta({\bm{\Theta}}-{\bm{\Theta}}^{*}_{\mathcal{L}})$, where ${\bm{\Theta}}^{*}_{\mathcal{L}}\in\arg\min_{\mathcal{P}}\left\langle-\ln\left|\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\right|\right\rangle_{\mathcal{D}}$, i.e.
+vanishing parameter uncertainty.
+The next section shows how this emerges from the inference posterior at finite $N/P$.
+I.4 PAC-Bayes estimation and connection to inference
+PAC-Bayes analysis Alquier (2021) provides concentration inequalities Hoeffding (1994) that bound expectations over ${\mathcal{D}}$ by empirical expectations over ${\mathcal{D}}_{N}$ with some probability.
+Whilst originally designed for classification problems, the approach has been extended to regression settings with unbounded losses Germain et al.
+(2016).
+As discussed in the SM Swinburne and Perez (2024), The expected loss ${\mathcal{L}}[\pi]$ admits the PAC-Bayes upper bound, holding with probability $1-\xi$, of
+${\mathcal{L}}[\pi]\leq\int_{\mathcal{P}}\hat{L}_{N}({\bm{\Theta}})+\frac{1}{N}\ln\left|\frac{\pi}{\pi_{0}}\right|{\rm d}\pi({\bm{\Theta}})-\frac{\ln|\xi|}{N}+C_{0},$ (8)
+where $\hat{L}_{N}({\bm{\Theta}})=-\langle\ln|\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\rangle_{N}$ is the empirical loss over $\rho_{N}({\mathbf{X}})$, $\pi_{0}({\bm{\Theta}})$ is some prior and $C_{0}$ is a constant Swinburne and Perez (2024).
+The minimizer of the upper bound (21) is the posterior from Bayesian inference Germain et al.
+(2016)
 $\pi_{{\mathcal{L}},N}({\bm{\Theta}})=\pi_{0}({\bm{\Theta}})\exp(\lambda-N\hat{L}_{N}({\bm{\Theta}})).$ (9)
-Appendix B discusses this bound in the context of Bayesian model selection. Using Laplace’s method (appendix C) as $N/P\to\infty$ gives $\hat{L}_{N}({\bm{\Theta}})\to L({\bm{\Theta}})$, with
+As $N/P\to\infty$ the posterior is strongly peaked around a loss minimizer ${\bm{\Theta}}^{*}_{\mathcal{L}}\in\arg\min_{{\bm{\Theta}}\in{\mathcal{P}}}L({\bm{\Theta}})$.
+Defining ${\bm{\Sigma}}^{-1}_{0}$ as the curvature of $-\ln\pi_{0}({\bm{\Theta}})$ around ${\bm{\Theta}}^{*}_{\mathcal{L}}$, we use Laplace’s method Swinburne and Perez (2024) as $N/P\to\infty$ to write
 […]
-[…]
-[…]
-$L({\bm{\Theta}})=-\langle\ln|\mathcal{N}({\mathcal{E}}({\mathbf{X}})|{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}}),{\bm{\Sigma}}_{\mathcal{Y}})|\rangle_{\mathcal{X}},$ (10)
 […]
@@
 ${\bm{\Sigma}}^{*}_{\mathcal{L}}=\langle[{\mathcal{E}}({\mathbf{X}})-{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}}^{*}_{\mathcal{L}})][{\mathcal{E}}({\mathbf{X}})-{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}}^{*}_{\mathcal{L}})]^{\top}\rangle$ is the error covariance around ${\bm{\Theta}}^{*}_{\mathcal{L}}$ and we use ${\rm Tr}({\bm{\Sigma}}_{\mathcal{Y}})=\mathcal{O}(\epsilon^{2})$.
-As noted by (Masegosa, 2020), it is clear that the maximum likelihood parameter distribution $\pi_{\mathcal{L}}({\bm{\Theta}})$ is not a minimizer of the generalization error under misspecification.
-We emphasize that the primary motivation of this section is to show that whilst the generalization error is the central quantity of interest for learning under misspecification, we typically only have access to an upper bound in the form of a PAC-Bayes concentration inequality (7) for the expected loss.
-Under model misspecification it is clear that minimization of this upper bound in sub-optimal, in that the generalization error diverges as $1/\epsilon^{2}$.
-In the next section we analyze the generalization error as $\epsilon\to 0$ to derive a condition for $\pi({\bm{\Theta}})$ that ensures a finite generalization error.
-We then design an ensemble ansatz which respects this condition and can be parametrized through a variational minimization of the PAC-Bayes upper bound (7).
-III Analysis of the generalization error in the near-deterministic limit
+Under misspecification it is thus clear that minimizing this upper bound is sub-optimal Masegosa (2020) as the generalization error diverges as $1/\epsilon^{2}$.
+In the next section we analyze the generalization error as $\epsilon\to 0$ to derive a condition for any $\pi({\bm{\Theta}})$ that ensures a finite generalization error.
+II Analysis of the generalization error in the near-deterministic limit
 We consider the generalization error (6) in the near-deterministic limit, ${\mathcal{G}}_{0}[\pi]\equiv\lim_{\epsilon\to 0}{\mathcal{G}}[\pi]$.
-The integral over ${\mathbf{Y}}$ in (6) concentrates at ${\mathbf{Y}}={\mathcal{E}}({\mathbf{X}})$; using $\langle\dots\rangle_{\mathcal{X}}\equiv\int\dots\rho_{\mathcal{D}}({\mathbf{X}}){\rm d}{\mathbf{X}}$ yields, using Laplace’s method
+The integral over ${\mathbf{Y}}$ in (6) concentrates at ${\mathbf{Y}}={\mathcal{E}}({\mathbf{X}})$; using $\langle\dots\rangle_{\mathcal{X}}\equiv\int\dots\rho_{\mathcal{D}}({\mathbf{X}}){\rm d}{\mathbf{X}}$ yields, using Laplace’s method Swinburne and Perez (2024)
 ${\mathcal{G}}_{0}[\pi]$ $=-\left\langle\ln\left|\rho_{\mathcal{M}}({\mathcal{E}}({\mathbf{X}})|{\mathbf{X}})\right|\right\rangle_{\mathcal{X}}$ (12)
@@
 Importantly, when (15) is satisfied, the predictive distribution $\rho_{\mathcal{M}}({\mathbf{Y}},{\mathbf{X}})$ covers every training data point, i.e.
-model predictions are guaranteed to envelope observations as $N/P\to\infty$.
-III.1 POPS-constrained loss minimizing ensemble ansatz
-The covering constraint (15) shows the minimizer of ${\mathcal{G}}_{0}$ must have mass in the POPS of each training point.
-This suggests that a representation in terms of a discrete set of models is likely to capture its key features.
-Selecting some set of pointwise fits $\mathcal{T}_{E}=\{{\bm{\Theta}}_{\mathbf{X}}\}_{{\mathbf{X}}\in{\mathcal{X}}}$, ${\bm{\Theta}}_{\mathbf{X}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})$ for each ${\mathbf{X}}\in{\mathcal{X}}$ and some weighting ${\rm w}_{E}({\mathbf{X}})>0$, $\langle{\rm w}_{E}({\mathbf{X}})\rangle_{\mathcal{X}}=1$, equation (15) is satisfied by the Gaussian ensemble ansatz
-$\pi_{E}({\bm{\Theta}})\equiv\langle{\rm w}_{E}({\mathbf{X}})\mathcal{N}({\bm{\Theta}}|{\bm{\Theta}}_{{\mathbf{X}}},\sigma^{2}\mathbb{I})\rangle_{\mathcal{X}},$ (16)
-where $\sigma$ is fixed and assumed small.
-Defining mass function for the ensemble which measures the number of pointwise fits ${\bm{\Theta}}_{{\mathbf{X}}^{\prime}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})$
-$m_{E}({\mathbf{X}})\equiv\left\langle{\rm w}_{E}({\mathbf{X}}^{\prime}){\rm m}_{\mathcal{E}}({\bm{\Theta}}_{{\mathbf{X}}^{\prime}},{\mathbf{X}})\right\rangle_{{\mathcal{X}}},$ (17)
-the generalization error of (16) reads, in the limit $\sigma\to 0$,
-${\mathcal{G}}_{0}[\pi_{E}]=-\left\langle\ln|m_{E}({\mathbf{X}})|\right\rangle_{{\mathcal{X}}}.$ (18)
-The above result shows to minimize ${\mathcal{G}}_{0}[\pi_{E}]$ we should select fits ${\bm{\Theta}}_{\mathbf{X}}\in\mathcal{T}_{E}$ to have maximum occupancy across all other POPS.
-In the worst case $\mathcal{T}_{E}\cap{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})={\bm{\Theta}}_{{\mathbf{X}}}$, $\forall{\mathbf{X}}\in{\mathcal{X}}$, ${\mathcal{G}}_{0}[\pi_{E}]$ still has a finite upper bound as $\sigma\to 0$ of
-${\mathcal{G}}_{0}[\pi_{E}]\leq-\left\langle\ln|{\rm w}_{E}({\mathbf{X}})|\right\rangle_{{\mathcal{X}}}+\mathcal{S}_{{\mathcal{X}}},$ (19)
-where $\mathcal{S}_{{\mathcal{X}}}=-\left\langle\ln|\rho({\mathbf{X}})|\right\rangle_{{\mathcal{X}}}$ is the entropy of the input distribution.
-In appendix D we show how each ${\bm{\Theta}}_{\mathbf{X}}\in\mathcal{T}_{E}$ can be determined in a variational setting at finite $\sigma$, using (16) in the PAC-Bayes loss bound (7) for the regression loss ${\mathcal{L}}[\pi_{E}]=\langle{\rm w}_{E}({\mathbf{X}})L({\bm{\Theta}}_{\mathbf{X}})\rangle_{\mathcal{X}}$, which as $P/N\to 0$, $\sigma\to 0$ reduces to POPS-constrained loss minimization
-${\bm{\Theta}}^{*}_{\mathbf{X}}$ $\equiv\arg\min_{{\bm{\Theta}}_{\mathbf{X}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})}L({\bm{\Theta}}_{\mathbf{X}}),$ (20)
-$\pi^{*}_{E}({\bm{\Theta}})$ $\equiv\langle{\rm w}_{E}({\mathbf{X}})\delta({\bm{\Theta}}-{\bm{\Theta}}^{*}_{{\mathbf{X}}})\rangle_{\mathcal{X}}.$ (21)
-The goal of this POPS-constrained loss-minimization is to ensure the pointwise fits will be as close as possible; for convex loss landscapes, this is measured by the Fisher information metric ${\bm{\mathcal{I}}}^{*}_{\mathcal{L}}$.
-By minimizing the ensemble spread, we aim to distribute an (albeit small) mass to other POPS, thereby reducing ${\mathcal{G}}_{0}$.
-In addition, for specified regression problems $\pi^{*}_{E}({\bm{\Theta}})$ then reduces to the true minimizer $\pi^{*}_{\mathcal{L}}({\bm{\Theta}})=\delta({\bm{\Theta}}-{\bm{\Theta}}^{*}_{\mathcal{L}})$.
-When the loss landscape is not convex additional constraints will have to be imposed to minimize (and measure) the ensemble spread.
-More generally, POPS-constrained loss minimization with our ensemble ansatz is potentially sub-optimal in that $\pi^{*}_{E}({\bm{\Theta}})$ chooses ${\bm{\Theta}}^{*}_{\mathbf{X}}$ to minimize the average prediction error across all other ${\mathbf{X}}^{\prime}\in{\mathcal{X}}$ but is not guaranteed to be a minimizer of (18).
-Exploring this tradeoff (multiple POPS occupancy vs global prediction error) would entail a combinatorial scheme which (21) avoids.
-One strategy, to be investigated in future work, would be to cluster the loss-minimizing pointwise fits, try to find POPS-intersections for each cluster, then replace all members of that cluster with a single POPS-intersecting fit.
-A further limitation of (21) is that many pointwise fits may be near-degenerate, raising the generalization error and, in general, underestimating the average deviation of any test point from the minimum loss prediction.
-However, the predicted max/min range of any test point is expected to be robust in the limit $N/P\to\infty$, as this is invariant to how we weight the ensemble.
-Furthermore, even if it were possible to find the true minimizer $\pi^{*}_{\mathcal{G}}({\bm{\Theta}})$ of ${\mathcal{G}}_{0}$, it is not clear a priori that the resulting distribution of predicted test errors would be close to observed test errors.
-The covering constraint (15) only ensures that any minimizer of ${\mathcal{G}}_{0}$ should produce predictions which bound any test point, i.e.
-provide a max/min envelope on model errors.
-In practice, we will show that uniformly sampling from either the parameters or posterior predictions of the POPS-constrained loss minimizing ensemble yields very good results for a diverse set of linear regression problems, which we demonstrate in the remainder of this paper.
-III.2 POPS-constrained ensemble for linear regression
+model predictions will envelope observations as $N/P\to\infty$.
+We now investigate two candidate distributions which satisfy (15), an $\mathcal{O}(N)$ ensemble ansatz and $\mathcal{O}(P)$ hypercube ansatz, finding approximate variational minima that allow for efficient deployment on high-dimensional datasets.
+II.1 POPS-constrained ensemble and hypercube ansatz
+The covering constraint (15) requires $\pi({\bm{\Theta}})$ to have mass in the POPS of each training point.
+This condition must be satisfied whilst concentrating mass in as small a region in ${\mathcal{P}}$ as possible to minimize ${\mathcal{G}}_{0}$.
+To achieve this, we first select a set of pointwise fits ${\bm{\Theta}}^{*}_{\mathbf{X}}\in\mathcal{T}^{*}_{E}$ that are each POPS-constrained loss minimizers
+${\bm{\Theta}}^{*}_{\mathbf{X}}\in\mathcal{T}^{*}_{E},\quad{\bm{\Theta}}^{*}_{\mathbf{X}}\equiv\arg\min_{{\bm{\Theta}}_{\mathbf{X}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})}L({\bm{\Theta}}_{\mathbf{X}}).$ (16)
+These pointwise fits will be closely clustered around the minimum loss solution ${\bm{\Theta}}^{*}_{\mathcal{L}}$, as measured by the Fisher metric ${\boldsymbol{\mathcal{I}}}^{*}_{\mathcal{L}}$.
+In the supplementary materials Swinburne and Perez (2024) we explore a weighted ensemble ansatz
+$\pi^{*}_{E}({\bm{\Theta}})=\langle{\rm w}({\mathbf{X}})\delta({\bm{\Theta}}-{\bm{\Theta}}^{*}_{\mathbf{X}})\rangle,$ (17)
+where ${\rm w}({\mathbf{X}})>0$.
+In particular, we show $\mathcal{T}^{*}_{E}$ naturally emerges when choosing pointwise fits ${\bm{\Theta}}_{\mathbf{X}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})$ to minimize the ansatz expected loss $\mathcal{L}[\pi_{E}]=\langle{\rm w}({\mathbf{X}})L({\bm{\Theta}}_{X})\rangle$.
+The ensemble ansatz produces highly informative max/min bounds for any test point, providing an envelope for worst case errors, as we show in figure 1.
+However, even when employing an optimal reweighting ${\rm w}^{*}({\mathbf{X}})$ (see Swinburne and Perez (2024)), the ensemble ansatz systematically underestimates moments of the test error distribution, as shown in figure 2.
+In addition, both resampling and storage are $\mathcal{O}(N)$, which can become problematic for the $N/P\to\infty$ limit of interest.
+For practical applications, the parameter distribution should ideally require at most $\mathcal{O}(P^{2})$ effort for storage and $\mathcal{O}(P)$ for resampling.
+We have found this can be achieved by finding the minimal hypercube $\mathcal{H}(\mathcal{T}^{*}_{E})\in\mathcal{P}$ which encompasses all members of the loss minimizing POPS-ensemble ${\bm{\Theta}}^{*}_{\mathbf{X}}\in\mathcal{T}^{*}_{E}$.
+As $\mathcal{H}(\mathcal{T}^{*}_{E})$ intersects all POPS by construction, equation (15) is satisfied by the hypercube ansatz uniform over $\mathcal{H}(\mathcal{T}^{*}_{E})$, i.e.
+$\pi^{*}_{\mathcal{H}}({\bm{\Theta}})=V^{-1}_{\mathcal{H}}\delta({\bm{\Theta}}\in\mathcal{H}(\mathcal{T}^{*}_{E})),$ (18)
+where $V_{\mathcal{H}}=\int_{\mathcal{P}}\delta({\bm{\Theta}}\in\mathcal{H}(\mathcal{T}^{*}_{E})){\rm d}{\bm{\Theta}}$.
+Crucially, resampling $\pi_{\mathcal{H}}({\bm{\Theta}})$ requires only $\mathcal{O}(P)$ effort whilst giving a conservative estimate and bounding of test errors.
+To determine $\mathcal{H}(\mathcal{T}^{*}_{E})$ we use singular value decomposition on the POPS-ensemble $\mathcal{T}^{*}_{E}$ to obtain the right eigenmatrix ${\bf V}\in\mathbb{R}^{R\times P}$, $R\leq P$, ${\bf V}{\bf V}^{\top}=\mathbb{I}_{R}$, requiring $\mathcal{O}(P^{3})$ effort for evaluation.
+Uniformly sampling the full max/min hypercube of the projected ensemble $\tilde{\mathcal{T}}^{*}_{E}=\{{\bf V}{\bm{\Theta}}|{\bm{\Theta}}\in{\mathcal{T}^{*}_{E}}\}$ returns samples $\tilde{{\bm{\Theta}}}^{\top}{\bf V}$ of $\pi^{*}_{\mathcal{H}}$ for $\mathcal{O}(R)$ effort and $\mathcal{O}(RP)$ storage.
+We find this efficient hypercube posterior to give excellent bounding and prediction of test errors for $P\in[3,1600]$, as shown in figures 2 and 3.
+An additional advantage is the simplicity of implementation for linear models; an example Python code is provided in the SM Swinburne and Perez (2024).
+II.2 POPS-constrained loss minimization for linear models
+Both the ensemble and hypercube ansatz require POPS-constrained loss minimizing fits ${\bm{\Theta}}^{*}_{\mathbf{X}}\in\mathcal{P}_{\mathcal{M}}({\mathbf{X}})$.
+Practical application therefore requires an efficient means to perform this constrained minimization.
 Linear models use feature functions $\mathcal{F}:{\mathcal{X}}\to{\mathcal{P}}\times{\mathcal{Y}}$ to map inputs ${\mathbf{X}}\in{\mathcal{X}}$ to features ${\mathbf{F}}({\mathbf{X}})\in{\mathcal{P}}\times{\mathcal{Y}}$, giving ${\mathcal{M}}({\mathbf{X}},{\bm{\Theta}})={\bm{\Theta}}\cdot{\mathbf{F}}({\mathbf{X}})$ with parameters ${\bm{\Theta}}\in{\mathcal{P}}$.
-Minimizing the PAC-Bayes bound (7) with a Gaussian prior $\pi_{0}({\bm{\Theta}})=\mathcal{N}({\bm{\Theta}}|{\bf 0},{\bm{\Sigma}}_{0})$ gives a minimizer ${\bm{\Theta}}^{*}_{\mathcal{L}}$ of
+Minimizing the PAC-Bayes bound (21) with a Gaussian prior $\pi_{0}({\bm{\Theta}})=\mathcal{N}({\bm{\Theta}}|{\bf 0},{\bm{\Sigma}}_{0})$ gives a global loss minimizer ${\bm{\Theta}}^{*}_{\mathcal{L}}$ of
 ${\bm{\Theta}}^{*}_{\mathcal{L}}$ $={\bf A}\langle{\mathbf{F}}({\mathbf{X}}){\mathcal{E}}({\mathbf{X}})\rangle_{\mathcal{X}},$ (19)
 ${\bf A}$ $=[{\bm{\Sigma}}_{\mathcal{Y}}{\bm{\Sigma}}^{-1}_{0}/N+\langle{\mathbf{F}}({\mathbf{X}})[{\mathbf{F}}({\mathbf{X}})]^{\top}\rangle]^{-1}_{\mathcal{X}}.$
-We discuss how the epistemic covariance for small $N/P$ can be incorporated in a variational PAC-Bayes scheme in Appendix D.
+The epistemic covariance for small $N/P$ can be incorporated in a variational PAC-Bayes scheme Swinburne and Perez (2024) that will be developed further in future work.
 However, in the following, we concentrate on misspecification uncertainties in the limit $N/P\to\infty$, where the influence of $\bm{\Sigma}_{0}$ is negligible.
@@
 The minimum loss solution is the leverage-weighted centroid of pointwise optima, ${\bm{\Theta}}^{*}_{\mathcal{L}}=\langle{\rm h}({\mathbf{X}}){\bm{\Theta}}^{*}_{{\mathbf{X}}}\rangle_{\mathcal{X}}$.
-Equation (23) is our third main result, a scheme to incorporate misspecification uncertainty into least-squares linear regression for minimal overhead, requiring only $\mathcal{O}(N)$ inner products over feature vectors.
-Implementation in least squares linear regression is trivial, as shown in appendix F.
-We now investigate how our POPS-constrained ensemble performs in a variety of high-dimensional regression tasks.
-IV Numerical Experiments
-The supplementary material provides python/JaX code to reproduce all figures, with the exception of the large atomic machine learning datasets.
-IV.1 Comparison with direct minimization of $\mathcal{G}$
-For low dimensional problems, it is possible to approximately minimize an empirical generalization error $\mathcal{G}$ over discrete model ensembles, using a finite aleatoric uncertainty ${\bm{\Sigma}}_{\mathcal{Y}}$ for regularization.
-With increasing ${\bm{\Sigma}}_{\mathcal{Y}}$ it is clear that $\mathcal{G}\to\mathcal{L}$ and thus we expect the minimum loss solution $\pi_{\mathcal{L}}({\bm{\Theta}})$ to emerge.
-We also compare these appraoches to Bayesian ridge regression, corresponding to minimizing a first-order PAC-Bayes loss bound, and an ensemble approach from citepmasegosa2020learning, section C.3, which corresponds to minimizing a second-order PAC-Bayes loss bound, whose loss function, as for $\mathcal{G}_{E}$, also requires exponentiating model errors.
-At small ${\bm{\Sigma}}_{\mathcal{Y}}$ direct minimization of $\mathcal{G}_{E}$ and the Masegosa ensemble bound typically fails, requiring careful parameter tuning and remaining at low $P<6$ to avoid vanishing gradients or other numerical issues.
-The POPS-ensemble, in contrast, can easily sample $P=\mathcal{O}(1000)$ or higher dimensional data.
-Figure 1 shows a simple example fitting a one dimensional quadratic polynomial to a sinusoidal function, with the constant, linear and quadratic features giving $P=3$ model parameters.
-However, the features are highly correlated, simplifying the regression task.
-We show results for ${\bm{\Sigma}}_{\mathcal{Y}}=1/3$, at the lower limit for which minimization converges for all methods.
-As expected from section II.3 Bayesian ridge regression significantly underestimates test errors even at the ±$\pm 3\sigma$ level, which as discussed above will worsen with increasing $N/P$.
-We find similar behaviour for the Masegosa ensemble, which we note is primarily designed for cases with finite aleatoric uncertainty.
-Direct minimization of $\mathcal{G}_{E}$ gives better bounding of test errors, though still with notable underestimates in high-density regions, though bounnds became tighter with lower ${\bm{\Sigma}}_{\mathcal{Y}}$ (whilst stable) However, the POPS-constrained ensemble gives almost perfect bounding of test errors, being stable for arbitrarly small ${\bm{\Sigma}}_{\mathcal{Y}}$ values.
+Equation (20) is our third main result, a scheme to incorporate misspecification uncertainty into least-squares linear regression for minimal overhead, requiring only $\mathcal{O}(N)$ inner products over feature vectors, which can then be used to construct the ensemble or hypercube ansatz given above.
+A simple python implementation is provided in the SM Swinburne and Perez (2024).
+III Numerical Experiments
+III.1 Comparison to Bayesian ridge regression and minimization of $\mathcal{G}_{E}$
+We applied Bayesian ridge regression and our POPS-ansatz to regress a quadratic polynomial against a sinusoidal function, the constant, linear and quadratic features giving $P=3$ model parameters.
+The SM Swinburne and Perez (2024) shows the same procedure for a model with $P=6$ independent features.
+For these simple models it is possible to approximately minimize the generalization error $\mathcal{G}_{E}$ of some discrete model ensemble, providing we regularize with a finite aleatoric uncertainty ${\bm{\Sigma}}_{\mathcal{Y}}=\sigma{\bm{\Sigma}}^{*}_{\mathcal{L}}$.
+With increasing $\sigma$ we have $\mathcal{G}\to\mathcal{L}$ and thus expect the ensemble to concentrate on ${\bm{\Theta}}^{*}_{\mathcal{L}}$, whilst in the small $\sigma$ regime of interest minimizing $\mathcal{G}_{E}$ becomes increasingly unstable.
+We found the same small $\sigma$ instability when applying an ensemble approach from Masegosa (2020) which targets a second-order PAC-Bayes loss bound, designed for misspecified probabilistic models where $\sigma$ can be large.
+These instabilities show existing misspecification-aware methods require careful hyperparameter tuning and low model dimension to avoid vanishing gradients or other numerical issues.
+Our POPS-ansatz, in contrast, can easily sample $P=\mathcal{O}(1000)$ or higher dimensional data.
+Figure 1 shows results for $\sigma=1,1/3,1/6$, at the lower limit for which minimization converges.
+As expected from section I.4, Bayesian ridge regression significantly underestimates test errors even at the 99.7% confidence level (three standard deviations) which as discussed above will worsen with increasing $N/P$.
+The max/min range of the numerical ensemble shows similar behavior for $\sigma=1,1/3$ whilst at $\sigma=1/6$ test errors are bounded but significantly overestimated, which worsens as $\sigma$ decreases (whilst stable).
+This was also seen on higher dimensional datasets Swinburne and Perez (2024).
+Importantly, in all cases our POPS ansatz $\pi^{*}_{E}$ and $\pi^{*}_{\mathcal{H}}$ give excellent bounding of test errors and are stable in the near-deterministic limit $\sigma\to 0$.
+As expected, the hypercube ansatz $\pi^{*}_{\mathcal{H}}$ gave slightly more conservative bounds.
 Figure 1: Regression of a deterministic quadratic polynomial ($P=3$) model onto a sinusoidal ”simulation engine”, trained on $N=100$ points.
-Clockwise from top left: mean and $3\sigma$ confidence interval from Bayesian ridge regression , direct minimization of $\mathcal{G}_{E}$, regularized with $\Sigma_{\mathcal{Y}}=1$, the min/max range of the POPS-ensemble (23), and $\mathcal{G}_{E}$ with $\Sigma_{\mathcal{Y}}=0.1$.
-We applied the same procedure to a model system with $P=6$ decorrelated features, unlike the polynomial example, to simulate more realistic datasets.
-The minimization routine could only treat a small range of $\Sigma_{\mathcal{Y}}$ values, which we normalize by the loss minimizer RMS error $\sigma_{MLE}$.
-Figure 2 shows the resulting parameter distributions.
-At large $\Sigma_{\mathcal{Y}}$ the minimizer quickly concentrates all ensemble members on the minimum loss solution, whilst at $\Sigma_{\mathcal{Y}}=\sigma_{MLE}$ the final distribution is very close to the POPS-ensemble.
-At lower values the final distribution was a little broader, but we attribute this largely to numerical instabilities as the final result is sensitive to initial conditions (unlike the first two cases).
-In addition, a minimizer initialized with the POPS-ensemble did not change appreciably.
-These albeit simple results give strong evidence that the POPS-ensemble is indeed a stable and highly efficient means to find approximate minimizers of ${\mathcal{G}}$ that would otherwise be intractable.
-Figure 2: Regression of a $P=6$ linear model onto a cubic ”simulation engine”, trained on $N=500$ points.
-The panels show the culmulative distribution of each parameter across the ensemble (same size as $N$) for the loss minimizer (MLE), the POPS-ensemble (23) and from direct minimization of ${\mathcal{G}}$ with $\Sigma_{\mathcal{Y}}/\sigma_{MLE}=2,1,1/2$.
-IV.2 Test errors in medium-dimensional linear regression
-We now look at higher/medium dimensional problems, $P=10-100$, where direct minimization of ${\mathcal{G}}$ is not practical.
-The simulation engines (see supplementary material) were constrained to be misspecified from the linear surrogate models by combining features in quadratic, cubic or sinusoidal functions or randomly sampling the linear feature coefficients from some predetermined subset of models.
+Top left: mean and $3\sigma$ interval from Bayesian ridge regression.
+All other plots show mean and max/min range.
+Top: POPS-ensemble and POPS-hypercube ansatz $\pi^{*}_{E}$ and $\pi^{*}_{\mathcal{H}}$.
+Bottom: numerically optimized $\mathcal{G}_{E}$ for a uniformly weighted $N$-ensemble, regularized with $\Sigma_{\mathcal{Y}}=\sigma\Sigma^{*}_{\mathcal{L}}$, for $\sigma=1,1/3,1/6$.
+Lower $\sigma$ values gave numerical instabilities.
+III.2 Test errors in medium-dimensional linear regression
+We now consider higher/medium dimensional problems, $P=10-50$, where numerical minimization of ${\mathcal{G}}_{E}$ is not possible.
+The simulation engines were constrained to be misspecified from the linear surrogate models by combining features in quadratic, cubic or sinusoidal functions or randomly sampling the linear feature coefficients from some predetermined subset of models.
 A noise term that simulates the presence of variability orthogonal to the chosen feature set was also used.
 In all cases, independent test and training data were generated, typically with a 10:90 test:train split.
-However, all conclusions remain robust to varying the precise nature of data generation and the classes of engine considered.
-Additional examples are provided in the supplementary material and appendix G.
-IV.2.1 POPS-ensemble prediction of test errors
-For each test point, it is simple to generate an ‘envelope’, or max/min range, from the POPS-ensemble.
-We find this envelope provides a highly informative bound on test errors for large $N/P$.
-The top left of figure 3 is a parity plot for the minimum loss model, showing around 93% coverage, or 7% envelope violation.
-The violation rate drops with $N/P$, as shown on the bottom of figure 3 for a range of $P$ and $N/P$ and two simulation engines.
-Whilst the envelope bounds are highly informative, it is also desirable to predict test error distributions.
-As discussed above, the limitation of our POPS-ensemble is that the weighting of ensemble members is unlikely to be optimal, whilst the max/min range can be expected to be robust.
-As a result, a conservative estimate of the test error can be generated by uniformly resampling the max/min bounds for each test point to generate predicted error histograms.
-As shown in the top (right) of figure 3, this prediction from the POPS-ensemble results in excellent agreement with the test error ensemble, as confirmed across multiple engines in appendix B.
-Whilst a single (re)sample per test point is sufficient, it is possible to effectively bootstrap error histograms through repeated resampling.
-IV.2.2 Resampled POPS (rPOPS) ensemble
-The parameter vectors of the POPS-ensemble across all simulation engines studied are observed to be dense across their simplex, or hypercube of the max/min range per parameter.
-Inspired by the success of the uniform resampling approach used to produce test error predictions of the POPS ensemble in figure 3, we have also investigated the effect of uniformly resampling this ‘POPS-simplex’, a process we name as ‘resampled’ POPS, or rPOPS.
-We emphasize that the rPOPS approach makes a strong assumption that all parameter vectors inside the POPS- simplex are close to valid POPS-ensemble members.
-We thus include these results as a motivation for further study; in particular, whilst the $\mathcal{O}(N)$ cost of the POPS-ensemble remains highly managable, the cost of the rPOPS-ensemble can be tailored through the number of resamples, offering large efficiencies as $N/P\to\infty$ when producing envelopes and test error distributions.
-As shown in figure 3, the rPOPS results are excellent, giving tight agreement with test error distributions and only slightly more conservative errors bounds, a feature which is again robust across simulation engines (appendix G).
-A detailed investigation of the POPS-simplex and theoretical analysis of the rPOPS-ensemble is left as a topic for future study.
-We conclude with an application to challenging high-dimensional $P=\mathcal{O}(1000)$ datasets from atomic machine learning.
-Figure 3: Top: test error predictions for a misspecified linear surrogate model (P=20) on a cubic simulation engine.
-Left: parity plot for the minimum loss model, along with envelope bounds from the POPS-ensemble and rPOPS-ensemble (see text), showing 93-98% coverage at $N/P=750$.
-Right: Test error histogram, along with predicted test errors from the POPS- and rPOPS-ensemble, as explained in the text.
-Below: POPS-ensemble envelope violation rates with $P$ and $N/P$ for two simulation engines.
-IV.3 Interatomic potentials for material simulations
-Figure 4: Characteristics of the ensemble method for the application to interatomic potentials.
-Left: Parity plot for energy and force predictions for $N/P=50$.
-The errorbars reflect the min/max envelope of the POPS-ensemble.
-Right: cumulative distribution of the ratio of the residual of the minimum-loss model to the envelope width (max/min range per test point) for different conditions.
+However, all conclusions remain robust to varying the precise nature of data generation and the classes of engine considered Swinburne and Perez (2024).
+As shown, the $\mathcal{O}(P)$ hypercube ansatz $\pi^{*}_{\mathcal{H}}$ provides an excellent prediction and bounding, of the test error distribution whilst the $\mathcal{O}(N)$ ensemble $\pi^{*}_{E}$ underestimates moments of the error distribution and the max/min envelope bound.
+Importantly, the envelope violation rate of the hypercube ansatz $\pi^{*}_{\mathcal{H}}$ drops with $N/P$, as shown in figure 2 for a range of $P$ and $N/P$.
+The simultaneous prediction and bounding of test errors is highly valuable for surrogate model deployment, allowing not only the prediction of expected errors but a robust assessment of worst case error scenarios.
+We conclude with an application of the hypercube ansatz to challenging high-dimensional $P=\mathcal{O}(1000)$ datasets from atomic machine learning.
+Figure 2: Test errors of a misspecified linear surrogate model on a cubic simulation engine.
+Left: test error histogram at $P=20$, $N/P=100$ for the minimum loss model (black) and predictions from the POPS-ensemble $\pi^{*}_{E}$ (orange) and POPS-hypercube $\pi^{*}_{\mathcal{H}}$ (green) ansatz.
+MAE: mean absolute error relative to the minimum loss solution.
+EV: envelope violation, points lying outside of the max/min bound.
+Right: Probability of envelope violation for the $\pi^{*}_{\mathcal{H}}$ ansatz with $P$ and $N/P$.
+III.3 High-dimensional models for atomic simulations
+Figure 3: Application to interatomic potentials.
+Left: test error histogram at $P=1596$, $N/P=81$ for the minimum loss model (black) and predictions from the POPS-hypercube $\pi^{*}_{\mathcal{H}}$ (green) ansatz.
+MAE: mean absolute error relative to the minimum loss solution.
+EV: envelope violation, points lying outside of the max/min bound.
+Right: Probability of envelope violation for the $\pi^{*}_{\mathcal{H}}$ ansatz with $N/P$.
 Atomic-scale simulations of materials were traditionally broadly classified into two types:
@@
 Estimating the errors incurred by such ML models is highly sought after to assess the robustness of conclusions drawn from the simulations (Wen and Tadmor, 2020; Li et al., 2018; Gabriel et al., 2021; Wan et al., 2021) .
-In the following, training data was generated using an information-theoretic approach (Karabin and Perez, 2020; Montes de Oca Zapiain et al., 2022).
+In the following, a linear surrogate model $V_{i}={\bm{\Theta}}\cdot{\bf F}_{i}$ was used to predict per-atom contributions.
+Numerical experiments were carried out by randomly sub-selecting $N$ training points (either energies or gradient components), and by generating a discrete ensemble of $N$ pointwise-optimal models according to Eq.
+20, up to a maximal value of $N=129,853$.
+The properties of the ensemble are then evaluated on a complementary set of 10,000 hold-out testing points.
+Error analysis is carried out in the context of the POPS-hypercube ansatz.
+Even on the full dataset, the analysis requires on the order of 20 minutes on 1 CPU core.
+It is therefore extremely lightweight in addition to being easily parallelizable, enabling its use for extremely large datasets.
+The results show that the envelope violation on test data varies between $2\%$ and $0.2\%$ for $N/P=0.06$ and $33$, respectively, in qualitative agreement with the results shown in figure 2.
+This demonstrates that worst-case model errors are extremely well captured by the POPS-hypercube ensemble.
+As reported in the left panel of figure 3, these excellent bounds do not trivially result from the resampled error distribution being overly pessimistic, as the distribution of errors induced by a uniform parameter resampling within the POPS-hypercube provides an excellent predictor of the actual error distribution with respect to the minimum-loss model.
+The MAE estimated from the resampled models exceeds the measured value by only 23%, while providing a very good description of the overall error distribution.
+The information generated by this simple and efficient approach will be invaluable to estimate the uncertainties on a range of different physical quantities inferred from atomistic simulations.
+IV Conclusion
+In spite of its ubiquity in practical applications, misspecification is often ignored in Bayesian regression approaches based on expected loss minimization, which can lead to erroneous conclusions regarding parameter uncertainty quantification, surrogate model selection, and error propagation.
+In this work, we show that the important near-deterministic underparameterized regime is amenable to approximation by two formally simple ansatz that exploits the concentration of the cross-entropy minimizing ensemble onto pointwise optimal parameter sets (POPS).
+For linear regression problems, this ensemble can be obtained extremely efficiently using rank-one perturbations of the expected loss solution, from which our optimally weighted POPS-ensemble or POPS-bounding hypercube ansatz can be efficiently obtained.
+Our POPS-constrained ansatz are shown to produce excellent parameter distributions which give accurate prediction of test error distributions and highly informative worst-case bounds on model predictions, in the important misspecified, near-deterministic regime where existing methods fail.
+Synthetic test data
+All synthetic data was generated with Python/JAX code, to be made available following peer review.
+Atomic simulation data
+Atomic training data was generated using an information-theoretic approach (Karabin and Perez, 2020; Montes de Oca Zapiain et al., 2022).
 10,000 atomistic configurations containing between 2 and 32 tungsten atoms were generated and characterized using Density Functional Theory (Kohn and Sham, 1965), yielding 167,922 training data points in total (10,000 energies and 157,922 gradient components).
 Per-atom energies are expressed as a linear combination of $P=1596$ so-called bispectrum components within the ‘Quadratic’ Spectral Neighborhood Analysis Potential (SNAP) formalism (Wood and Thompson, 2018), which can be expressed as a linear surrogate model $V_{i}={\bm{\Theta}}\cdot{\bf F}_{i}(\{{\bf x}_{j}|r_{ij}<r_{c}\})$.
-Numerical experiments were carried out by randomly sub-selecting $N$ training points (either energies or gradient components), and by generating a discrete ensemble of $N$ pointwise-optimal models according to Eq.
-23.
-The properties of the ensemble are then evaluated on a complementary set of 1,000 testing points.
-The random test/train partition is repeated 10 times for each value of $N/P$.
-The results show that the envelope violation on test data varies between $40\%$ and $10\%$ for $N/P=1$ and $50$, respectively, in qualitative agreement with the results shown in figure 3, demonstrating that worst-case model errors are increasingly well captured by the ensemble envelope.
-Indeed, the scaling is actually superior to the medium test examples.
-As seen in these examples, the envelope induced by the model ensemble yields a useful estimate of the actual errors, e.g., as incurred by the ordinary minimum-loss model, as reported in figure 4.
-For example, the mean ratio of the residual of the minimum-loss model to the envelope width (max/min range) is around $0.45$ at $N/P=1$, i.e, the envelope width overestimates the actual minimum-loss residual only by a factor of around 2 on average.
-This figure drops to about $0.25$ at $N/P=50$, following the broadening of the envelopes that occurs as rare ”outliers” that are particularly difficult to capture within the model class are gradually added to the training set.
-This was empirically verified by sampling ensembles containing only $P$ models, independently of the size of the training set: in this case, the mean ratio of the residual of the minimum-loss model to the envelope width remains around $0.45$ (c.f., figure 4).
-This numerical experiment indicates that the width of the model envelope predictions provides reliable worst-case confidence intervals as well as representative (albeit somewhat conservative) estimates of the actual errors at testing points.
-V Conclusion
-stIn spite of its ubiquity in practical applications, misspecification is often ignored in Bayesian regression approaches based on expected loss minimization, which can lead to erroneous conclusions regarding parameter uncertainty quantification, surrogate model selection, and error propagation.
-In this work, we show that the important near-deterministic underparameterized regime is amenable to approximation by a formally simple discrete ensemble ansatz that exploits the concentration of the cross-entropy minimizing ensemble onto pointwise optimal parameter sets.
-For linear regression problems, this ensemble can be obtained extremely efficiently using rank-one perturbations of the expected loss solution.
-POPS-constrained loss-minimizing ensembles are shown to produce excellent parameter distributions, worst-case bounds on model predictions, and approximations of the pointwise errors.
-Software and Data
-Will be released after peer review
 Acknowledgements
-This work was initiated during a visit at the Institute for Pure and Applied Mathematics (IPAM) at the University of California in Los Angeles (UCLA).
+This work was initiated during a visit at the Institute for Pure and Applied Mathematics at the University of California, Los Angeles (supported by National Science Foundation (NSF) grant DMS-1925919) and continued at the Institute for Mathematical and Statistical Innovation, University of Chicago (supported by the NSF grant DMS-1929348).
 Their hospitality is graciously acknowledged.
-TDS gratefully acknowledges support from ANR grant ANR-19-CE46-0006-1, IDRIS allocations A0120913455 and an Emergence@INP grant from the CNRS.
+TDS gratefully acknowledges support from ANR grants ANR-19-CE46-0006-1 and ANR-23-CE46-0006-1, IDRIS allocation A0120913455 and an Emergence@INP grant from the CNRS.
 D.P.
@@
 We refer the reader to (Germain et al., 2016) and subsequent works (Shalaeva et al., 2020; Masegosa, 2020) for further discussion.
+For completeness, we recall here the bound used in the main text:
+${\mathcal{L}}[\pi]\leq\int_{\mathcal{P}}\hat{L}_{N}({\bm{\Theta}})+\frac{1}{N}\ln\left|\frac{\pi}{\pi_{0}}\right|{\rm d}\pi({\bm{\Theta}})-\frac{\ln|\xi|}{N}+C_{0},$ (21)
 Appendix B Model selection criteria for deterministic models
@@
 We note that the value of the PAC-Bayes loss bound (21) satifies $BIC=2N\hat{\mathcal{L}}_{N}[\pi_{{\mathcal{L}},N}]-2C_{0}$, with $C_{0}$ model independent, as discussed in (Germain et al., 2016).
-In standard derivations ${\bm{\Sigma}}_{\mathcal{Y}}$ is set to the variational maximum likelihood solution ${\bm{\Sigma}}_{\mathcal{Y}}={\bm{\Sigma}}^{*}_{\mathcal{L}}$, yielding the familar expression $BIC=P\ln N+N\ln\|{\bm{\Sigma}}^{*}_{\mathcal{L}}\|+\ln\|{\bm{\mathcal{I}}}^{*}_{\mathcal{L}}\|$, where model independent terms are neglected.
+In standard derivations ${\bm{\Sigma}}_{\mathcal{Y}}$ is set to the variational maximum likelihood solution ${\bm{\Sigma}}_{\mathcal{Y}}={\bm{\Sigma}}^{*}_{\mathcal{L}}$, yielding the familar expression $BIC=P\ln N+N\ln\|{\bm{\Sigma}}^{*}_{\mathcal{L}}\|+\ln\|{\boldsymbol{\mathcal{I}}}^{*}_{\mathcal{L}}\|$, where model independent terms are neglected.
 However, for deterministic regression problems, ${\bm{\Sigma}}_{\mathcal{Y}}$ is fixed.
 Neglecting model-independent terms, this gives an alternative BIC for deterministic models of
-$BIC=N{\rm Tr}({\bm{\Sigma}}^{*}_{\mathcal{E}}{\bm{\Sigma}}^{-1}_{\mathcal{Y}})+P\ln N+\ln\|{\bm{\mathcal{I}}}^{*}_{\mathcal{L}}\|.$ (25)
+$BIC=N{\rm Tr}({\bm{\Sigma}}^{*}_{\mathcal{E}}{\bm{\Sigma}}^{-1}_{\mathcal{Y}})+P\ln N+\ln\|{\boldsymbol{\mathcal{I}}}^{*}_{\mathcal{L}}\|.$ (23)
 Appendix C Summary of Laplace’s method
@@
 $\lim_{N\to\infty}\int_{a}^{b}\exp[Nf(x)]{\rm d}{x}=\sqrt{\frac{2\pi}{N}}\sqrt{\frac{{\rm d}^{2}f((x)}{{\rm d}x^{2}}}^{-1}\exp[N\max_{x\in[a,b]}f(x)].$ (24)
-Appendix D POPS-constrained variational minimization of ensemble ansatz
+Appendix D POPS-constrained ensemble ansatz
+Selecting some set of pointwise fits and $\mathcal{T}_{E}=\{{\bm{\Theta}}_{\mathbf{X}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})\}_{\mathcal{X}}$ and weights $\mathcal{W}_{E}=\{{\rm w}_{E}({\mathbf{X}})>0\}_{\mathcal{X}}$ such that $\langle{\rm w}_{E}({\mathbf{X}})\rangle_{\mathcal{X}}=1$, the covering constraint is satisfied by
+[…]
+For any choice of weights $\mathcal{W}_{E}$, an approximate optimal set ${\bm{\Theta}}^{*}_{\mathbf{X}}\in\mathcal{T}^{*}_{E}$ can be determined in a variational setting at finite $N$, using (25) in the PAC-Bayes loss bound (21) for the loss ${\mathcal{L}}[\pi_{E}]=\langle{\rm w}_{E}({\mathbf{X}})L({\bm{\Theta}}_{\mathbf{X}})\rangle_{\mathcal{X}}$.
+As $P/N\to 0$ we show in the sub-section below this reduces to the set of POPS-constrained loss minimizers
+${\bm{\Theta}}^{*}_{\mathbf{X}}\in\mathcal{T}^{*}_{E},\quad{\bm{\Theta}}^{*}_{\mathbf{X}}\equiv\arg\min_{{\bm{\Theta}}_{\mathbf{X}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})}L({\bm{\Theta}}_{\mathbf{X}})$ (26)
+The POPS-constrained loss-minimization ensures the fits will be as close as possible under the Fisher metric ${\boldsymbol{\mathcal{I}}}^{*}_{\mathcal{L}}$, such that ${\bm{\Theta}}^{*}_{\mathbf{X}}\to{\bm{\Theta}}^{*}_{\mathcal{L}}$ for specified models.
+Under misspecification, the generalization error of (25) using $\mathcal{T}^{*}_{E}$ has the $P/N\to 0$ limit
+${\mathcal{G}}_{0}[\pi_{E}]=-\left\langle\ln\left\langle{\rm w}_{E}({\mathbf{X}}^{\prime}){\rm m}_{\mathcal{E}}({\bm{\Theta}}^{*}_{{\mathbf{X}}^{\prime}},{\mathbf{X}})\right\rangle_{{\mathbf{X}}^{\prime}\in{\mathcal{X}}}\right\rangle_{{\mathbf{X}}\in{\mathcal{X}}}.$ (27)
+In general we expect a uniform weighting ${\rm w}_{E}({\mathbf{X}})=1$ to be sub-optimal for generalization as the loss-minimizing fits ${\bm{\Theta}}^{*}_{\mathbf{X}}$ will be too close to ${\bm{\Theta}}^{*}_{\mathcal{L}}$.
+To optimize ${\rm w}_{E}({\mathbf{X}})$ we define a density $\rho_{{\mathbf{X}}}^{*}({\mathbf{X}}^{\prime})={\rm m}_{\mathcal{E}}({\bm{\Theta}}^{*}_{{\mathbf{X}}^{\prime}},{\mathbf{X}})/\langle{\rm m}_{\mathcal{E}}({\bm{\Theta}}^{*}_{{\mathbf{X}}^{\prime}},{\mathbf{X}})\rangle_{{\mathbf{X}}^{\prime}\in{\mathcal{X}}}$ of pointwise fits ${\bm{\Theta}}^{*}_{{\mathbf{X}}^{\prime}}$ satisfying $|{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}}^{*}_{{\mathbf{X}}^{\prime}})-{\mathcal{E}}({\mathbf{X}})|=\mathcal{O}(\epsilon)$, then use Jensen’s inequality to write the ”loss-like” bound
+${\mathcal{G}}_{0}[\pi_{E}]$ $={\mathcal{G}}_{U}-\left\langle\ln\left\langle{\rm w}_{E}({\mathbf{X}}^{\prime})\rho_{{\mathbf{X}}}^{*}({\mathbf{X}}^{\prime})\right\rangle_{{\mathbf{X}}^{\prime}\in{\mathcal{X}}}\right\rangle_{{\mathbf{X}}\in{\mathcal{X}}}$
+$\leq{\mathcal{G}}_{U}-\left\langle\ln|{\rm w}_{E}({\mathbf{X}}^{\prime})|\left\langle\rho_{{\mathbf{X}}}^{*}({\mathbf{X}}^{\prime})\right\rangle_{{\mathbf{X}}\in{\mathcal{X}}}\right\rangle_{{\mathbf{X}}^{\prime}\in{\mathcal{X}}},$ (28)
+where ${\mathcal{G}}_{U}$ is the generalization error (27) with uniform weighting ${\rm w}_{E}({\mathbf{X}})=1$.
+Under the constraint $\langle{\rm w}_{E}({\mathbf{X}})\rangle_{\mathcal{X}}=1$ it is simple to show this upper bound is minimized by
+${\rm w}^{*}_{E}({\mathbf{X}})\in\mathcal{W}_{E}^{*},\quad{\rm w}^{*}_{E}({\mathbf{X}})=\lambda\left\langle\rho_{{\mathbf{X}}^{\prime}}^{*}({\mathbf{X}})\right\rangle_{{\mathbf{X}}^{\prime}\in{\mathcal{X}}}.$ (29)
+where $\lambda$ ensures $\langle{\rm w}^{*}_{E}({\mathbf{X}})\rangle_{\mathcal{X}}=1$.
+Evaluation requires at most $\mathcal{O}(N^{2})$ effort but only $\mathcal{O}(N)$ for storage and resampling.
+The final ensemble ansatz parameter posterior then reads
+$\pi^{*}_{E}({\bm{\Theta}})\equiv\langle{\rm w}^{*}_{E}({\mathbf{X}})\delta({\bm{\Theta}}-{\bm{\Theta}}^{*}_{{\mathbf{X}}})\rangle_{\mathcal{X}}..$ (30)
+Equation (30) is a variationally optimized ensemble ansatz for misspecified regression problems.
+The worst case of extreme misspecification, corresponding to non-intersecting POPS ${\mathcal{P}}_{\mathcal{M}}({\mathbf{X}}^{\prime})\cap{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})=\emptyset$, gives $\rho_{{\mathbf{X}}^{\prime}}^{*}({\mathbf{X}})=\delta_{{\mathbf{X}}{\mathbf{X}}^{\prime}}$ and thus uniform weights ${\rm w}^{*}_{E}({\mathbf{X}})=1$.
+Interestingly, it is simple to see we have uniform weighting at large $\epsilon$ as $\rho_{{\mathbf{X}}^{\prime}}^{*}({\mathbf{X}})$ is uniform.
+For real systems we find ${\rm w}^{*}_{E}({\mathbf{X}})$ is highly non-uniform and smallest for points near to ${\bm{\Theta}}^{*}_{\mathcal{L}}$, as expected.
+However, numerical test reveal moments of the predicted ensemble errors are still underestimated, as the ”loss like” bounds still bring all points too close to the minimum loss value, which motivates the hypercube approach described in the text.
+However, the max/min bounds of the ensemble, which are insensitive to our choice of ${\rm w}({\bf X})$, remain robust, as shown in figure 1.
+D.1 Variational optimization
 For finite $N/P$, we have access to training data ${\mathcal{D}}_{N}=\{{\mathbf{Y}}_{i},{\mathbf{X}}_{i}\}_{i=1}^{N}$ to give averages over $\langle\dots\rangle_{N}$, with some set of pointwise fits $\mathcal{T}_{E}=\{{\bm{\Theta}}_{i}\}_{i=1}^{N}$, ${\bm{\Theta}}_{i}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}}_{i})$.
@@
 Aside from the POPS constrained variation, this is a standard variational result which in principle allows for variational minimization of $\pi_{E}({\bm{\Theta}})$ at finite $N$ for both linear and non-linear models.
-An alternative POPS-constrained variational approach for linear regression at finite $N$ is provided in appendix E.
+An alternative POPS-constrained variational approach for linear regression at finite $N$ is provided in appendix LABEL:app:variational-gaussian.
 In the limit $\sigma\to 0$ we make the approximation that, for ${\bm{\Theta}}-{\bm{\Theta}}_{i}=\delta{\bm{\Theta}}$, $\pi_{E}({\bm{\Theta}}_{i}+\delta{\bm{\Theta}})\simeq(1/N)\mathcal{N}(\delta{\bm{\Theta}}|{\bf 0},\sigma^{2}\mathbb{I})$.
@@
 $\int\nabla_{\bm{\Theta}}\ln\left|\frac{\pi_{E}({\bm{\Theta}})}{\pi_{0}({\bm{\Theta}})}\right|\mathcal{N}({\bm{\Theta}}|{\bm{\Theta}}_{i},\sigma^{2}\mathbb{I}){\rm d}{\bm{\Theta}}=\int\left(\frac{{\bm{\Theta}}-{\bm{\Theta}}_{i}}{\sigma^{2}}+{\bm{\Sigma}}_{0}^{-1}{\bm{\Theta}}\right)\mathcal{N}({\bm{\Theta}}|{\bm{\Theta}}_{i},\sigma^{2}\mathbb{I}){\rm d}{\bm{\Theta}}\to{\bm{\Sigma}}_{0}^{-1}{\bm{\Theta}}_{i},\quad\sigma\to 0.$ (35)
-In the present work we are primarily concerned with the underparametrized, or large data limit $P/N\to 0$, where the influence of the prior distribution is negligible, as demonstrated in section II.3.
+In the present work we are primarily concerned with the underparametrized, or large data limit $P/N\to 0$, where the influence of the prior distribution is negligible, as demonstrated in the main text.
 In this regime, for small $\sigma$ the variation of the upper bound (21) has a limiting form of
@@
 ${\bm{\Theta}}^{*}_{i}\equiv\arg\min_{{\bm{\Theta}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}}_{i})}L({\bm{\Theta}}),$ (37)
-in agreement with the result given in equation (21).
-Appendix E POPS-constrained variational ansatz for PAC-Bayesian linear regression
-The POPS-constrained ensemble ansatz can also be used to incorporate both epistemic and misspecification parameter uncertainties in a variational scheme robust at small $N/P$, which we detail here for completeness.
-We assume a Gaussian posterior $\pi_{V}({\bm{\Theta}})=\mathcal{N}({\bm{\Theta}}|{\bm{\mu}}_{V},{\bm{\Sigma}}_{V})$, such that the ensemble ansatz is characterized by the ensemble mean ${\bm{\mu}}_{E}=\langle{\bm{\Theta}}^{*}_{\mathbf{X}}\rangle$ and covariance ${\bm{\Sigma}}_{E}$.
-A first approximation simply combines the covariance from minimization of the PAC-Bayes loss bound (7), ${\bm{\Sigma}}^{*}_{\mathcal{L}}=\left[{\bm{\Sigma}}_{\mathcal{Y}}{\bm{\Sigma}}^{-1}_{0}+N\langle{\mathbf{F}}({\mathbf{X}})[{\mathbf{F}}({\mathbf{X}})]^{\top}\rangle_{\mathcal{X}}\right]^{-1}$, with the ensemble covariance ${\bm{\Sigma}}_{E}$.
-This yield a posterior covariance ${\bm{\Sigma}}^{*}_{\mathcal{L}}+{\bm{\Sigma}}_{E}$, which remains robust at finite $N/P$ and limits to ${\bm{\Sigma}}_{E}$ as $N/P\to\infty$.
-However, the full variational scheme has comparable complexity to Bayesian ridge regression for linear models (Tipping, 2001).
-Our variational ansatz replaces ${\bm{\Theta}}^{*}_{\mathcal{L}}$ in the pointwise fits (23) with a variational parameter ${\bm{\Theta}}_{V}$, giving a variational posterior mean and covariance of
-${\bm{\mu}}_{V}({\bm{\Theta}}_{V})\equiv\left\langle{\bf v}_{\mathbf{X}}\right\rangle_{\mathcal{X}}+\left\langle{\bf W}_{\mathbf{X}}\right\rangle_{\mathcal{X}}{\bm{\Theta}}_{V},\quad{\bm{\Sigma}}_{V}({\bm{\Theta}}_{V})\equiv\left\langle\left[\delta{\bf v}_{\mathbf{X}}+\delta{\bf W}_{\mathbf{X}}{\bm{\Theta}}_{V}\right]\left[\delta{\bf v}_{\mathbf{X}}+\delta{\bf W}_{\mathbf{X}}{\bm{\Theta}}_{V}\right]^{\top}\right\rangle_{\mathcal{X}},$ (34)
-where ${\bf v}_{\mathbf{X}},{\bf W}_{\mathbf{X}}$ are derived from (23) such that ${\bm{\mu}}_{V}({\bm{\Theta}}^{*}_{\mathcal{L}})={\bm{\mu}}_{E}$ and ${\bm{\Sigma}}_{V}({\bm{\Theta}}^{*}_{\mathcal{L}})={\bm{\Sigma}}_{E}$, reading
-${\bf v}_{\mathbf{X}}\equiv{\bf A}\frac{{\mathbf{F}}({\mathbf{X}}){\mathcal{E}}({\mathbf{X}})}{{\rm h}({\mathbf{X}})},\quad{\bf W}_{\mathbf{X}}\equiv\mathbb{I}-{\bf A}\frac{{\mathbf{F}}({\mathbf{X}}){\mathbf{F}}^{\top}({\mathbf{X}})}{{\rm h}({\mathbf{X}})},\quad\delta{\bf v}_{\mathbf{X}}\equiv{\bf v}_{\mathbf{X}}-\left\langle{\bf v}_{\mathbf{X}}\right\rangle_{\mathcal{X}},\quad\delta{\bf W}_{\mathbf{X}}\equiv{\bf W}_{\mathbf{X}}-\left\langle{\bf W}_{\mathbf{X}}\right\rangle_{\mathcal{X}}.$
-This gives a variational distribution which can be used in the PAC-Bayes loss bound ^$\hat{{\mathcal{L}}}_{N}[\pi]$ (7)
-$\pi_{V}({\bm{\Theta}}|{\bm{\Theta}}_{V})\equiv\mathcal{N}({\bm{\Theta}}|{\bm{\mu}}_{V}({\bm{\Theta}}_{V}),{\bm{\Sigma}}_{V}({\bm{\Theta}}_{V})),\quad{\bm{\Theta}}_{V}^{*}\equiv\arg\min_{{\bm{\Theta}}_{V}\in{\mathcal{P}}}\hat{{\mathcal{L}}}_{N}[\pi_{V}(.|{\bm{\Theta}}_{V})].$ (35)
-Using the empirical loss $\hat{L}_{N}({\bm{\Theta}})$, quadratic in ${\bm{\Theta}}$ for linear models, the variational minimum satisfies
-$\left\langle{\bf W}_{\mathbf{X}}\right\rangle_{\mathcal{X}}[N{\bm{\nabla}}_{\bm{\Theta}}\hat{L}_{N}({\bm{\Theta}}^{*}_{V})+{\bm{\Sigma}}_{0}^{-1}{\bm{\mu}}_{V}({\bm{\Theta}}^{*}_{V})]+\left[N{\bf A}^{-1}+{\bm{\Sigma}}_{0}^{-1}+{\bm{\Sigma}}^{-1}_{V}({\bm{\Theta}}^{*}_{V})\right]:{\bm{\Sigma}}^{\prime}_{V}({\bm{\Theta}}^{*}_{V})={\bf 0},$ (36)
-where :::: denotes a contraction over the first two indicies of a third-order tensor ${\bm{\Sigma}}^{\prime}_{V}({\bm{\Theta}})$ linear in ${\bm{\Theta}}$:
-${\bm{\Sigma}}^{\prime}_{V}({\bm{\Theta}})=\left\langle\delta{\bf W}_{\mathbf{X}}\delta{\bf v}^{\top}_{\mathbf{X}}+\delta{\bf v}_{\mathbf{X}}\delta{\bf W}_{\mathbf{X}}+\delta{\bf W}_{\mathbf{X}}{\bm{\Theta}}\delta{\bf W}_{\mathbf{X}}+\delta{\bf W}_{\mathbf{X}}\delta{\bf W}_{\mathbf{X}}{\bm{\Theta}}\right\rangle_{\mathcal{X}}.$ (37)
-As for Bayesian ridge regression, the only non-linearity in (36) is the term ${\bm{\Sigma}}^{-1}_{V}({\bm{\Theta}}):{\bm{\Sigma}}^{\prime}_{V}({\bm{\Theta}}^{*}_{V})$, meaning any implementation can use existing iterative schemes e.g.
-(Tipping, 2001),(Wipf and Nagarajan, 2007).
-A full study of the low data regime is left for future work.
-Appendix F Implementation of the ensemble ansatz in Python
-The following python code, where A.T is the transpose of A and we assume a uniform weighting ${\rm w}_{E}({\bf X})=1$:
+in agreement with the result given in the main text.
+Appendix E Implementation of the ansatz in Python
+The hypercube projector ${\bf V}\in\mathbb{R}^{R\times P}$ and bounds ${\bf B}\in\mathbb{R}^{R\times 2}$ can be produced with the following python code, where A.T is the transpose of A:
 1import numpy as np
-2def pointwise_optima(X,y):
+2def hypercube(X,y,prior,thresh=1e-8):
 3 # X.shape = (N,P), y.shape = (N,)
-4    # Inverse of A
-5    A_i = np.dot(X.T,X) +
+4 # Inverse of A with prior matrix
+5 A_i = np.dot(X.T,X) + prior / y.size
 6 # Minimum loss solution
@@
 14 T = np.multiply(X_dot_A, e/h )
-15    return T
-An ensemble of test predictions can then be made simply by contracting with a constant tensor, a process which can easily be vectorized/parallelized.
-Appendix G Additional plots of test error prediction
-The plots below are identical in presentation to the top row of figure 3, but for the quadratic and random linear simulation engines described in the main text.
+15 # Find hypercube projector
+16 nu,V = np.linalg.eigh(T@T.T)
+17 nu_sel = nu<nu.max()*thresh
+18 V = V[nu_sel,:] # shape = (R,P)
+19 # bounds
+20 B = np.array([(T@V.T).max(0),(T@V.T).min(0)]).T
+21 return V,B
+To sample the ensemble we use uniform random numbers
+1import numpy as np
+2def resample(V,B,n=20,seed=123):
+3 np.random.seed(seed)
+4 u = np.random.uniform(\
+5 min=B[:,0],
+6 max=B[:,1],
+7 size=(n,V.shape[0]))
+8 return u@V
+Appendix F Additional plots of test error prediction
+F.1 Low dimensional models
+We applied the same procedure as for the polynomical example to a model system with $P=6$ independent features to simulate more realistic datasets.
+As before, the numerical minimization routine could only treat a small range of $\sigma$ values.
+Figure 4 shows the resulting parameter distributions.
+At large $\sigma=2$ the minimizer quickly concentrates all ensemble members on the minimum loss solution; at $\sigma=1$ the ensemble distribution is notably sharper than the POPS-ensemble $\pi^{*}_{E}$, whilst for low sigma $\sigma=1/2$ the minimization only has marginal stability, but becomes closer to the POPS-bounding hypercube $\pi^{*}_{\mathcal{H}}$.
+In addition, a minimizer initialized with the POPS-ensemble did not change appreciably.
+These albeit simple results give strong evidence that our ansatz is indeed a stable and highly efficient means to find approximate minimizers of ${\mathcal{G}}$ that would otherwise be intractable.
+Figure 4: Regression of a linear model onto a cubic ”simulation engine” with $P=6$ independent features, trained on $N=500$ points.
+The panels show the cumulative distribution of each parameter across the $N$-member ensemble.
+Black, gray : ensemble and hypercube ansatz $\pi^{*}_{E}$ and $\pi^{*}_{\mathcal{H}}$.
+Blue, orange and green : numerically optimized $\mathcal{G}_{E}$ for a uniformly weighted $N$-ensemble, regularized with $\Sigma_{\mathcal{Y}}=\sigma\Sigma^{*}_{\mathcal{L}}$, for $\sigma=2,1,1/2$.
+As in Figure 1 in the main text, lower $\sigma$ gave numerical instabilities.
+F.2 Medium dimensional models
+The plots below are identical in presentation to the top row of figures in the main text, but for the quadratic and random linear simulation engines described in the main text.
 Figure 5: Test error prediction for medium dimensional problems.
-From left: Observed test errors for a misspecified linear surrogate model (P=20) on a quadratic simulation engine, along with predicted test errors from the POPS-ensemble.
-Parity plots for the minimum loss model, along with envelope bounds from the POPS-ensemble.
+Left: Observed test errors for a misspecified linear surrogate model (P=15) on a quadratic (top) and random linear (bottom) simulation engine, along with predicted test errors from the POPS-ensemble.
+Right: Parity plots for the minimum loss model, along with envelope bounds from the POPS-ensemble and hypercube.

### v5

@@
 Bayesian regression determines model parameters by minimizing the expected loss, an upper bound to the true generalization error.
-However, the loss ignores misspecification, where models are imperfect.
-Parameter uncertainties from Bayesian regression are thus significantly underestimated and vanish in the large data limit.
-This is particularly problematic when building models of low-noise, or near-deterministic, calculations, as the main source of uncertainty is neglected.
-We analyze the generalization error of misspecified, near-deterministic surrogate models, a regime of broad relevance in science and engineering.
-We show posterior distributions must cover every training point to avoid a divergent generalization error and design an ansatz that respects this constraint, which for linear models incurs minimal overhead.
-This is demonstrated on model problems before application to thousand dimensional datasets in atomistic machine learning.
-Our efficient misspecification-aware scheme gives accurate prediction and bounding of test errors where existing schemes fail, allowing this important source of uncertainty to be incorporated in computational workflows.
-Surrogate models are widely used across science and engineering to efficiently approximate the action of computationally intensive simulation engines  Alizadeh et al.
+However, this loss ignores model form error, or misspecification, meaning parameter uncertainties are significantly underestimated and vanish in the large data limit.
+As misspecification is the main source of uncertainty for surrogate models of low-noise calculations, such as those arising in atomistic simulation, predictive uncertainties are systematically underestimated.
+We analyze the true generalization error of misspecified, near-deterministic surrogate models, a regime of broad relevance in science and engineering.
+We show that posterior parameter distributions must cover every training point to avoid a divergence in the generalization error and design a compatible ansatz which incurs minimal overhead for linear models.
+The approach is demonstrated on model problems before application to thousand-dimensional datasets in atomistic machine learning.
+Our efficient misspecification-aware scheme gives accurate prediction and bounding of test errors in terms of parameter uncertainties, allowing this important source of uncertainty to be incorporated in multi-scale computational workflows.
+Surrogate models are widely used across science and engineering to efficiently approximate the output of computationally intensive simulation engines  Alizadeh et al.
 (2020); Deringer et al.
@@
 In a broad range of applications, the inference problem shares three key characteristics:
-a) Simulation engines are near deterministic: outputs have vanishing aleatoric uncertainty, e.g. the atomic energy $V({\bf X})$ is deterministic to atomic positions ${\bf X}$.
-b) Surrogate models are misspecified: no single choice of the $P$ parameters can match all $N$ observations, meaning parameters are intrinsically uncertain.
-c) Surrogate models are underparametrized: large quantities of training data are available, such that standard epistemic uncertainty estimates from Bayesian inference vanish, typically corresponding to the regime $P/N\ll 1$.
+  1. a)
+    Simulation engines are near deterministic: outputs have vanishing aleatoric uncertainty, e.g.
+the atomic energy $V({\bf X})$ from some quantum chemistry calculation is a near-deterministic function of atomic positions ${\bf X}$, with very weak uncertainty due to e.g.
+convergence of the electronic minimization scheme.
+  2. b)
+    Surrogate models are misspecified: no single choice of the $P$ parameters can match all $N$ observations, meaning parameters are intrinsically uncertain.
+  3. c)
+    Surrogate models are underparametrized: large quantities of training data are available, such that standard epistemic uncertainty estimates from Bayesian inference vanish, typically corresponding to the regime $P/N\ll 1$.
 Misspecification affects both finite capacity models and deep learning approaches with finite training resources  Lahlou et al.
@@
 (2021).
-In this paper, we analyze the generalization error of deterministic, misspecified and under-parametrized surrogate models, i.e.
+In this paper, we analyze the generalization error of near-deterministic, misspecified and under-parametrized surrogate models, i.e.
 those typically used to approximate simulation engines in science and engineering.
-As posterior distributions strictly become unbounded in the deterministic limit, this is only taken as a theoretical device to derive conditions that any posterior must obey to avoid a divergent generalization error.
+Our central application is to the problem of fitting interatomic potentials for atomic simulationsGoryaeva et al.
+(2021); Montes de Oca Zapiain et al.
+(2022); Li et al.
+(2020); Wood and Thompson (2018), where the training data is generated by quantum chemistry calculations.
+However, our analysis applies to the fitting of any surrogate model to a near-deterministic ground truth.
+Posterior distributions strictly become unbounded as the aleatoric uncertainty vanishes, meaning the low-noise/near-deterministic limit is only taken as a theoretical device to derive conditions that any posterior distribution must obey to avoid a logarithmic divergence in the generalization error of that distribution.
 We then design an ansatz that respect this condition and admits a robust learning scheme at finite $N$, with standard treatment of epistemic ($N$-dependent) errors, reducing to a constrained loss minimization problem for large $N/P$.
-Our final misspecification-aware parameter posterior has finite uncertainty even as $N/P\to\infty$ with a provably lower generalization error than the Gibbs posterior of Bayesian inference.
+Our final misspecification-aware parameter posterior has finite uncertainty even as $N/P\to\infty$, whilst the generalization error of the posterior from Bayesian inference diverges.
+The approach thus captures misspecification uncertainties missed by standard Bayesian regression schemes.
+Our implementation for linear models is a robust and to our knowledge unique means to estimate misspecification uncertainties, and thus is the optimal available choice when fitting high-dimensional linear models to low-noise data.
 Our main contributions are:
-a) We define pointwise optimal parameter sets (POPS) for each training point, within which model predictions are exact.
+  1. a)
+    We define pointwise optimal parameter sets (POPS) for each training point, within which model predictions are exact.
 Parameter distributions must have mass in every POPS to avoid a divergent generalization error.
-b) We use the ensemble of loss minimizers from each POPS to design two ansatz posteriors, a reweighted $\mathcal{O}(N)$ ensemble and a bounding $\mathcal{O}(P)$ hypercube, both of which respect POPS occupancy and give a finite generalization error, lower than the Gibbs posterior from inference.
-c) For linear models, our ansatz can be efficiently evaluated via rank-one updates to a leverage-weighted loss minimizer.
-This gives an efficient scheme to incorporate misspecification into linear regression with minimal computational overhead.
-d) We test the approach on synthetic datasets, showing how our scheme provides robust bounds on test errors.
-This performance is maintained under application to thousand-dimensional datasets from atomic machine learning, untreatable with existing methods.
-.1 Related work
-Germain and coworkers Germain et al.
+  2. b)
+    We use the ensemble of loss minimizers from each POPS to design two ansatz posteriors, a reweighted $\mathcal{O}(N)$ ensemble and a bounding $\mathcal{O}(P)$ hypercube, both of which respect POPS occupancy and give a finite generalization error, lower than the posterior from Bayesian inference.
+  3. c)
+    For linear models, our ansatz can be efficiently evaluated via rank-one updates to a leverage-weighted loss minimizer.
+This gives an efficient scheme to incorporate misspecification into linear regression, which we show gives minimal computational overhead (approximately a factor of two in training) over Bayesian inference.
+  4. d)
+    A pip-installable Python implementation of our misspecification-aware POPS regression algorithm, following the Scikit-learn linear_model API Pedregosa et al.
+(2011), is provided at https://github.com/tomswinburne/POPS-Regression.git.
+  5. e)
+    We test the approach on synthetic datasets, showing how our scheme provides robust bounds on test errors.
+This performance is maintained under application to thousand-dimensional datasets for fitting machine learning interatomic potentialsMontes de Oca Zapiain et al.
+(2022); Li et al.
+(2020) and informatics modelsTynes et al.
+(2024); Ruddigkeit et al.
+(2012), untreatable with existing misspecification-aware methods.
+The remainder of this paper is organized as follows.
+In Section I, we review Bayesian regression and PAC-Bayes bounds, highlighting the limitations of standard approaches for misspecified models.
+Section II presents our analysis of the generalization error for near-deterministic surrogate models, introducing the concept of pointwise optimal parameter sets (POPS).
+In Section III, we develop our misspecification-aware ansatz, providing efficient implementations for linear models.
+Section IV demonstrates the effectiveness of our approach on both synthetic datasets and high-dimensional problems in atomic machine learning.
+Section V discusses our method in the context of existing uncertainty quantification and propagation schemes for atomistic simulation.
+Finally, Section VI summarizes our findings and highlights future directions for research in misspecification-aware regression.
+I Bayesian fitting of surrogate models
+This section summarizes known results from Bayesian analysis and PAC-Bayes inequalitiesAlquier (2021), focusing on the limit of interest for this paper: fitting a misspecified surrogate model in the case that the ground truth is (near-)deterministic, i.e.
+known to have vanishing aleatoric error.
+We provide a glossary of all terms in Appendix A
+I.1 Deterministic simulation engines
+A simulation engine ${\mathcal{E}}:{\mathcal{X}}\to{\mathcal{Y}}$ takes input ${\mathbf{X}}\in{\mathcal{X}}$ and returns ${\mathbf{Y}}\in{\mathcal{Y}}$.
+We denote the dimension of ${\mathcal{Y}}$ and ${\mathcal{X}}$ by $Y$ and $X$ respectively.
+We consider the regime where the aleatoric uncertainty of ${\mathcal{E}}$ is weak, such that the output ${\mathbf{Y}}$ follows a Gaussian distribution
+$\rho_{\mathcal{E}}({\mathbf{Y}}|{\mathbf{X}})=\mathcal{N}({\mathbf{Y}}|{\mathcal{E}}({\mathbf{X}}),{\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})),$ (1)
+which is fully defined by the output mean $\langle{\mathbf{Y}}|{\mathbf{X}}\rangle\equiv{\mathcal{E}}({\mathbf{X}})$ and covariance $\langle{\mathbf{Y}}{\mathbf{Y}}^{\top}|{\mathbf{X}}\rangle-{\mathcal{E}}({\mathbf{X}}){\mathcal{E}}^{\top}({\mathbf{X}})\equiv{\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})$, where $\|{\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})\|=\mathcal{O}(\epsilon^{2Y})$, and $\epsilon$ is assumed small.
+As the aleatoric error is vanishingly weak, we assume homoskedatiscity ${\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})={\bm{\Sigma}}_{\mathcal{Y}}$ for simplicity.
+We will later take the deterministic, underparametrised limit $\epsilon\to 0$, $N/P\to\infty$ assuming that $\epsilon=\mathcal{O}(\sqrt{P/N})$.
+With weights $\sum_{i=1}^{N}{\rm w}_{i}=1$, the training data ${\mathcal{D}}_{N}=\{{\mathbf{Y}}_{i},{\mathbf{X}}_{i}\}_{i=1}^{N}$ form an input distribution in ${\mathbf{X}}$ of
+$\rho_{N}({\mathbf{X}})\equiv\sum_{i}{\rm w}_{i}\delta({\mathbf{X}}-{\mathbf{X}}_{i}),\,\,\rho_{{\mathcal{D}}}({\mathbf{X}})\equiv\lim_{N\to\infty}\rho_{N}({\mathbf{X}}).$ (2)
+Learning requires that we can bound expectations over independent identically distributed (i.i.d.) samples ${\mathbf{X}},{\mathbf{Y}}\in{\mathcal{D}}$, where ${\mathcal{D}}$ is the set of all possible input-output pairs, corresponding to the limit of infinite training data,
+$\langle f({\mathbf{Y}},{\mathbf{X}})\rangle_{{\mathcal{D}}}\equiv\int_{\mathcal{D}}f({\mathbf{Y}},{\mathbf{X}})\rho_{\mathcal{E}}({\mathbf{Y}}|{\mathbf{X}})\rho_{{\mathcal{D}}}({\mathbf{X}}){\rm d}{\mathbf{X}}{\rm d}{\mathbf{Y}}.$ (3)
+For timeseries or other signal data the i.i.d.
+property is only satisfied for decorrelated segments ${\mathbf{X}}=\{{\bf x}_{t}\}_{t=1}^{T}$, meaning ${\bm{\Sigma}}_{\mathcal{Y}}$ should capture signal correlations as in generalized least squares regressionAlquier (2021); Germain et al.
+(2016); Masegosa (2020).
+I.2 Deterministic surrogate models
+A surrogate model ${\mathcal{M}}:{\mathcal{X}}\to{\mathcal{Y}}$ for ${\mathcal{E}}$ is defined by $P$ parameters ${\bm{\Theta}}\in{\mathcal{P}}$.
+Under a parameter distribution $\pi({\bm{\Theta}})$, the predicted distribution has the same form as (1), reading
+$\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})$ $=\mathcal{N}({\mathbf{Y}}|{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}}),{\bm{\Sigma}}_{\mathcal{Y}}),$ (4)
+$\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}})$ $=\int_{\mathcal{P}}\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}.$ (5)
+Unlike typical Bayesian schemesBhat and Kumar (2010); Von Toussaint (2011), we do not treat the covariance term ${\bm{\Sigma}}_{\mathcal{Y}}$ as fitting parameter but as an intrinsic property of ${\mathcal{E}}$.
+This has implications for model selection criteria, discussed in appendix B.
+I.3 The generalization error and expected loss
+The generalization error ${\mathcal{G}}[\pi]$ is simply the cross entropy of the predicted distribution $\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}})\rho_{N}({\mathbf{X}})$ to the observed distribution $\rho({\mathbf{Y}}|{\mathbf{X}})\rho_{N}({\mathbf{X}})$; omiting the constant term ${\mathcal{G}}_{\mathcal{E}}\equiv\left\langle\ln\left|\rho_{\mathcal{E}}({\mathbf{Y}}|{\mathbf{X}})\right|\right\rangle_{\mathcal{D}}$ we have
+${\mathcal{G}}[\pi]=-\left\langle\ln\left|\int_{\mathcal{P}}\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}\right|\right\rangle_{\mathcal{D}}.$ (6)
+In principle any learning scheme should aim to minimize the functional ${\mathcal{G}}[\pi]$, but in practice learning schemes target a tractable upper bound, the expected loss ${\mathcal{L}}[\pi]\geq{\mathcal{G}}[\pi]$, often known as the negative log likelihood.
+Using the Jensen inequality $-\langle\ln x\rangle\leq-\ln\langle x\rangle$, as used to derive Gibbs’ famous inequality for free energy estimationGibbs (1902), we find that
+$\mathcal{G}[\pi]\leq{\mathcal{L}}[\pi]=-\int_{\mathcal{P}}\left\langle\ln\left|\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\right|\right\rangle_{\mathcal{D}}\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}.$ (7)
+${\mathcal{L}}[\pi]$ is clearly minimized by a delta function $\delta({\bm{\Theta}}-{\bm{\Theta}}^{*}_{\mathcal{L}})$, where ${\bm{\Theta}}^{*}_{\mathcal{L}}\in\arg\min_{\mathcal{P}}\left\langle-\ln\left|\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\right|\right\rangle_{\mathcal{D}}$.
+As the delta function has formally zero width, the loss minimizing distribution thus has vanishing parameter uncertainty.
+The next section shows how this emerges from the inference posterior at finite $N/P$.
+I.4 PAC-Bayes estimation and connection to inference
+PAC-Bayes analysis Alquier (2021) provides concentration inequalities Hoeffding (1994) that bound expectations over ${\mathcal{D}}$ by empirical expectations over ${\mathcal{D}}_{N}$ with some probability.
+Whilst originally designed for classification problems, Germain and coworkers Germain et al.
 (2016) established an important connection between Bayesian inference and probability approximately correct (PAC-Bayes) analysis for regression problems with unbounded losses, which bounds true expectations by empirical likelihood expectations through concentration inequalities.
-This provides a robust learning scheme without test-train splitting, and gives a rationalization for Bayesian inference as a form of regularization.
+A short additional note on the PAC-Bayes framework for regression is provided in appendix LABEL:app:pac_constant.
+The PAC-Bayes framework provides a robust learning scheme without test-train splitting, and gives a rationalization for Bayesian inference as a form of regularization.
 In particular, minimizing the PAC-Bayes bound for the expected loss gives the familiar relation between posterior, prior and likelihood from Bayesian inference.
+This subsection adapts the results of Germain et al.
+Germain et al.
+(2016) and Masegosa et al.
+Masegosa (2020) in the near-deterministic limit of interest here.
+The central result ofGermain et al.
+(2016) is that the expected loss ${\mathcal{L}}[\pi]$ admits the PAC-Bayes upper bound, holding with probability $1-\xi$, of
+${\mathcal{L}}[\pi]\leq\int_{\mathcal{P}}\hat{L}_{N}({\bm{\Theta}})+\frac{1}{N}\ln\left|\frac{\pi}{\pi_{0}}\right|{\rm d}\pi({\bm{\Theta}})-\frac{\ln|\xi|}{N}+C_{0},$ (8)
+where $\hat{L}_{N}({\bm{\Theta}})=-\langle\ln|\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\rangle_{N}$ is the empirical loss over $\rho_{N}({\mathbf{X}})$, $\pi_{0}({\bm{\Theta}})$ is some prior and $C_{0}$ is a constant.
+The value of the constant $C_{0}$ requires certain assumptions on the distribution of the error between the empirical and true expected loss.
+Following (Germain et al., 2016), we assume that this error distribution is unbounded but sub-normal, i.e.
+all error fluctuations are bounded by a normal distribution of variance $s^{2}$, giving $C_{0}[\pi_{0}]<s^{2}/2$.
+We refer the reader to (Germain et al., 2016) and subsequent works (Shalaeva et al., 2020; Masegosa, 2020) for further discussion.
+The minimizer of the upper bound (8) is the posterior from Bayesian inference Germain et al.
+(2016), reading, with a parameter $\lambda$ to enforce normalization
+$\pi_{{\mathcal{L}},N}({\bm{\Theta}})=\pi_{0}({\bm{\Theta}})\exp(\lambda-N\hat{L}_{N}({\bm{\Theta}})).$ (9)
+As $N/P\to\infty$ the posterior is strongly peaked around a loss minimizer ${\bm{\Theta}}^{*}_{\mathcal{L}}\in\arg\min_{{\bm{\Theta}}\in{\mathcal{P}}}L({\bm{\Theta}})$.
+Defining ${\bm{\Sigma}}^{-1}_{0}$ as the curvature of $-\ln\pi_{0}({\bm{\Theta}})$ around ${\bm{\Theta}}^{*}_{\mathcal{L}}$, we use Laplace’s method (appendix C) as $N/P\to\infty$ to write
+[…]
+[…]
+$\lim_{N/P\to\infty}{\mathcal{G}}[\pi_{\mathcal{L}}]=\frac{1}{2}{\rm Tr}({\bm{\Sigma}^{*}_{\mathcal{L}}}{\bm{\Sigma}}_{\mathcal{Y}}^{-1})=\mathcal{O}(1/\epsilon^{2}).$ (11)
+${\bm{\Sigma}}^{*}_{\mathcal{L}}=\langle[{\mathcal{E}}({\mathbf{X}})-{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}}^{*}_{\mathcal{L}})][{\mathcal{E}}({\mathbf{X}})-{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}}^{*}_{\mathcal{L}})]^{\top}\rangle$ is the error covariance around ${\bm{\Theta}}^{*}_{\mathcal{L}}$ and we use ${\rm Tr}({\bm{\Sigma}}_{\mathcal{Y}})=\mathcal{O}(\epsilon^{2})$.
+Under misspecification it is thus clear that minimizing this upper bound is sub-optimal Masegosa (2020) as the generalization error diverges as $1/\epsilon^{2}$.
+In the next section we analyze the generalization error as $\epsilon\to 0$ to derive a condition for any $\pi({\bm{\Theta}})$ that ensures a finite generalization error.
+I.5 Related work on misspecification-aware Bayesian regression
 In recent years, multiple groups have used the PAC-Bayes framework to derive tighter bounds for the generalization error than that provided by the expected loss  Masegosa (2020); Germain et al.
@@
 (2022); Morningstar et al.
-(2022).
+(2022), to build a misspecified-aware regression scheme.
 However, these efforts focus on probabilistic regression settings, often with neural network models far from the underparametrized limit, meaning aleatoric, epistemic and misspecification errors must be considered jointly.
@@
 As various definitions for the distinction between epistemic and misspecification uncertainties exist Kato et al.
-(2022), we define misspecification as non-aleatoric uncertainties which survive in the underparametrized limit, where the minimizer of PAC-Bayes bounds for the expected loss, which coincides with the posterior from Bayesian inference, predict vanishing parameter uncertainties.
-This definition is made more precise in section I.4.
+(2022), we have defined (section I.4) misspecification as non-aleatoric uncertainties which survive in the underparametrized limit, where the minimizer of PAC-Bayes bounds for the expected loss, which coincides with the posterior from Bayesian inference, predicts vanishing parameter uncertainties.
 Our epistemic uncertainties thus coincide with standard estimates from loss minimization.
-I Bayesian surrogate models
-I.1 Deterministic simulation engines
-A simulation engine ${\mathcal{E}}:{\mathcal{X}}\to{\mathcal{Y}}$ takes input ${\mathbf{X}}\in{\mathcal{X}}$ and returns ${\mathbf{Y}}\in{\mathcal{Y}}$.
-We consider the regime where the aleatoric uncertainty of ${\mathcal{E}}$ is weak and can be described as a Gaussian distribution with mean $\langle{\mathbf{Y}}|{\mathbf{X}}\rangle\equiv{\mathcal{E}}({\mathbf{X}})$ and covariance $\langle{\mathbf{Y}}{\mathbf{Y}}^{\top}|{\mathbf{X}}\rangle-{\mathcal{E}}({\mathbf{X}}){\mathcal{E}}^{\top}({\mathbf{X}})\equiv{\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})$, where $\|{\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})\|=\mathcal{O}(\epsilon^{2Y})$, and $\epsilon$ is assumed small.
-The true conditional output distribution thus reads
-$\rho_{\mathcal{E}}({\mathbf{Y}}|{\mathbf{X}})=\mathcal{N}({\mathbf{Y}}|{\mathcal{E}}({\mathbf{X}}),{\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})).$ (1)
-We assume homoskedatiscity ${\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})={\bm{\Sigma}}_{\mathcal{Y}}$ of this weak aleatoric error to investigate the deterministic, underparametrised limit $\epsilon\to 0$, $N/P\to\infty$ for $\epsilon=\mathcal{O}(\sqrt{P/N})$.
-With weights $\sum_{i=1}^{N}{\rm w}_{i}=1$, the training configurations ${\mathcal{D}}_{N}=\{{\mathbf{Y}}_{i},{\mathbf{X}}_{i}\}_{i=1}^{N}$ form an input distribution in ${\mathbf{X}}$ of
-$\rho_{N}({\mathbf{X}})\equiv\sum_{i}{\rm w}_{i}\delta({\mathbf{X}}-{\mathbf{X}}_{i}),\,\,\rho_{{\mathcal{D}}}({\mathbf{X}})\equiv\lim_{N\to\infty}\rho_{N}({\mathbf{X}}).$ (2)
-Learning requires that we can bound expectations over independent identically distributed (i.i.d.) samples ${\mathbf{X}},{\mathbf{Y}}\in{\mathcal{D}}$
-$\langle f({\mathbf{Y}},{\mathbf{X}})\rangle_{{\mathcal{D}}}\equiv\int_{\mathcal{D}}f({\mathbf{Y}},{\mathbf{X}})\rho_{\mathcal{E}}({\mathbf{Y}}|{\mathbf{X}})\rho_{{\mathcal{D}}}({\mathbf{X}}){\rm d}{\mathbf{X}}{\rm d}{\mathbf{Y}}.$ (3)
-For timeseries or other signal data the i.i.d.
-property is only satisfied for decorrelated segments ${\mathbf{X}}=\{{\bf x}_{t}\}_{t=1}^{T}$, meaning ${\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})$ captures signal correlations as in generalized least squares regression.
-I.2 Deterministic surrogate models
-A surrogate model ${\mathcal{M}}:{\mathcal{X}}\to{\mathcal{Y}}$ for ${\mathcal{E}}$ is defined by $P$ parameters ${\bm{\Theta}}\in{\mathcal{P}}$.
-Under a parameter distribution $\pi({\bm{\Theta}})$, the predicted distribution has the same form as (1), reading
-$\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})$ $=\mathcal{N}({\mathbf{Y}}|{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}}),{\bm{\Sigma}}_{\mathcal{Y}}),$ (4)
-$\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}})$ $=\int_{\mathcal{P}}\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}.$ (5)
-Unlike typical Bayesian schemes, we do not treat the covariance term ${\bm{\Sigma}}_{\mathcal{Y}}$ as fitting parameter but as an intrinsic property of ${\mathcal{E}}$.
-This has implications for model selection criteria, discussed in the supplementary material (SM) Swinburne and Perez (2024).
-I.3 The generalization error and expected loss
-The generalization error ${\mathcal{G}}[\pi]$ is simply the cross entropy of the predicted distribution $\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}})\rho_{N}({\mathbf{X}})$ to the observed distribution $\rho({\mathbf{Y}}|{\mathbf{X}})\rho_{N}({\mathbf{X}})$; omiting the constant term ${\mathcal{G}}_{\mathcal{E}}\equiv\left\langle\ln\left|\rho_{\mathcal{E}}({\mathbf{Y}}|{\mathbf{X}})\right|\right\rangle_{\mathcal{D}}$ we have
-${\mathcal{G}}[\pi]=-\left\langle\ln\left|\int_{\mathcal{P}}\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}\right|\right\rangle_{\mathcal{D}}.$ (6)
-In principle any learning scheme should aim to minimize the non-convex functional ${\mathcal{G}}[\pi]$, but in practice learning schemes target a convex upper bound, the expected loss ${\mathcal{L}}[\pi]\geq{\mathcal{G}}[\pi]$, often known as the negative log likelihood.
-Using the Jensen inequality $-\langle\ln x\rangle\leq-\ln\langle x\rangle$ gives
-$\mathcal{G}[\pi]\leq{\mathcal{L}}[\pi]=-\int_{\mathcal{P}}\left\langle\ln\left|\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\right|\right\rangle_{\mathcal{D}}\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}.$ (7)
-${\mathcal{L}}[\pi]$ is clearly minimized by a delta function $\delta({\bm{\Theta}}-{\bm{\Theta}}^{*}_{\mathcal{L}})$, where ${\bm{\Theta}}^{*}_{\mathcal{L}}\in\arg\min_{\mathcal{P}}\left\langle-\ln\left|\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\right|\right\rangle_{\mathcal{D}}$, i.e.
-vanishing parameter uncertainty.
-The next section shows how this emerges from the inference posterior at finite $N/P$.
-I.4 PAC-Bayes estimation and connection to inference
-PAC-Bayes analysis Alquier (2021) provides concentration inequalities Hoeffding (1994) that bound expectations over ${\mathcal{D}}$ by empirical expectations over ${\mathcal{D}}_{N}$ with some probability.
-Whilst originally designed for classification problems, the approach has been extended to regression settings with unbounded losses Germain et al.
-(2016).
-As discussed in the SM Swinburne and Perez (2024), The expected loss ${\mathcal{L}}[\pi]$ admits the PAC-Bayes upper bound, holding with probability $1-\xi$, of
-${\mathcal{L}}[\pi]\leq\int_{\mathcal{P}}\hat{L}_{N}({\bm{\Theta}})+\frac{1}{N}\ln\left|\frac{\pi}{\pi_{0}}\right|{\rm d}\pi({\bm{\Theta}})-\frac{\ln|\xi|}{N}+C_{0},$ (8)
-where $\hat{L}_{N}({\bm{\Theta}})=-\langle\ln|\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\rangle_{N}$ is the empirical loss over $\rho_{N}({\mathbf{X}})$, $\pi_{0}({\bm{\Theta}})$ is some prior and $C_{0}$ is a constant Swinburne and Perez (2024).
-The minimizer of the upper bound (21) is the posterior from Bayesian inference Germain et al.
-(2016)
-$\pi_{{\mathcal{L}},N}({\bm{\Theta}})=\pi_{0}({\bm{\Theta}})\exp(\lambda-N\hat{L}_{N}({\bm{\Theta}})).$ (9)
-As $N/P\to\infty$ the posterior is strongly peaked around a loss minimizer ${\bm{\Theta}}^{*}_{\mathcal{L}}\in\arg\min_{{\bm{\Theta}}\in{\mathcal{P}}}L({\bm{\Theta}})$.
-Defining ${\bm{\Sigma}}^{-1}_{0}$ as the curvature of $-\ln\pi_{0}({\bm{\Theta}})$ around ${\bm{\Theta}}^{*}_{\mathcal{L}}$, we use Laplace’s method Swinburne and Perez (2024) as $N/P\to\infty$ to write
-[…]
-[…]
-$\lim_{N/P\to\infty}{\mathcal{G}}[\pi_{\mathcal{L}}]=\frac{1}{2}{\rm Tr}({\bm{\Sigma}^{*}_{\mathcal{L}}}{\bm{\Sigma}}_{\mathcal{Y}}^{-1})=\mathcal{O}(1/\epsilon^{2}).$ (11)
-${\bm{\Sigma}}^{*}_{\mathcal{L}}=\langle[{\mathcal{E}}({\mathbf{X}})-{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}}^{*}_{\mathcal{L}})][{\mathcal{E}}({\mathbf{X}})-{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}}^{*}_{\mathcal{L}})]^{\top}\rangle$ is the error covariance around ${\bm{\Theta}}^{*}_{\mathcal{L}}$ and we use ${\rm Tr}({\bm{\Sigma}}_{\mathcal{Y}})=\mathcal{O}(\epsilon^{2})$.
-Under misspecification it is thus clear that minimizing this upper bound is sub-optimal Masegosa (2020) as the generalization error diverges as $1/\epsilon^{2}$.
-In the next section we analyze the generalization error as $\epsilon\to 0$ to derive a condition for any $\pi({\bm{\Theta}})$ that ensures a finite generalization error.
 II Analysis of the generalization error in the near-deterministic limit
 We consider the generalization error (6) in the near-deterministic limit, ${\mathcal{G}}_{0}[\pi]\equiv\lim_{\epsilon\to 0}{\mathcal{G}}[\pi]$.
-The integral over ${\mathbf{Y}}$ in (6) concentrates at ${\mathbf{Y}}={\mathcal{E}}({\mathbf{X}})$; using $\langle\dots\rangle_{\mathcal{X}}\equiv\int\dots\rho_{\mathcal{D}}({\mathbf{X}}){\rm d}{\mathbf{X}}$ yields, using Laplace’s method Swinburne and Perez (2024)
-${\mathcal{G}}_{0}[\pi]$ $=-\left\langle\ln\left|\rho_{\mathcal{M}}({\mathcal{E}}({\mathbf{X}})|{\mathbf{X}})\right|\right\rangle_{\mathcal{X}}$ (12)
-$=-\left\langle\ln\left|\int_{\mathcal{P}}\rho_{\mathcal{M}}({\mathcal{E}}({\mathbf{X}})|{\mathbf{X}},{\bm{\Theta}})\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}\right|\right\rangle_{\mathcal{X}}.$
+The integral over ${\mathbf{Y}}$ in (6) concentrates at ${\mathbf{Y}}={\mathcal{E}}({\mathbf{X}})$; using $\langle\dots\rangle_{\mathcal{X}}\equiv\int\dots\rho_{\mathcal{D}}({\mathbf{X}}){\rm d}{\mathbf{X}}$ yields, using Laplace’s method (appendix C)
+${\mathcal{G}}_{0}[\pi]=-\left\langle\ln\left|\rho_{\mathcal{M}}({\mathcal{E}}({\mathbf{X}})|{\mathbf{X}})\right|\right\rangle_{\mathcal{X}}=-\left\langle\ln\left|\int_{\mathcal{P}}\rho_{\mathcal{M}}({\mathcal{E}}({\mathbf{X}})|{\mathbf{X}},{\bm{\Theta}})\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}\right|\right\rangle_{\mathcal{X}}.$ (12)
 Whilst ${\mathbf{Y}}={\mathcal{E}}({\mathbf{X}})$ is a unique maximum, to ensure $\min_{{\mathbf{X}}\in{\mathcal{X}}}\rho_{\mathcal{M}}({\mathcal{E}}({\mathbf{X}})|{\mathbf{X}})>0$ and thus avoid divergence in ${\mathcal{G}}_{0}[\pi]$ as $\epsilon\to 0$ we require $\pi({\bm{\Theta}})$ has mass in every pointwise optimal parameter set (POPS)
 ${\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})\equiv\left\{{\bm{\Theta}}\,|{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}})={\mathcal{E}}({\mathbf{X}})\right\}.$ (13)
-Any model with a constant term will ensure ${\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})\neq\emptyset$, but under misspecification it clear that $\cap_{{\mathbf{X}}\in{\mathcal{X}}}{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})=\emptyset$.
+Any model with a constant term will ensure ${\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})\neq\emptyset$, i.e.
+that the POPS is not empty, but under misspecification it clear that $\cap_{{\mathbf{X}}\in{\mathcal{X}}}{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})=\emptyset$, i.e.
+that no one parameter choice can be a member of all POPS.
 To express the final generalization error, we introduce a pointwise mass function
-${\rm m}_{\mathcal{E}}({\bm{\Theta}},{\mathbf{X}})$ $\equiv\int_{\mathcal{P}}\delta({\bm{\Theta}}-{\bm{\Theta}}^{\prime})\rho_{\mathcal{M}}({\mathcal{E}}({\mathbf{X}})|{\mathbf{X}},{\bm{\Theta}}^{\prime}){\rm d}{\bm{\Theta}}^{\prime},$
+${\rm m}_{\mathcal{E}}({\bm{\Theta}},{\mathbf{X}})$ $\equiv\int_{\mathcal{P}}\delta({\bm{\Theta}}-{\bm{\Theta}}^{\prime})\rho_{\mathcal{M}}({\mathcal{E}}({\mathbf{X}})|{\mathbf{X}},{\bm{\Theta}}^{\prime}){\rm d}{\bm{\Theta}}^{\prime},$ (14)
 which gives our first main result, a generalization error for deterministic surrogate models of
@@
 We now investigate two candidate distributions which satisfy (16), an $\mathcal{O}(N)$ ensemble ansatz and $\mathcal{O}(P)$ hypercube ansatz, finding approximate variational minima that allow for efficient deployment on high-dimensional datasets.
-II.1 POPS-constrained ensemble and hypercube ansatz
+III POPS-constrained ensemble and hypercube ansatz
 The covering constraint (16) requires $\pi({\bm{\Theta}})$ to have mass in the POPS of each training point.
@@
 ${\bm{\Theta}}^{*}_{\mathbf{X}}\in\mathcal{T}^{*}_{E},\quad{\bm{\Theta}}^{*}_{\mathbf{X}}\equiv\arg\min_{{\bm{\Theta}}_{\mathbf{X}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})}L({\bm{\Theta}}_{\mathbf{X}}).$ (17)
-These pointwise fits will be closely clustered around the minimum loss solution ${\bm{\Theta}}^{*}_{\mathcal{L}}$, as measured by the Fisher metric ${\boldsymbol{\mathcal{I}}}^{*}_{\mathcal{L}}$.
-In the supplementary materials Swinburne and Perez (2024) we explore a weighted ensemble ansatz
+These pointwise fits will be closely clustered around the minimum loss solution ${\bm{\Theta}}^{*}_{\mathcal{L}}$, as measured by the Fisher metric ${\bm{\mathcal{I}}}^{*}_{\mathcal{L}}$.
+In appendix D we explore a weighted ensemble ansatz
 $\pi^{*}_{E}({\bm{\Theta}})=\langle{\rm w}({\mathbf{X}})\delta({\bm{\Theta}}-{\bm{\Theta}}^{*}_{\mathbf{X}})\rangle,$ (18)
@@
 The ensemble ansatz produces highly informative max/min bounds for any test point, providing an envelope for worst case errors, as we show in figure 1.
-However, even when employing an optimal reweighting ${\rm w}^{*}({\mathbf{X}})$ (see Swinburne and Perez (2024)), the ensemble ansatz systematically underestimates moments of the test error distribution, as shown in figure 2.
+However, even when employing the variational optimal reweighting ${\rm w}^{*}({\mathbf{X}})$ (see appendix D.1), the ensemble ansatz systematically underestimates moments of the test error distribution, as shown in figure 2.
 In addition, both resampling and storage are $\mathcal{O}(N)$, which can become problematic for the $N/P\to\infty$ limit of interest.
+As a result, whilst further investigation of ensemble schemes and alternatives to the variational optimal reweighting ${\rm w}^{*}({\mathbf{X}})$ (appendix D.1), we now look for an alternative ansatz beyond the ensemble approach.
 For practical applications, the parameter distribution should ideally require at most $\mathcal{O}(P^{2})$ effort for storage and $\mathcal{O}(P)$ for resampling.
@@
 We find this efficient hypercube posterior to give excellent bounding and prediction of test errors for $P\in[3,1600]$, as shown in figures 2 and 3.
-An additional advantage is the simplicity of implementation for linear models; an example Python code is provided in the SM Swinburne and Perez (2024).
-II.2 POPS-constrained loss minimization for linear models
+III.1 POPS-constrained loss minimization for linear models
 Both the ensemble and hypercube ansatz require POPS-constrained loss minimizing fits ${\bm{\Theta}}^{*}_{\mathbf{X}}\in\mathcal{P}_{\mathcal{M}}({\mathbf{X}})$.
@@
 Minimizing the PAC-Bayes bound (8) with a Gaussian prior $\pi_{0}({\bm{\Theta}})=\mathcal{N}({\bm{\Theta}}|{\bf 0},{\bm{\Sigma}}_{0})$ gives a global loss minimizer ${\bm{\Theta}}^{*}_{\mathcal{L}}$ of
-${\bm{\Theta}}^{*}_{\mathcal{L}}$ $={\bf A}\langle{\mathbf{F}}({\mathbf{X}}){\mathcal{E}}({\mathbf{X}})\rangle_{\mathcal{X}},$ (19)
-${\bf A}$ $=[{\bm{\Sigma}}_{\mathcal{Y}}{\bm{\Sigma}}^{-1}_{0}/N+\langle{\mathbf{F}}({\mathbf{X}})[{\mathbf{F}}({\mathbf{X}})]^{\top}\rangle]^{-1}_{\mathcal{X}}.$
-The epistemic covariance for small $N/P$ can be incorporated in a variational PAC-Bayes scheme Swinburne and Perez (2024) that will be developed further in future work.
+${\bf A}=[{\bm{\Sigma}}_{\mathcal{Y}}{\bm{\Sigma}}^{-1}_{0}/N+\langle{\mathbf{F}}({\mathbf{X}})[{\mathbf{F}}({\mathbf{X}})]^{\top}\rangle]^{-1}_{\mathcal{X}},\quad{\bm{\Theta}}^{*}_{\mathcal{L}}={\bf A}\langle{\mathbf{F}}({\mathbf{X}}){\mathcal{E}}({\mathbf{X}})\rangle_{\mathcal{X}},$ (20)
+The epistemic covariance for small $N/P$ can be incorporated in a variational PAC-Bayes scheme (appendix D.1) that will be developed further in future work.
 However, in the following, we concentrate on misspecification uncertainties in the limit $N/P\to\infty$, where the influence of $\bm{\Sigma}_{0}$ is negligible.
@@
 ${\bm{\Theta}}^{*}_{{\mathbf{X}}}={\bm{\Theta}}^{*}_{\mathcal{L}}+\frac{{\mathcal{E}}({\mathbf{X}})-{\bm{\Theta}}^{*}_{\mathcal{L}}\cdot{\mathbf{F}}({\mathbf{X}})}{{\rm h}({\mathbf{X}})}{\bf A}{\mathbf{F}}({\mathbf{X}}),$ (21)
-where ${\rm h}({\mathbf{X}})=[{\mathbf{F}}({\mathbf{X}})]^{\top}{\bf A}{\mathbf{F}}({\mathbf{X}})$ is the leverage, an outlier measure closely related to the distance of (Mahalanobis, 1936).
+where ${\rm h}({\mathbf{X}})=[{\mathbf{F}}({\mathbf{X}})]^{\top}{\bf A}{\mathbf{F}}({\mathbf{X}})$ is the leverage, an outlier measure closely related to the distance of Ref.
+Mahalanobis, 1936.
 The minimum loss solution is the leverage-weighted centroid of pointwise optima, ${\bm{\Theta}}^{*}_{\mathcal{L}}=\langle{\rm h}({\mathbf{X}}){\bm{\Theta}}^{*}_{{\mathbf{X}}}\rangle_{\mathcal{X}}$.
 Equation (21) is our third main result, a scheme to incorporate misspecification uncertainty into least-squares linear regression for minimal overhead, requiring only $\mathcal{O}(N)$ inner products over feature vectors, which can then be used to construct the ensemble or hypercube ansatz given above.
-A simple python implementation is provided in the SM Swinburne and Perez (2024).
-III Numerical Experiments
-III.1 Comparison to Bayesian ridge regression and minimization of $\mathcal{G}_{E}$
+Details of the implementation are provided in appendix E.
+In appendix F we compare the performance of our POPSRegression implementation against BayesianRidge from sklearn, finding an minimal overhead of around 2×2$\times$2 × for fitting and essentially identical performance for prediction, even when returning max/min bounds in addition to standard deviations.
+IV Numerical Experiments
+In this section we apply our POPS approach to a variety of synthetic and real datasets, comparing our POPS-ensemble and POPS-hypercube ansatz to Bayesian inference and the minimization of the PAC-Bayes generalization error $\mathcal{G}_{E}$.
+We also discuss our approach in the context of existing methods for uncertainty quantification in atomistic machine learning and discuss how these parameter uncertainties can be propagated up scales to predict material properties with uncertaintyMaliyov et al.
+(2024).
+IV.1 Comparison to Bayesian ridge regression and minimization of $\mathcal{G}_{E}$
 We applied Bayesian ridge regression and our POPS-ansatz to regress a quadratic polynomial against a sinusoidal function, the constant, linear and quadratic features giving $P=3$ model parameters.
-The SM Swinburne and Perez (2024) shows the same procedure for a model with $P=6$ independent features.
+Appendix G shows the same procedure for a model with $P=6$ independent features.
 For these simple models it is possible to approximately minimize the generalization error $\mathcal{G}_{E}$ of some discrete model ensemble, providing we regularize with a finite aleatoric uncertainty ${\bm{\Sigma}}_{\mathcal{Y}}=\sigma{\bm{\Sigma}}^{*}_{\mathcal{L}}$.
-With increasing $\sigma$ we have $\mathcal{G}\to\mathcal{L}$ and thus expect the ensemble to concentrate on ${\bm{\Theta}}^{*}_{\mathcal{L}}$, whilst in the small $\sigma$ regime of interest minimizing $\mathcal{G}_{E}$ becomes increasingly unstable.
-We found the same small $\sigma$ instability when applying an ensemble approach from Masegosa (2020) which targets a second-order PAC-Bayes loss bound, designed for misspecified probabilistic models where $\sigma$ can be large.
+With increasing $\sigma$ we have $\mathcal{G}\to\mathcal{L}$ and thus expect the ensemble to concentrate on ${\bm{\Theta}}^{*}_{\mathcal{L}}$.
+In the small $\sigma$ regime of interest minimizing $\mathcal{G}_{E}$ becomes increasingly unstable, meaning that the final minimization is highly sensitive to initial conditions and frequently diverges, leading to conditioning issues.
+We found the same small $\sigma$ instability when applying an ensemble approach from Ref.
+Masegosa, 2020 which targets a second-order PAC-Bayes loss bound, designed for misspecified probabilistic models where $\sigma$ can be large.
 These instabilities show existing misspecification-aware methods require careful hyperparameter tuning and low model dimension to avoid vanishing gradients or other numerical issues.
@@
 The max/min range of the numerical ensemble shows similar behavior for $\sigma=1,1/3$ whilst at $\sigma=1/6$ test errors are bounded but significantly overestimated, which worsens as $\sigma$ decreases (whilst stable).
-This was also seen on higher dimensional datasets Swinburne and Perez (2024).
 Importantly, in all cases our POPS ansatz $\pi^{*}_{E}$ and $\pi^{*}_{\mathcal{H}}$ give excellent bounding of test errors and are stable in the near-deterministic limit $\sigma\to 0$.
 As expected, the hypercube ansatz $\pi^{*}_{\mathcal{H}}$ gave slightly more conservative bounds.
+In appendix G we provide a similar test of directly minimizing the generalization error for higher dimensional datasets, with $P=6$ independent features.
+As for the polynomial case, we find that strong regularization is required for minimization to be stable, and in this limit the solution closely follows that of loss minimization.
+At weaker regularization, the direct minimization becomes systematically closer to the prediction of the POPS ansatz $\pi^{*}_{\mathcal{H}}$.
+These albeit simple results give strong evidence that our ansatz is indeed a stable and highly efficient means to find approximate minimizers of ${\mathcal{G}}$ that would otherwise be intractable.
 Figure 1: Regression of a deterministic quadratic polynomial ($P=3$) model onto a sinusoidal ”simulation engine”, trained on $N=100$ points.
@@
 Bottom: numerically optimized $\mathcal{G}_{E}$ for a uniformly weighted $N$-ensemble, regularized with $\Sigma_{\mathcal{Y}}=\sigma\Sigma^{*}_{\mathcal{L}}$, for $\sigma=1,1/3,1/6$.
-Lower $\sigma$ values gave numerical instabilities.
-III.2 Test errors in medium-dimensional linear regression
-We now consider higher/medium dimensional problems, $P=10-50$, where numerical minimization of ${\mathcal{G}}_{E}$ is not possible.
-The simulation engines were constrained to be misspecified from the linear surrogate models by combining features in quadratic, cubic or sinusoidal functions or randomly sampling the linear feature coefficients from some predetermined subset of models.
-A noise term that simulates the presence of variability orthogonal to the chosen feature set was also used.
+Lower $\sigma$ values gave numerical instabilities and are not presented (see appendix D).
+IV.2 Test errors in medium-dimensional linear regression
+We now consider higher/medium dimensional problems, $P=10-100$, where numerical minimization of ${\mathcal{G}}_{E}$ is not possible.
+The simulation engines were constrained to be misspecified from the linear surrogate models by combining features in a cubic polynomial.
+We also investigated the influence of adding a simple Gaussian noise term to simulates the presence of variability orthogonal to the chosen feature set.
 In all cases, independent test and training data were generated, typically with a 10:90 test:train split.
-However, all conclusions remain robust to varying the precise nature of data generation and the classes of engine considered Swinburne and Perez (2024).
-As shown, the $\mathcal{O}(P)$ hypercube ansatz $\pi^{*}_{\mathcal{H}}$ provides an excellent prediction and bounding, of the test error distribution whilst the $\mathcal{O}(N)$ ensemble $\pi^{*}_{E}$ underestimates moments of the error distribution and the max/min envelope bound.
-Importantly, the envelope violation rate of the hypercube ansatz $\pi^{*}_{\mathcal{H}}$ drops with $N/P$, as shown in figure 2 for a range of $P$ and $N/P$.
-The simultaneous prediction and bounding of test errors is highly valuable for surrogate model deployment, allowing not only the prediction of expected errors but a robust assessment of worst case error scenarios.
-We conclude with an application of the hypercube ansatz to challenging high-dimensional $P=\mathcal{O}(1000)$ datasets from atomic machine learning.
+However, all conclusions remain robust to varying the precise nature of data generation.
+In appendix G we provide examples showing this performance remains when varying the form of simulation engine to quadratic and sinusoidal functions or randomly sampling the linear feature coefficients from some predetermined subset of models.
 Figure 2: Test errors of a misspecified linear surrogate model on a cubic simulation engine.
@@
 Right: Probability of envelope violation for the $\pi^{*}_{\mathcal{H}}$ ansatz with $P$ and $N/P$.
-III.3 High-dimensional models for atomic simulations
-Figure 3: Application to interatomic potentials.
-Left: test error histogram at $P=1596$, $N/P=81$ for the minimum loss model (black) and predictions from the POPS-hypercube $\pi^{*}_{\mathcal{H}}$ (green) ansatz.
+As shown, the $\mathcal{O}(P)$ hypercube ansatz $\pi^{*}_{\mathcal{H}}$ provides an excellent prediction and bounding, of the test error distribution whilst the $\mathcal{O}(N)$ ensemble $\pi^{*}_{E}$ underestimates moments of the error distribution and the max/min envelope bound.
+Importantly, the envelope violation rate of the hypercube ansatz $\pi^{*}_{\mathcal{H}}$ drops with $N/P$, as shown in figure 2 for a range of $P$ and $N/P$.
+The simultaneous prediction and bounding of test errors is highly valuable for surrogate model deployment, allowing not only the prediction of expected errors but a robust assessment of worst case error scenarios.
+IV.3 Application to machine learning interatomic potentials
+We now demonstrate the performance of the method on challenging high-dimensional $P=\mathcal{O}(100-1000)$ regression tasks from atomic machine learning.
+Atomic-scale simulations of materials were traditionally broadly classified into two types:
+  1. a)
+    Empirical simulations with physically-motivated parametric models of interatomic interactions. These allow for fast simulations at a qualitative level of accuracy.
+  2. b)
+    First-principle quantum simulations obtained from approximate solutions of the full Schrodinger equation. These incur a high computational cost and exhibit poor scalability with system size.
+The use of machine learning techniques in atomic simulation has blurred the lines between these two limits by promising near-quantum accuracy at a much more favorable computational cost and scaling (Deringer et al., 2019).
+The total energy of the system $V$ is typically factored into a sum of per-atom contributions $V_{i}$, which are themselves parameterized in term of a feature vector that describes the local atomic environment ${\bf F}_{i}(\{{\bf x}_{j}|r_{ij}<r_{c}\})$, where $r_{c}$ is a interaction range.
+Training a machine learning model proceeds by minimizing a squared error loss between reference energies and gradients.
+While certain approaches are systematically improvable, the computational cost associated with repeated evaluations can limit the complexity of the models that are selected for applications.
+Further, quantum calculations can be tightly converged, so that their intrinsic error (w.r.t.
+a fully converged calculation) can be made small compared to typical errors incurred by computationally-efficient models.
+This regime corresponds to the underparametrized deterministic setting considered here.
+Estimating the errors incurred by such machine learning models is highly sought after to assess the robustness of conclusions drawn from the simulations (Wen and Tadmor, 2020; Li et al., 2018; Gabriel et al., 2021; Wan et al., 2021).
+In the following, we explore the application of our POPS approach for a linear interatomic potential model $V_{i}={\bm{\Theta}}\cdot{\bf F}_{i}$, with forces obtained by taking spatial gradients.
+Whilst we choose a particlar form of linear interatomic potential model, the presented approach is general to the wide range of available implementationsLysogorskiy et al.
+(2021); Wood and Thompson (2018); Podryabinkin and Shapeev (2017); Goryaeva et al.
+(2021).
+Figure 3: Fitting a qSNAPWood and Thompson (2018) interatomic potential to a diverse tungsten datasetKarabin and Perez (2020).
+Left: representative training configurationsMontes de Oca Zapiain et al.
+(2022).
+Center: test error histogram at $P=1596$, $N/P=81$ for the minimum loss model (black) and predictions from the POPS-hypercube $\pi^{*}_{\mathcal{H}}$ (green) ansatz.
 MAE: mean absolute error relative to the minimum loss solution.
@@
 Right: Probability of envelope violation for the $\pi^{*}_{\mathcal{H}}$ ansatz with $N/P$.
-Atomic-scale simulations of materials were traditionally broadly classified into two types:
-  • Empirical simulations with physically-motivated parametric models of interatomic interactions. These allow for fast simulations at a qualitative level of accuracy.
-  • First-principle quantum simulations obtained from approximate solutions of the full Schrodinger equation. These incur a high computational cost and exhibit poor scalability with system size.
-The advent of ML has blurred the lines between these two limits by promising near-quantum accuracy at a much more favorable computational cost and scaling (Deringer et al., 2019).
-The total energy of the system $V$ is typically factored into a sum of per-atom contributions $V_{i}$, which are themselves parameterized in term of a feature vector that describes the local atomic environment ${\bf F}_{i}(\{{\bf x}_{j}|r_{ij}<r_{c}\})$, where $r_{c}$ is a interaction range.
-Training an ML model then proceeds by minimizing a squared error loss between reference energies and gradients.
-While certain ML approaches are systematically improvable, the computational cost associated with repeated evaluations can limit the complexity of the models that are selected for applications.
-Further, quantum calculations can be tightly converged, so that their intrinsic error $\epsilon$ (w.r.t.
-a fully converged calculation) can be made small compared to typical errors incurred by computationally-efficient models.
-This regime corresponds to the underparametrized deterministic setting considered here.
-Estimating the errors incurred by such ML models is highly sought after to assess the robustness of conclusions drawn from the simulations (Wen and Tadmor, 2020; Li et al., 2018; Gabriel et al., 2021; Wan et al., 2021) .
-In the following, a linear surrogate model $V_{i}={\bm{\Theta}}\cdot{\bf F}_{i}$ was used to predict per-atom contributions.
+Our first application is to pure tungsten, a key material for nuclear fusion applicationsGoryaeva et al.
+(2021).
+Atomic training data was generated using a recently introduced information-theoretic approach (Karabin and Perez, 2020; Montes de Oca Zapiain et al., 2022) that aims to maximize an estimated feature entropy for a given set of feature functions, giving an an extremely (ideally, maximally) diverse training set.
+10,000 atomistic configurations containing between 2 and 32 tungsten atoms were generated and characterized using Density Functional Theory (Kohn and Sham, 1965), yielding 167,922 training data points in total (10,000 energies and 157,922 gradient components).
+Per-atom energies are expressed as a linear combination of $P=1596$ so-called bispectrum components within the ‘Quadratic’ Spectral Neighborhood Analysis Potential (SNAP) formalism (Wood and Thompson, 2018), known as qSNAP, which can (in spite of what the name suggests) be expressed in linear model form $V_{i}={\bm{\Theta}}\cdot{\bf F}_{i}$ by defining the ${\bf F}_{i}$ as products of standard SNAP features.
 Numerical experiments were carried out by randomly sub-selecting $N$ training points (either energies or gradient components), and by generating a discrete ensemble of $N$ pointwise-optimal models according to Eq.
-20, up to a maximal value of $N=129,853$.
+21, up to a maximal value of $N=129,853$.
 The properties of the ensemble are then evaluated on a complementary set of 10,000 hold-out testing points.
 Error analysis is carried out in the context of the POPS-hypercube ansatz.
-Even on the full dataset, the analysis requires on the order of 20 minutes on 1 CPU core.
+Even on the full dataset, the analysis requires only on the order of 20 minutes on 1 CPU core.
 It is therefore extremely lightweight in addition to being easily parallelizable, enabling its use for extremely large datasets.
@@
 The MAE estimated from the resampled models exceeds the measured value by only 23%, while providing a very good description of the overall error distribution.
-The information generated by this simple and efficient approach will be invaluable to estimate the uncertainties on a range of different physical quantities inferred from atomistic simulations.
-IV Conclusion
+Our second test case fits a simple SNAP potential to ab initio data for the NbMoTaW high-entropy alloy from Li et al.Li et al.
+(2020).
+In this case, we use a simple form of multi-component descriptor that is linear in the number of species, and additionally choose a small set of descriptor functions, giving a total of $P=120$ adjustable parameters to fit $N=43,250$ force observations of this chemically complex system, with a further $N=121248$ force observations held out for testing.
+As can be seen in figure 4, the POPS-hypercube shows excellent performance in bounding prediction errors and in estimating the test error distribution, albeit with a slight underestimation in the tails of the error distribution.
+To investigate the ability of our POPS approach to capture these large-error predictions, we also plot a common metric for uncertainty quantification schemes, the correlation between predicted and observed errors.
+Whilst such a correspondence is not expected to be exact, as the observed error is a stochastic quantity, there is clear benfit in the ability to asses, with confidence, whether a given prediction is expected to have small or large error.
+In the present setting, an appropriate comparison is the absolute error of the loss minimizing model against the predicted by the hypercube posterior prediction.
+A well perfoming model should a) bound the absolute error from above, i.e.
+not predict low error when the actual error is high and b) have a max-min range which increases for points with large error.
+As shown in the right of figure 4, we see these properties are indeed observed in the HEA example.
+An in-depth study of our POPS apporach in active learning schemes will be the subject of future work.
+Figure 4: Fitting a SNAPWood and Thompson (2018) interatomic potential to a NbMoTaW high-entropy alloy (HEA) training setLi et al.
+(2020).
+Far Left: representative HEA configurations.
+Left: test error histogram at $P=120$, $N/P=376$ for the minimum loss model (black) and predictions from the POPS-hypercube $\pi^{*}_{\mathcal{H}}$ (green) ansatz.
+MAE: mean absolute error relative to the minimum loss solution.
+EV: envelope violation, points lying outside of the max/min bound.
+Right: probability of envelope violation for the $\pi^{*}_{\mathcal{H}}$ ansatz with $N/P$.
+Far right: correlation of actual vs predicted error in the test set.
+See main text.
+IV.4 Application to molecular cheminformatics models
+As a final application, we apply the POPS scheme to a recently introduced linear graphlet modelTynes et al.
+(2024) for predicting molecular properties.
+Whilst superficially similar to the interatomic potential application, the graphlet features are quite distinct as they are designed to produce explanatory models that detail the contribution of different local bond environment to the total atomic energy rather than providing a smoothly differentiable interatomic potential.
+The graphlet features, which are essentially discrete counts of the different possible subgraphs of the molecular graph, have a much higher sparsity than the atomic descriptors employed in the previous examples.
+Here, we consider the formation energies of around 50,000 small organic molecules from the well-known QM9 dataset of small organic moleculesRuddigkeit et al.
+(2012); Ramakrishnan et al.
+(2014).
+Restricting feature extraction to graphlets of maximum order 3 or 4 gave a total of $P=247$ or $P=1239$ features, respectively.
+Once again, the formation energy model is linear in the graphlet features.
+As shown in figure 5, the POPS-hypercube again provides an excellent prediction and bounding of test errors, with the probability of envelope violation decreasing with increasing $N/P$.
+A detailed study of the use of our POPS approach for linear models in a wide range of applications will be the subject of a separate publication.
+Figure 5: Fitting a linear graphlet modelTynes et al.
+(2024) to predict energies from the QM9 datasetRuddigkeit et al.
+(2012).
+Left: representative small organic molecules.
+Center: test error histogram at $P=247$, $N/P=42$ for the minimum loss model (black) and predictions from the POPS-hypercube $\pi^{*}_{\mathcal{H}}$ (green) ansatz.
+MAE: mean absolute error relative to the minimum loss solution.
+EV: envelope violation, points lying outside of the max/min bound.
+Right: Probability of envelope violation for the $\pi^{*}_{\mathcal{H}}$ ansatz with $N/P$.
+V Comparison against existing approaches for uncertainty quantification and propagation in atomic simulation
+As discussed in the previous section, uncertainty quantification schemes for machine learning interatomic potentials are actively sought after to assess the robustness of conclusions drawn from atomic simulationsWen and Tadmor (2020); Li et al.
+(2018); Gabriel et al.
+(2021); Wan et al.
+(2021); Best et al.
+(2024); Kellner and Ceriotti (2024); Thaler et al.
+(2023).
+Existing approaches for uncertainty quantification can be broadly categorized into two types, those that aim to directly estimate the error distribution of the surrogate model, and those (such as the POPS approach) that aim to capture model errors through the determination of parameter uncertainties.
+Whilst direct error estimation is central to active learning workflowsLi et al.
+(2015); Podryabinkin and Shapeev (2017); Vandermause et al.
+(2020), these estimates can not typically be used to propagate uncertainty through the entire simulation engine to predict uncertainty on quantities of interest.
+This is because the result of any atomic simulation workflow is some quantity of interest (e.g.
+formation energies, elastic constants, or relaxed structures).
+Each quantity of interest is an (implicit) function of the parameters of the chosen interatomic potential, and will in general vary with the parameters in a non-trivial manner due to the strong correlations in the simulation routine.
+For example, computing defect formation energies require comparing two very similar, and thus highly correlated simulations.
+A direct estimate of energy error for each configuration will in general significantly overestimate the error in the formation energy, as the correlation between the two structures is not accounted for.
+As a result, schemes which estimate uncertainty in model parameters are desirable for uncertainty propagation.
+The propagation can be achieved if we can predict how simulations are affected by variation in parameter, either through i) implicit differentiation methodsMaliyov et al.
+(2024) ii) reweighting of existing simulations using e.g.
+thermodynamic perturbation theoryImbalzano et al.
+(2021), or iii) direct resampling simulations by drawing samples from the parameter posterior distribution Best et al.
+(2024).
+A detailed study combining the first two of these approaches (implicit differentiation and thermodynamic perturbation) with the POPS procedure will be the subject of future work.
+The POPS regression procedure presented above is a simple and efficient approach which returns uncertainties on the parameters of misspecified models.
+To the best of our knowledge, existing schemes for uncertainty quantification use as their objective function some form of loss, and thus cannot account for misspecification in parameter uncertainties of the model.
+In this sense, the POPS scheme is unique in its ability to robustly capture the parameter uncertainty of misspecified linear models.
+The most directly-comparable scheme to POPS is ensembling, which typically train an ensemble of models on random subsets of the training data, a process known as baggingBreiman (1996).
+Various forms of ensemble methods are a popular approach to quantify uncertainty in neural network modelsAbe et al.
+(2022) and have been successfully applied to uncertainty quantification for neural network interatomic potentialsImbalzano et al.
+(2021); Kellner and Ceriotti (2024); Vita et al.
+(2024).
+It is known that bagging and ensemble methods more generally are most effective for high-capacity models in presence of appreciable aleatoric uncertaintyBreiman (1996).
+In the low-noise regime, i.e.
+with weak aleatoric uncertainty, ensemble uncertainty estimates can be expected to significantly underestimate the true uncertainty, even when using high-capacity models such as neural networks.
+For low-capacity models, such as the underparametrized linear models considered here, bagging methods are not expected to produce useful uncertainty estimates.
+In most settings, loss-based uncertainty estimates require significant calibrationImbalzano et al.
+(2021); Kellner and Ceriotti (2024); Vita et al.
+(2024), or the use of closely-related methods such as conformal predictionBest et al.
+(2024), which similarly require some form of calibration.
+In addition, all loss-based schemes should give vanishing parameter uncertainty in the large data limit, as discussed in section I.
+Whilst bagging schemes for neural networks do provide finite parameter uncertainty, this is predominantly due to the multi-modal loss landscape, and has unclear relationship to the true uncertainty and generalization error of the model.
+Indeed, even the POPS-ensemble ansatz was found to underestimate moments of the error distribution, motivating the hypercube ansatz.
+An interesting direction for future work is to investigate how the POPS coverage criterion, equation (16), can be used to systematically construct improved ensemble methods in the low-noise regime.
+We end this section with a comment on the wider validity of the POPS method for atomic simulation.
+Clearly, the requirement to be in the underparametrised limit $N\gg P$, can in practice lead to a large amount of training data when $P$ is large.
+In addition, this condition implicitly assumes that the model is not applied in a strongly extrapolative regime, which is formally true for any ensemble methodKellner and Ceriotti (2024).
+Whilst we have not found this to be an issue in practical applications, this theoretical requirement is also common to ensemble methodsKellner and Ceriotti (2024).
+In practical tests we do not find this to be problematic as the POPS scheme will typically give an overestimate of model uncertainty in regions of low training data density, which is a robust means to quantify extrapolation in high dimensionZeni et al.
+(2022).
+Nevertheless, we recommend using one of the many extrapolation grades / outlier measures for high-dimensional linear modelsPodryabinkin and Shapeev (2017); Lysogorskiy et al.
+(2021); Goryaeva et al.
+(2020); Zeni et al.
+(2022) to ensure that the POPS conditions are satisfied.
+VI Conclusion
 In spite of its ubiquity in practical applications, misspecification is often ignored in Bayesian regression approaches based on expected loss minimization, which can lead to erroneous conclusions regarding parameter uncertainty quantification, surrogate model selection, and error propagation.
@@
 Our POPS-constrained ansatz are shown to produce excellent parameter distributions which give accurate prediction of test error distributions and highly informative worst-case bounds on model predictions, in the important misspecified, near-deterministic regime where existing methods fail.
-Synthetic test data
-All synthetic data was generated with Python/JAX code, to be made available following peer review.
-Atomic simulation data
-Atomic training data was generated using an information-theoretic approach (Karabin and Perez, 2020; Montes de Oca Zapiain et al., 2022).
-10,000 atomistic configurations containing between 2 and 32 tungsten atoms were generated and characterized using Density Functional Theory (Kohn and Sham, 1965), yielding 167,922 training data points in total (10,000 energies and 157,922 gradient components).
-Per-atom energies are expressed as a linear combination of $P=1596$ so-called bispectrum components within the ‘Quadratic’ Spectral Neighborhood Analysis Potential (SNAP) formalism (Wood and Thompson, 2018), which can be expressed as a linear surrogate model $V_{i}={\bm{\Theta}}\cdot{\bf F}_{i}(\{{\bf x}_{j}|r_{ij}<r_{c}\})$.
+Importantly, POPS encode model error as misspecification uncertainty on model parameters, which is essential for the propagation of uncertainty to simulation resultsMaliyov et al.
+(2024); Imbalzano et al.
+(2021).
+The application of POPS parameter uncertainties to multi-scale uncertainty propagation will be the subject of forthcoming work.
+In addition, a clear extension of the POPS approach is to non-linear models, for interatomic potentialsLysogorskiy et al.
+(2021); Vita et al.
+(2024) or more widely in scientific machine learning.
+As the loss function is non-convex, the central challenge will be to efficiently generate an optimal set of POPS-constrained loss minimizers, which can be achieved analytically in the linear case.
+Data availability
+A Python implementation of the POPS-constrained linear regression algorithm, following the Scikit-learn linear_model APIPedregosa et al.
+(2011), is available on GitHub at https://github.com/tomswinburne/POPS-Regression.git, where we also provide a Jupyter notebook to reproduce a selection of results from the paper.
+The library can also be installed via the pip package manager via pip install POPSRegression.
 Acknowledgements
 This work was initiated during a visit at the Institute for Pure and Applied Mathematics at the University of California, Los Angeles (supported by National Science Foundation (NSF) grant DMS-1925919) and continued at the Institute for Mathematical and Statistical Innovation, University of Chicago (supported by the NSF grant DMS-1929348).
-Their hospitality is graciously acknowledged.
-TDS gratefully acknowledges support from ANR grants ANR-19-CE46-0006-1 and ANR-23-CE46-0006-1, IDRIS allocation A0120913455 and an Emergence@INP grant from the CNRS.
+Their hospitality is gratefully acknowledged.
+TDS was supported by ANR grants ANR-19-CE46-0006-1 and ANR-23-CE46-0006-1, IDRIS allocation A0120913455 and an Emergence@INP grant from the CNRS.
 D.P.
@@
 89233218CNA000001).
-Appendix A PAC-Bayes concentration inequalities for regression
-PAC-Bayes inequalities were originally designed for classification problems where the loss function is bounded to $[0,1]$.
-The extension to regression requires treatment of unbounded losses, which is discussed in detail in (Germain et al., 2016).
-Determining the value of the upper bound for $C_{0}[\pi_{0}]$ in the PAC-Bayes concentration inequality (21) requires certain assumptions on the distribution of the error between the empirical and true expected loss (Germain et al., 2016).
-Following (Germain et al., 2016), we assume that this error distribution is unbounded but sub-normal, i.e.
-all error fluctuations are bounded by a normal distribution of variance $s^{2}$, giving $C_{0}[\pi_{0}]<s^{2}/2$.
-We refer the reader to (Germain et al., 2016) and subsequent works (Shalaeva et al., 2020; Masegosa, 2020) for further discussion.
-For completeness, we recall here the bound used in the main text:
-${\mathcal{L}}[\pi]\leq\int_{\mathcal{P}}\hat{L}_{N}({\bm{\Theta}})+\frac{1}{N}\ln\left|\frac{\pi}{\pi_{0}}\right|{\rm d}\pi({\bm{\Theta}})-\frac{\ln|\xi|}{N}+C_{0},$ (21)
+Appendix A Glossary of terms
+Table 1 contains a glossary of the key algebraic symbols used in the paper.
+Symbol Description
+${\mathcal{E}}$ Simulation engine
+${\mathcal{M}}$ Surrogate model
+${\mathcal{X}}$ Input space
+${\mathcal{Y}}$ Output space
+${\mathbf{X}}$ Input vector
+${\mathbf{Y}}$ Output vector
+$X$ Input vector dimension
+$Y$ Output vector dimension
+${\bm{\Theta}}$ Model parameters
+${\mathcal{P}}$ Parameter space
+$N$ Number of training points
+$P$ Number of model parameters
+$\rho_{\mathcal{E}}({\mathbf{Y}}|{\mathbf{X}})$ Output distribution of simulation engine
+$\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})$ Output distribution of surrogate model for given parameter choice
+$\pi({\bm{\Theta}})$ Posterior parameter distribution
+$\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}})$ Model output distribution $\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})$ averaged over $\pi({\bm{\Theta}})$
+${\mathcal{G}}[\pi]$ Generalization error of $\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}})$ to $\rho_{\mathcal{E}}({\mathbf{Y}}|{\mathbf{X}})$ under $\pi({\bm{\Theta}})$
+${\mathcal{L}}[\pi]$ Expected loss of $\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}})$ to $\rho_{\mathcal{E}}({\mathbf{Y}}|{\mathbf{X}})$ under $\pi({\bm{\Theta}})$
+$\hat{L}_{N}({\bm{\Theta}})$ Empirical loss of $\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})$ to $\rho_{\mathcal{E}}({\mathbf{Y}}|{\mathbf{X}})$ with $N$ training points
+${\bm{\mathcal{I}}}^{*}_{\mathcal{L}}$ $X\times X$ Fisher information matrix from empirical loss
+${\bm{\Sigma}}_{\mathcal{Y}}$ $Y\times Y$ Aleatoric covariance matrix (assumed vanishing, determinant $\mathcal{O}(\epsilon^{2Y})$)
+$\epsilon$ Dimensionless measure of aleatoric uncertainty (assumed small)
+${\bm{\Sigma}}^{*}_{\mathcal{L}}$ $Y\times Y$ Error covariance matrix from loss minimization
+$\pi_{0}({\bm{\Theta}})$ Prior parameter distribution
+$\pi_{{\mathcal{L}},N}({\bm{\Theta}})$ Posterior parameter distribution from Bayesian inference
+$\pi^{*}_{E}({\bm{\Theta}})$ Posterior parameter distribution from POPS-constrained ensemble ansatz
+$\pi^{*}_{H}({\bm{\Theta}})$ Posterior parameter distribution from POPS-constrained hypercube ansatz
+Table 1: Glossary of key algebraic symbols used in the paper
 Appendix B Model selection criteria for deterministic models
-The Bayesian information criterion (BIC), used to discriminate between possible surrogate models, is derived by approximating (twice) the negative log evidence as $N/P\to\infty$, assuming a slowly varying prior $\pi_{0}({\bm{\Theta}})=\mathcal{O}(1)$.
-Using the above definitions we find
+The Bayesian information criterion (BIC), used to discriminate between surrogate models, is derived by approximating (twice) the negative log evidence as $N/P\to\infty$, assuming a slowly varying prior $\pi_{0}({\bm{\Theta}})=\mathcal{O}(1)$.
+We have
 […]
-We note that the value of the PAC-Bayes loss bound (21) satifies $BIC=2N\hat{\mathcal{L}}_{N}[\pi_{{\mathcal{L}},N}]-2C_{0}$, with $C_{0}$ model independent, as discussed in (Germain et al., 2016).
-In standard derivations ${\bm{\Sigma}}_{\mathcal{Y}}$ is set to the variational maximum likelihood solution ${\bm{\Sigma}}_{\mathcal{Y}}={\bm{\Sigma}}^{*}_{\mathcal{L}}$, yielding the familar expression $BIC=P\ln N+N\ln\|{\bm{\Sigma}}^{*}_{\mathcal{L}}\|+\ln\|{\boldsymbol{\mathcal{I}}}^{*}_{\mathcal{L}}\|$, where model independent terms are neglected.
+We note that the value of the PAC-Bayes loss bound (8) satisfies $BIC=2N\hat{\mathcal{L}}_{N}[\pi_{{\mathcal{L}},N}]-2C_{0}$, with $C_{0}$ model independent, as discussed in Ref.
+Germain et al., 2016.
+In standard derivations, ${\bm{\Sigma}}_{\mathcal{Y}}$ is set to the variational maximum likelihood solution ${\bm{\Sigma}}_{\mathcal{Y}}={\bm{\Sigma}}^{*}_{\mathcal{L}}$, yielding the familar expression $BIC=P\ln N+N\ln\|{\bm{\Sigma}}^{*}_{\mathcal{L}}\|+\ln\|{\bm{\mathcal{I}}}^{*}_{\mathcal{L}}\|$, where model independent terms are neglected (principally, terms containing only $N$ and the aleatoric covariance $\Sigma_{\mathcal{Y}}$).
 However, for deterministic regression problems, ${\bm{\Sigma}}_{\mathcal{Y}}$ is fixed.
 Neglecting model-independent terms, this gives an alternative BIC for deterministic models of
-$BIC=N{\rm Tr}({\bm{\Sigma}}^{*}_{\mathcal{E}}{\bm{\Sigma}}^{-1}_{\mathcal{Y}})+P\ln N+\ln\|{\boldsymbol{\mathcal{I}}}^{*}_{\mathcal{L}}\|.$ (23)
+$BIC=N{\rm Tr}({\bm{\Sigma}}^{*}_{\mathcal{L}}{\bm{\Sigma}}^{-1}_{\mathcal{Y}})+P\ln N+\ln\|{\bm{\mathcal{I}}}^{*}_{\mathcal{L}}\|.$ (23)
+In physical terms, this corresponds to an estimation of a partition function based on an harmonic approximation of the Hamiltonian around its minimum.
 Appendix C Summary of Laplace’s method
@@
 This encompasses all applications in this paper; the extension to multidimensional functions is straightforward.
+The method approximates the integral using a second-order Taylor expansion of $f(x)$ around the maximum of the function, which reduces the problem to a Gaussian integral.
 Laplace’s method then reads
-$\lim_{N\to\infty}\int_{a}^{b}\exp[Nf(x)]{\rm d}{x}=\sqrt{\frac{2\pi}{N}}\sqrt{\frac{{\rm d}^{2}f((x)}{{\rm d}x^{2}}}^{-1}\exp[N\max_{x\in[a,b]}f(x)].$ (24)
+$\lim_{N\to\infty}\int_{a}^{b}\exp[Nf(x)]{\rm d}{x}=\sqrt{\frac{2\pi}{N}}\sqrt{\frac{{\rm d}^{2}f(x^{*})}{{\rm d}x^{2}}}^{-1}\exp[Nf(x^{*})],\quad x^{*}=\arg\max_{x\in[a,b]}f(x).$ (24)
+For further information we refer the reader to e.g. (Wong, 2001).
 Appendix D POPS-constrained ensemble ansatz
@@
 For any choice of weights $\mathcal{W}_{E}$, an approximate optimal set ${\bm{\Theta}}^{*}_{\mathbf{X}}\in\mathcal{T}^{*}_{E}$ can be determined in a variational setting at finite $N$, using (25) in the PAC-Bayes loss bound (8) for the loss ${\mathcal{L}}[\pi_{E}]=\langle{\rm w}_{E}({\mathbf{X}})L({\bm{\Theta}}_{\mathbf{X}})\rangle_{\mathcal{X}}$.
-As $P/N\to 0$ we show in the sub-section below this reduces to the set of POPS-constrained loss minimizers
+As $P/N\to 0$, we show in the sub-section below that this reduces to the set of POPS-constrained loss minimizers
 ${\bm{\Theta}}^{*}_{\mathbf{X}}\in\mathcal{T}^{*}_{E},\quad{\bm{\Theta}}^{*}_{\mathbf{X}}\equiv\arg\min_{{\bm{\Theta}}_{\mathbf{X}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})}L({\bm{\Theta}}_{\mathbf{X}})$ (26)
-The POPS-constrained loss-minimization ensures the fits will be as close as possible under the Fisher metric ${\boldsymbol{\mathcal{I}}}^{*}_{\mathcal{L}}$, such that ${\bm{\Theta}}^{*}_{\mathbf{X}}\to{\bm{\Theta}}^{*}_{\mathcal{L}}$ for specified models.
+The POPS-constrained loss-minimization ensures the fits will be as close as possible under the Fisher metric ${\bm{\mathcal{I}}}^{*}_{\mathcal{L}}$, such that ${\bm{\Theta}}^{*}_{\mathbf{X}}\to{\bm{\Theta}}^{*}_{\mathcal{L}}$ for specified models.
 Under misspecification, the generalization error of (25) using $\mathcal{T}^{*}_{E}$ has the $P/N\to 0$ limit
-${\mathcal{G}}_{0}[\pi_{E}]=-\left\langle\ln\left\langle{\rm w}_{E}({\mathbf{X}}^{\prime}){\rm m}_{\mathcal{E}}({\bm{\Theta}}^{*}_{{\mathbf{X}}^{\prime}},{\mathbf{X}})\right\rangle_{{\mathbf{X}}^{\prime}\in{\mathcal{X}}}\right\rangle_{{\mathbf{X}}\in{\mathcal{X}}}.$ (27)
+${\mathcal{G}}_{0}[\pi_{E}]=-\left\langle\ln\left\langle{\rm w}_{E}({\mathbf{X}}^{\prime}){\rm m}_{\mathcal{E}}({\bm{\Theta}}^{*}_{{\mathbf{X}}^{\prime}},{\mathbf{X}})\right\rangle_{{\mathbf{X}}^{\prime}\in{\mathcal{X}}}\right\rangle_{{\mathbf{X}}\in{\mathcal{X}}},$ (27)
+where ${\rm m}_{\mathcal{E}}({\bm{\Theta}},{\mathbf{X}})$ is the mass function introduced in (14).
 In general we expect a uniform weighting ${\rm w}_{E}({\mathbf{X}})=1$ to be sub-optimal for generalization as the loss-minimizing fits ${\bm{\Theta}}^{*}_{\mathbf{X}}$ will be too close to ${\bm{\Theta}}^{*}_{\mathcal{L}}$.
@@
 The final ensemble ansatz parameter posterior then reads
-$\pi^{*}_{E}({\bm{\Theta}})\equiv\langle{\rm w}^{*}_{E}({\mathbf{X}})\delta({\bm{\Theta}}-{\bm{\Theta}}^{*}_{{\mathbf{X}}})\rangle_{\mathcal{X}}..$ (30)
+$\pi^{*}_{E}({\bm{\Theta}})\equiv\langle{\rm w}^{*}_{E}({\mathbf{X}})\delta({\bm{\Theta}}-{\bm{\Theta}}^{*}_{{\mathbf{X}}})\rangle_{\mathcal{X}}.$ (30)
 Equation (30) is a variationally optimized ensemble ansatz for misspecified regression problems.
 The worst case of extreme misspecification, corresponding to non-intersecting POPS ${\mathcal{P}}_{\mathcal{M}}({\mathbf{X}}^{\prime})\cap{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}})=\emptyset$, gives $\rho_{{\mathbf{X}}^{\prime}}^{*}({\mathbf{X}})=\delta_{{\mathbf{X}}{\mathbf{X}}^{\prime}}$ and thus uniform weights ${\rm w}^{*}_{E}({\mathbf{X}})=1$.
-Interestingly, it is simple to see we have uniform weighting at large $\epsilon$ as $\rho_{{\mathbf{X}}^{\prime}}^{*}({\mathbf{X}})$ is uniform.
+Interestingly, it is simple to show that optimal weights are also uniform in the large $\epsilon$ limit, as $\rho_{{\mathbf{X}}^{\prime}}^{*}({\mathbf{X}})$ is then uniform.
 For real systems we find ${\rm w}^{*}_{E}({\mathbf{X}})$ is highly non-uniform and smallest for points near to ${\bm{\Theta}}^{*}_{\mathcal{L}}$, as expected.
-However, numerical test reveal moments of the predicted ensemble errors are still underestimated, as the ”loss like” bounds still bring all points too close to the minimum loss value, which motivates the hypercube approach described in the text.
+However, while this approaches improves the description of the parameter uncertainty compared to the uniform weight variant described in the main text, numerical tests reveal that moments of the predicted ensemble errors are still underestimated, as the ”loss like” bounds still bring all points too close to the minimum loss value.
+This observation (as well as the reduced evaluation cost) motivates the hypercube approach described in the text.
 However, the max/min bounds of the ensemble, which are insensitive to our choice of ${\rm w}({\bf X})$, remain robust, as shown in figure 1.
@@
 Aside from the POPS constrained variation, this is a standard variational result which in principle allows for variational minimization of $\pi_{E}({\bm{\Theta}})$ at finite $N$ for both linear and non-linear models.
-An alternative POPS-constrained variational approach for linear regression at finite $N$ is provided in appendix LABEL:app:variational-gaussian.
 In the limit $\sigma\to 0$ we make the approximation that, for ${\bm{\Theta}}-{\bm{\Theta}}_{i}=\delta{\bm{\Theta}}$, $\pi_{E}({\bm{\Theta}}_{i}+\delta{\bm{\Theta}})\simeq(1/N)\mathcal{N}(\delta{\bm{\Theta}}|{\bf 0},\sigma^{2}\mathbb{I})$.
@@
 This result shows the POPS-constrained variation of the PAC-Bayes loss upper bound converges to the gradient of the squared loss evaluated at a POPS-constrained parameter choice, to within a positive multiplicative constant.
-As a result, in the underparametrized limit the variational optimal choice of of POPS-constrained parameter ${\bm{\Theta}}_{\mathbf{X}}$ will be given by
+As a result, in the underparametrized limit, the variational optimal choice of of POPS-constrained parameter ${\bm{\Theta}}_{\mathbf{X}}$ will be given by
 ${\bm{\Theta}}^{*}_{i}\equiv\arg\min_{{\bm{\Theta}}\in{\mathcal{P}}_{\mathcal{M}}({\mathbf{X}}_{i})}L({\bm{\Theta}}),$ (37)
 in agreement with the result given in the main text.
-Appendix E Implementation of the ansatz in Python
-The hypercube projector ${\bf V}\in\mathbb{R}^{R\times P}$ and bounds ${\bf B}\in\mathbb{R}^{R\times 2}$ can be produced with the following python code, where A.T is the transpose of A:
-1import numpy as np
-2def hypercube(X,y,prior,thresh=1e-8):
-3 # X.shape = (N,P), y.shape = (N,)
-4 # Inverse of A with prior matrix
-5 A_i = np.dot(X.T,X) + prior / y.size
-6 # Minimum loss solution
-7 T_L = np.linalg.solve(A_i,np.dot(X.T,y))
-8 # Minimum loss errors
-9 e = y - np.dot(X, T_L)
-10 # Leverage of each point
-11 X_dot_A = np.linalg.solve(A_i,X.T)
-12 h = ( X * X_dot_A ).sum(1)
-13 # Pointwise fits, shape = (P,N)
-14 T = np.multiply(X_dot_A, e/h )
-15 # Find hypercube projector
-16 nu,V = np.linalg.eigh(T@T.T)
-17 nu_sel = nu<nu.max()*thresh
-18 V = V[nu_sel,:] # shape = (R,P)
-19 # bounds
-20 B = np.array([(T@V.T).max(0),(T@V.T).min(0)]).T
-21 return V,B
-To sample the ensemble we use uniform random numbers
-1import numpy as np
-2def resample(V,B,n=20,seed=123):
-3 np.random.seed(seed)
-4 u = np.random.uniform(\
-5 min=B[:,0],
-6 max=B[:,1],
-7 size=(n,V.shape[0]))
-8 return u@V
-Appendix F Additional plots of test error prediction
-F.1 Low dimensional models
+Appendix E Implementation of the ansatz
+{algorithm}
+[h!] Find POPS-Constrained loss minimizers
+1:${\bf X}\in\mathbb{R}^{N\times P}$, ${\bf y}\in\mathbb{R}^{N}$, ${\bm{\Sigma}}_{0}\in\mathbb{R}^{P\times P}$ (prior covariance)
+2:${\bf C}\leftarrow{\bf X}^{\top}{\bf X}+\frac{{\bm{\Sigma}}_{0}}{N}$triangleright Compute regularized feature covariance matrix
+3:${\bf w}\leftarrow\text{solve}({\bf C},{\bf X}^{\top}{\bf y})$triangleright Solve ${\bf C}{\bf w}={\bf X}^{\top}{\bf y}$ for least squares solution
+4:${\bf e}\leftarrow{\bf y}-{\bf Xw}$triangleright Calculate residuals
+5:${\bf A}\leftarrow\text{solve}({\bf C},{\bf X}^{\top})$triangleright Solve ${\bf C}{\bf A}={\bf X}^{\top}$ for influence matrix
+6:${\bf h}\leftarrow\text{diag}({\bf X}{\bf A})$triangleright Compute vector of leverage scores
+7:${\bf T}\leftarrow{\bf A}\odot({\bf e}\oslash{\bf h})$triangleright Element-wise product and division for POPS-constrained corrections to ${\bf w}$
+8:return ${\bf w},{\bf T}$triangleright Return the loss minimizer and POPS-constrained corrections
+{algorithm}
+[h!] Generate POPS-constrained hypercube samples
+1:${\bf T}\in\mathbb{R}^{N\times P}$ (POPS-constrained corrections from Algorithm E)
+2:${\bf U},{\bf S},{\bf V}^{\top}\leftarrow\text{SVD}({\bf T})$triangleright Perform SVD on T
+3:$R\leftarrow\text{rank}({\bf S})$triangleright Determine rank of S
+[…]
+5:$\tilde{\bf T}\leftarrow{\bf T}{\bf V}_{R}$triangleright Project T onto the singular vectors
+6:${\bf l}\leftarrow\min(\tilde{\bf T},\text{axis}=0)$triangleright Find min bounds of projections
+7:${\bf u}\leftarrow\max(\tilde{\bf T},\text{axis}=0)$triangleright Find max bounds of projections
+8:for $i=1$ to $M$ do $\triangleright$ Generate M samples
+9:     ${\bf x}\leftarrow\text{UniformRandom}({\bf l},{\bf u})$triangleright Sample uniformly within bounds
+10:     ${\bf T}_{\text{sample},i}\leftarrow{\bf x}{\bf V}_{R}^{\top}$triangleright Project back to original basis
+11:end for
+12:return ${\bf T}_{\text{sample}}$triangleright Return M hypercube samples
+A slightly modified implementation of this POPS-constrained linear regression algorithm, wrapping the Scikit-learn BayesianRidge classPedregosa et al.
+(2011), is available on GitHub at https://github.com/tomswinburne/POPS-Regression.git.
+It can be installed (requires Scikit-learn) via pip: pip install POPSRegression.
+Figure 6: Comparision of POPSRegression against BayesianRidge from sklearn.
+Left: Mean and standard deviation of the time taken for each method to fit a $P$ dimensional linear model against $N=10000$ training points, with $P$ ranging from 10 to 1000.
+Right: Mean and standard deviation of the time taken to make $1000$ predictions with mean and epistemic uncertainties for BayesianRidge, whilst POPSRegression returns mean, misspecification uncertainty and max/min bounds.
+Appendix F Performance comparison against Bayesian Ridge
+To demonstrate the performance of POPSRegression against BayesianRidge from sklearn, we fit against the medium-dimensional cubic ”simulation engine” described in the main text, with $P$ ranging from 10 to 1000 and $N=10000$.
+Figure 6 shows POPSRegression incurs a minimal (2×2$\times$2 ×) overhead over BayesianRidge, whilst existing misspecification aware schemes do not converge in the low-noise regime.
+Appendix G Additional plots of test error prediction
+G.1 Low dimensional models
 We applied the same procedure as for the polynomical example to a model system with $P=6$ independent features to simulate more realistic datasets.
@@
 Figure 7 shows the resulting parameter distributions.
-At large $\sigma=2$ the minimizer quickly concentrates all ensemble members on the minimum loss solution; at $\sigma=1$ the ensemble distribution is notably sharper than the POPS-ensemble $\pi^{*}_{E}$, whilst for low sigma $\sigma=1/2$ the minimization only has marginal stability, but becomes closer to the POPS-bounding hypercube $\pi^{*}_{\mathcal{H}}$.
+At large $\sigma=2$, the minimizer quickly concentrates all ensemble members on the minimum loss solution; at $\sigma=1$ the ensemble distribution is notably sharper than the POPS-ensemble $\pi^{*}_{E}$, whilst for low sigma $\sigma=1/2$ the minimization only has marginal stability, but becomes closer to the POPS-bounding hypercube $\pi^{*}_{\mathcal{H}}$.
 In addition, a minimizer initialized with the POPS-ensemble did not change appreciably.
@@
 As in Figure 1 in the main text, lower $\sigma$ gave numerical instabilities.
-F.2 Medium dimensional models
+G.2 Medium dimensional models
 The plots below are identical in presentation to the top row of figures in the main text, but for the quadratic and random linear simulation engines described in the main text.